A Complete Bibliography of Publications in ACM

SIGSOFT Software Engineering Notes

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 August 2018 Version 1.00

Title word cross-reference

$24.69 [Kel96c]. α [Kok99]. K [Mus91b]. -metric [Kok99]. 0-201-87746-5 [Kel96c]. 12th [Con99]. 13th [Con99]. 17th [Tra95a]. 1993 [Ede93]. 1997 [LNS98]. 2 [Pfl95d]. 2167A [Ove90]. 21st [BBB^{+}99]. 3Rs [Yu91]. 4 [Ole93]. 4th [Rei97, Sit96].

5th [PD98]. 6th [PT94]. 7-layered [Rya96]. 7th [LW96b]. 8th [EW97, KW96]. ’9 [PBL99]. ’91 [Wie92]. 94 [Vil95]. 95 [Swa96]. 97 [Fug96]. ’99 [Ano99]. 9th [PBL99]. = [KKLK98]. A011 [TTC95]. ability [FW91]. Abstract [Jac93, Bec99, Blu93, BG94, BDM90, Cor96, Day96, Jac96, Par97, PV97, Pet96, Rec96]. Abstracting [Gun96]. abstraction [BZ99, HS95, Wil90]. abstractions [AA98, DCH97, GCBM96]. abstracts [PW95]. academia [Gri98]. accommodates [YHR90]. accreditation [ELB^{+}98]. accuracy [CC96]. Acedia [Van95]. achieve [SS90]. achieving [MB96]. ACM [Tra95b, Coo98a, Coo98b, CM98, Eic99b, Eic99a, Fin99, Lec99, Mes98, Rei97]. ACME [GM97]. acquiring [Jia92]. acquisition [Gam97, Kel96b]. Acquisitions [Add98]. acronyms [Hec90]. action [GMK99]. actions [XGG98]. active [Hua98b]. activities [JDL^{+}99]. AD [Mac91b]. AD/Cycle [Mac91b]. Ada [Che92a, DBDS93, HSWZ94, Hug91, JS90, Lai91, Sca91, YT90, Hol96]. Adaptable [Dav95a, GI90]. adaptation [IGE97]. adaptive [Bas97, MDE97]. added [O’N97]. Adding [Wil90]. ADL [CRS96]. administration [CDP95]. advanced [Joh96b]. advantage [LSJ97]. afford [Sco92]. Age [Kel97a]. agent [JDL^{+}99, LZL98, Nou96, XGG98]. agent-oriented [LZL98]. agents [HZG99, Nou98]. agility [Bru96]. agility/Maturity [Bru96]. agreement [LW99]. agreement-driven [LW99]. AI [Aik90, Nou96]. Aided [Pou93, TL90, GN92, Lin92, Tat96]. Airport [Swa96]. al [MS96]. Alexander [Lea94]. Alexandria [PSTZ99]. Algebra [Hoa93, Zel94, Zel95a, Vaz96]. Algebraic [WF99, Den96, HS96d, LMD90, Vaz96]. algorithm [CP93, HF98, NA98, NAC99, YHR90]. algorithms [Gre97, RvH91, WF93]. alias [HR96]. aliasing [ZRL96]. All-du-path [YSP98]. all-uses [MB96]. along [FF97]. ami [Kel96c]. amplifying [HV93]. analogical [LU95]. analyses [LH99b, ZRL96]. analysing [KJ97c, KP98b]. Analysis [Aye96c, MS97b, Pfl95c, RI99, AGM97, AGI98, AG98, Bal98, Bal99a, BJ93, BF92, CC96, CGK97a, CK93, CK95, CGK97b, CGHS99, CA94, DTF96, DBDS93, DC94, FH95, Gam97, GHJ98, Gor93, Hei96, HRS95, HS96c, Jac93, Jac98, Jan94, KL96, Kit97, KJ97c, LM97, Lan93, LF95, LT93, LH99b, LH99a, Lut93, MLMK97, MM96, Mus91a, NCO96, NACO97, NCC99, PY96, Pfl94, Pfl95d, Pfl95a, Pfl95b, PY93, PA96, PD90, RRL99, Smi91, ST98, SPH99, SMDP96, TvKP90, Tra92, Tra96b, UOH93, VG94, WTW^{+}91, WF93, WH99]. analyst [VPC98]. analytic [AS95]. analyzed [Rya95]. Analyzing [AG93, FGP93, RM99b, CL91, JD96, Pfl95b]. Anderson [Kim97]. Andrew [Lei98]. animals [RP98]. annealing [TCM98]. Annotated [Tra95c, Arn95, BP92, BW93, Edw94, FP94, GJ94, HLO^{+}92, Rol94, TS90, vdBKV97]. Annual [Add98, EW97, GT93, LW96b, PT94]. Anstey [Pou98]. answers [Kol98]. anyone [Buc91]. applet [pP98]. applet-like [pP98]. applicable [HF98]. Application [Woo90, BTS94, CJPV90, GCN92, Jar95, KW95, MR90, SRZ97, SAMS97, TC97, WGHS99]. application-specific [GCN92]. applications [BEKM97, BM98, Che92a, JSZ97, Joh96b, Pol90, RBDL97, pP98, Doe98a]. applied [TFW93]. Applying [Aye96c, FH95, LS90a, MSU98, NACO97, PT97, Hef93]. apportionment [AS95]. apprentice [WT91]. Approach [Kel96d, Kel96c, AL95, BK95a, Bha95, Bla97, Buk91, BGL98, CM96, DKW92, EHTE97, FB97a, Gam97, GCN92, GS95, Hei92, HHJ^{+}99, HZG99, HK97, JGS99, KSZ91, KT96, Lam97b, LS94, LZL98, LU95, MFCS95, MF96, Pet91, PW97, Ram96, SNS99, SVR97, Tom94, Tur93, Vaz96, Woo93, ZW96, Aye96b, Doe97a, Doe97c, Doe97d]. approaches [JK99]. appropriate [Kit96a, Kit96b]. Arcadia [Kad92, ROMA92]. architecting [Rec96, Tra97a]. Architectural [Sha95a, Gaf99, GAO94, MORT96, SvK98]. Architecture [Edw96, RI99, Tra95e, Wol97, AAG93, And95, BGW^{+}99, BSKH92, Clé90, DW97, FC98, Gar95b, Gri99a, Kar98, Le 96, LA98, LZL99, MT97, MHF^{+}97, Nor97, PW92, THRE^{+}94, TCY93, Tra94, WF99, GTP95]. Architectures [Add98, CW98a, Gar95a, BCGS95, DBM97, FLP99, Gac95, Jar95, Lam97a, MK96, MOT97, MQ94, NACO97, TTC95, Tra96b, Tho98]. archiving [BBP99]. Ariane [Dow97]. arising [Bly97]. Arithmetic [Fos91]. Arranga [Ray98a]. Art [Tra97a, DKLS96, Mod92, Ros96]. articulation [SMG99]. artifacts [SVR97]. asked [Tra94]. aspect [Jac93]. assess [PSTV95]. Assessing [FW91]. assessment [MS97b, Wey95]. assessments [Kra97]. assessors [Kra97]. assets [BGW^{+}99]. assistance [War92]. assistant [QJD90]. assisted [RF90]. assurance [Buk91, FB97a, Lok92, Smi91]. ASTRAL [CPK93]. asynchronous [MFR94]. attribute [TSK90]. authoring [LP97]. Automated [BKW96, BG94, GMS98, Kor96, KAY98, SSP95, TCM98, WTW^{+}91, AGM97, CR99b, CLL96, LP97, Swa96, Ter93a, Vog93]. Automatic [GBR98, Guo97, JH98, MAH^{+}98b, Shi91, VH93, BF99, Mar98a]. automatically [MiS97]. Automation [SGJ^{+}97, Jaz95]. avionics [BCGS95, Lam97b]. avoid [Boe99, JGS99, Mus91a]. aware [Ste99]. AXE [KT97].

bachelor [Sol98]. Bad [Kel96a]. baggage [Swa96]. band [Tak99]. Barbara [Eic99a]. Barriers [Dav96, Fin97]. Based [Doe97a, Kos98, LNS98, UCD99, Aik90, AG91, BGV90, BJ93, BK95a, BL97, BZ99, CC96, CABN98, CRS96, CR99b, Che94a, CW98b, Den96, DK99, DP98, EHTE97, Fek95, FGP93, FHS^{+}92, Fie99, FB97a, Hol90, JK99, Jør95, JBVW94, LW93, LvO92, Mar97, Mat97, MV96, MMS96, MTO^{+}92, NCC99, PO96, Pet91, SRZ97, SvG98, Sha95b, SW94, Spr91, SVR97, SC93b, TSK90, Ter90, THM99, TWC92, VG94, WW98, AK99, Fav96]. basis [Den96, JSZ97]. be [Hol99]. beat [GHJ98]. before [Bra90]. behaves [Rya95]. behavior [KKLK98, SPL96, Wal92]. behavioral [CL91, Kil99]. behaviorally [Mat97]. behaviorally-based [Mat97]. Benefits [Ste99, BO99, PSTV95]. better [Jaz95]. between [CA93, Gun96, HJN94, JK99, Kri95, Lam98, MNS95]. Beyond [AA95, Sha95b, SH96, O’L96a]. Biblio [Mrd90]. bibliography [Arn95, BP92, BW93, Edw94, FP94, GJ94, HLO^{+}92, Rol94, TS90, Ter90, Tri90, vdBKV97]. binary [DJJ96]. board [Bag98a]. Booch [Coo98b]. Book [Aye95, Aye96a, Aye96b, Aye96c, Aye97, Bor95, Doe97a, Doe97c, Doe97d, Doe97b, Doe98a, Fin96a, Fin96b, Hol96, Ire97, Ive97, Kel96e, Kel96d, Kel96b, Kel96c, Kel97a, Kos95, Kos98, Lec96a, Lei96a, Lei96b, Lei98, Neu95c, O’L96a, O’L96b, O’L96c, O’L97a, O’L97b, Ola98, Per96, Per97, Pou95, Pou98, Ray96, Ray97a, Ray97b, Ray98a, Ray98b, Sur97, Tra95b, Tra95c, Tra97a, Tra98a]. Books [Tra95b]. Boolean [BGL98]. both [MZ98]. bounded [WTW^{+}91]. bounding [CA93]. box [Mat92]. brewery [SC93a]. Brian [Kos98]. bridges [Pet96]. bridging [Lam98, MNS95, Tra98b]. briefly [Arn95]. Bringing [Aye96a]. Brock [MS96]. browsing [HLS^{+}99]. Bruce [Kim97]. BS [LN99]. bug [Pre99a]. Bugfind [CD90]. Bugs [Neu95a]. Building [And95, CDP95, EAD99, Kel96d, ZR96, Kad92, KSL90, LWM99, Que92, SRZ97]. built [WKP^{+}98]. built-in [WKP^{+}98]. Business [Ben96, Hef93, Vil95]. by-product [Was95].

C [Ive97, Lei98, Ray98a, BZ95, CR99a, CGK97a, HSWZ94, Hug91, Kar98, SR96, SCB^{+}99, Wyb90, O’L96b]. C/C [O’L96b]. C2 [MOT97, MORT96]. C2-style [MOT97]. CAIS [MFR94]. calculus [Fro90, MFDM97, Woo93, YCZ98]. calibrating [Ves99]. California [GS96]. call [Bru96, Kil99, Kol98, Zho90]. Cambridge [Ive97]. Can [Sco92, Ham98, Kic96, Loy93]. cancer [Bou93]. Capability [Aye95]. card [BM98]. card-oriented [BM98]. cards [GLV99]. Carroll [Ray97b]. CASE [Pou93, TL90, Ber93, DTF96, Fek95, FGP93, HS96a, KP98c, KP98a, KP98b, Lam97b, Pfl94, PK97, Rin92, Ros97, RGR92, SC93b, WVF95, XZLY95, BK95a, Bla97, Dro97b, GRW90, Hua98b, Jan94, McL93, Ole93, Pol90, Sha94, TS90, Tri90, Wyb91, YH96]. case-study [Lam97b]. case-tool [PK97, Ole93]. cases [KPP^{+}99, Mar98b, MMS96]. casts [SCB^{+}99]. Cataloguing [HSF92]. cautionary [DeT91]. CAx [Que92]. CAx-framework [Que92]. CCS [BCJ96]. CDRl [TTC95]. Center [Doe97a]. centered [Bar92, Bla97, GJ94, SHO90]. centred [Lam98]. Century [BBB^{+}99, Ola98]. certify [Kra97]. Challenge [Ola98]. challenges [Fin98]. Chamond [Ray98b]. Change [Joe97, Kel96a, Mar90, OH90, SC98, SvK98, SHO90, VN96]. change/configuration [Mar90]. changes [SC98, TSK90]. changing [MiS97]. channels [CFI94]. chaos [Rac95a, Rac95b]. characteristics [Car93]. characterization [DW97]. chart [Guo97]. Chasing [Neu95a]. check [Blu93, EGHT94]. checked [Bok99]. Checking [DJJ96, GMK99, AY98, ABB^{+}96, AG91, BFG96, CABN98, CR99a, DY94, DCH97, DP98, FLP99, GH99, GHJ98, JGS99, JXXG98, WVF95]. checklists [Bry99]. chemical [Kog95]. children [Hef93]. CHIME [DK99]. choosing [Pfl95a]. chopping [RR95]. Christopher [Lea94]. CIP [Fie99]. claim [ZS99]. claims [DS97]. Clarifying [Kra97]. clashes [Boe99]. class [PK97, ST98, SPH99]. classes [Che93b, HR94]. classification [HSF92, Lai92, LU95, MAKGM97]. classifying [MT97]. classroom [JBVW94, Kel97a]. Cleanroom [Smi96]. clinical [Jac90]. cluster [PY93]. clusters [Wau97]. CMM [Sej95]. co [LW99]. co-ordinator [LW99]. Coad [Doe98a]. Cobbler [Hef93]. Cobol [FR99, Tom94, Ray98a]. COCA [LWM99]. COCOMO [IGE97]. CODE [CHS90, Kam91, CGK97a, Dav95a, ACM97, EGHT94, FC98, GMR99, Got99, HS96a, JGS99, LSJ97, MPR99, MS97a, Par99a, PRM97, PSTV95, Shu93, WWBJ98]. coding [Bra90]. CoffeeStrainer [Bok99]. coffin [MAKGM97]. Cohesion [BK95b]. Collaboration [WBG^{+}99]. collaborative [Her99, KTC^{+}92, MAM93, MFR94]. collection [Agg97]. combining [BF99, Mar98b, WRBM97, XJG98]. Coming [Pet91]. comment [MS96]. Comments [Mac91a, Off90, YH96]. commercial [Tra98b]. committee [Not96]. communication [CFI94, Che92c, HJ99, LA98, NCO96, Ter93a, VPC98]. Communications [Rei97]. community [Edw99, HKMR98]. compact [Cor98]. companionship [BdLLvdS90]. Comparative [Sim96]. Comparing [SRLZ98, WRBM97, MT97]. Comparison [JGS99, Den96, Loy90, SC93a]. compatibility [CL91]. competence [And95]. Compiler [O’L97b, BTS94, Kar98]. compilers [CD90, VB97]. complete [Vaz94, XDH^{+}98]. Completeness [mWE91]. Complex [Nor97, Tra96a, BTS94, SS95]. complexity [KBZP98, Law98c, Rac95c]. Component [LNS98, SW94, UCD99, AGI98, Fie99, HJ99, JK99, MAKGM97, SN90, Whi95]. component- [Fie99]. Component-Based [LNS98, UCD99, SW94, JK99]. Components [CE99, Edw96, Joh97, BHKW94, CDP95, Dai95, EHL^{+}94, HC99, HSWZ94, HZG99, MiS97, MOT97, PW95, Sug95, Tra97b, ZW95]. componentware [Kim98]. composability [CPK93]. composite [BGL98]. composition [AA98, MDE97, MQ94, WBG^{+}99, KLS99]. Compositional [BFG96, CK95, JXXG98, Tho98, CK93, CGK97b, CA94]. comprehension [MRS98]. comprehensive [HHJ^{+}99]. Computation [HRS98, HR96]. Computer [Con99, GN92, Kri95, Neu95a, Nor97, Pou93, Rei97, TL90, Gor94, JBVW94, Lin92, Neu92, Neu94, Neu96a, Tat96, War92, Tra95b]. Computer-Aided [Pou93, TL90, GN92, Lin92, Tat96]. computer-based [JBVW94]. Computerized [Kel97a, Day96, MS97b]. Computers [Neu95e, Neu95f, Neu95b, Neu90b, Neu90c, Neu91a, Neu91b, Neu93, Neu95d, Neu95g, Neu96b, Neu96c, Neu96d, Neu97a, Neu97b, Neu97c, Neu97d, Neu98a, Neu98b, Neu98c, Neu98d, Neu99a, Neu99b]. computing [NAC99]. concept [Bal99a, Pet96, RP98, ST98]. concepts [KGL98]. conceptual [CS95, CFF93, Kro98, Tra91]. concerning [Zuc94]. concurrency [CW98b, LT93]. concurrent [BMS93, CC96, CA93, Cor96, Cor98, DS94a, DY94, DC94, GPS96, KT96, Lai91, NAC99, Que92, XGG98]. conditions [Fek95]. conducting [Hol90]. Conference [Con99, Dov97, Nor97, Rei97, Tat96, Tra95a, Tra95f, Ano99, CQW^{+}95, PD98, Sit96, Sit97]. conferences [Tra97c]. Confessions [Lei95, Tra95d]. Configuration [Kob90, MR90, Som96, Gun96, HM97, Joe97, ML90, Mar90, SNS99, Spr91, Zel95b]. Configuring [KRT97]. congruent [Bro90, Ket90, Lot90]. connected [Bla97]. connectors [DR97]. consensus [MR90]. consequence [Wey95]. conservative [NA98]. considered [Hec90, PB94]. Consistency [FLP99, HK92, QJD90]. constraint [GBR98]. constraints [Bok99, BGL98, CK93, CK95, Day96, Kel96a, XP91]. Constructing [Cor96, Cor98, Fro90, Sug95]. Construction [Hol96, Che91b, Che91c, Fie99, HMR93, HZG99, JZZ^{+}95, LS94, Mar98a, Rin92, O’L97b]. constructs [Smi91, SHO90]. containment [Kel97b]. content [HM97, MM96]. content-derived [HM97]. context [CK93, Kit96c, Rem97, SPL96, SRLZ98]. contexts [Kob90]. contextual [PW97]. contribute [Kic96]. contributions [BK95a]. Control [Arn95, Bal98, Che92c, HRS98, Jac90, LHR99, MAM93, RM99a, Sha95b, WP92]. controlled [MAKGM97]. controlled-vocabulary [MAKGM97]. convenient [Kil91]. Conventional [Kil91, Dro99]. ConversationBuilder [KTC^{+}92]. conversations [BO99]. conversion [Sar93]. cooperation [Bar92]. Cooperative [Kri95, BL97]. Coordinating [JDL^{+}99]. coordination [BZ99, Ver96]. COOZ [XDH^{+}98, YCZ98]. Coping [SCB^{+}99, TSK90]. core [Sim95]. Corporate [CHS90]. correctly [Rya95]. Correctness [Bro90, Ket90, Lot90, MQ94]. Cost [BG90, Ari93, DKW92, GI90, Gor93, IGE97, JBVW94, PSTV95, RW96]. cost-benefits [PSTV95]. cost-effective [DKW92, Gor93]. cost-effectiveness [RW96]. costs [BO99, VG94]. COTS [Fav96]. COTS-based [Fav96]. counter [BFG96]. counter-examples [BFG96]. counterexample [JD96]. coverage [Agr99, DS97, LGRM91, MB96, YSP98]. Coyle [Ray98a]. Craft [Kos98]. Creating [Amo95, BCGS95, Lam97a]. Creativity [Neu95c]. Crisis [Ola98, Joh96a]. criteria [ELB^{+}98, Kit96a]. Critical [DPS96, Sef94, AGM97, BF91, Che92a, Gor94, MMS96, RHR91, RvH91]. criticality [Ebe97]. criticisms [KL90]. cross [LW99]. cross-organisational [LW99]. Cryptographic [DS97]. CSCD [LZL99]. CSCW [Joh96b, Kri95]. CSE [Con99]. CSEET [Con99]. Cultures [Gri96]. Current [Dav96, Fin97, JG96]. curriculum [TTC95]. curve [Rac96]. customer [McL93]. Customization [AA98]. customized [LP97]. Cute [Hec90]. Cycle [Mac91b, Car93, HSE93, Rac95a, Sim95]. cycles [BG90]. cyclotron [Jac90].

Dagstuhl [GTP95, THP93]. daily [Ano99]. Daistish [HS96d]. Data [DC94, RRL99, Agg97, BTS94, BGS97, BM99, BG94, BDM90, CGK97a, Che92c, Cla90, DJ98, GBR98, Gri93, GCBM96, Guo97, GMS98, HF98, HR94, Kor96, LH99b, MF96, MV91, NCO96, NA98, NCC99, Pfl95b, Raj98, RHR91, Rem97, SS90, Ste99, Tur93, Vaz94, Vaz96, Ves98, VB97, WWBJ98, XDB99]. data-driven [Cla90]. Data-flow [RRL99, LH99b, RHR91]. data-sensitive [WWBJ98]. Database [Sar93, JSZ97, TTB^{+}90]. Databases [Pou95]. dataflow [HRS95]. David [Pou98, Eic99b]. Davies [Ire97]. day [Kim98]. DCF [Sej95]. dead [CGK97a]. Deadline [Tra98a]. deadlock [DBDS93]. Debugging [Bec91, CD90, DE91a, DE91b, GBF99, JGS99, Law97, Mac91a, Shi91, Lei96a]. decision [Ada99, BZ99, DJJ96, KBZP98, MSU98, Tak99]. decisions [FP98]. Declarative [HLS^{+}99]. decomposition [ZRL96]. Decoupling [VN96]. dedicated [GLV99]. deduction [Vaz96]. deductive [AGM97]. def [SPH99]. def-use [SPH99]. Defect [Neu95a, Kel97b, LSJ97, WRBM97]. defects [Rin96]. define [Sha94]. definition [Bok99]. degree [Sol98]. Demand [HRS95]. DeMarco [Tra98a]. democracy [Buc91]. Dennis [Coo98a]. dependability [Ham94, Ham96]. dependence [HMR93, HRS98]. dependences [JR94]. dependencies [Gun96]. depends [HB94]. deployment [Sar91]. derivation [BdLLvdS90, Red90]. derived [HM97]. Deriving [CDHW93]. describing [Che93b]. Description [Wau97, CJPV90, MT97, Sar92c, Whi95]. descriptions [AAG93]. deserve [TBK92]. Design [Aye96a, Bur96, Doe97c, Doe97d, Kim97, Kog95, Pfl94, Pfl95c, Sch92a, Tra96a, Ana98, Bra90, Coo96, Day96, Dro96, DR97, Fis93, GAO94, Gar95b, Hua98a, JD96, Jac98, KSZ91, Kok96, KS91, MLMK97, Mar98b, MSU98, MS96, McL91, McL92, MORT96, MDE97, Pfl95d, Pfl95a, Pfl95b, PT97, PSTZ99, RRG97, RDO98, Rin93, RW97, SS95, SvG98, Sha95b, Smi91, Son95, SO97, Ter93b, TC97, TvKP90, TWC92, VN96, WT91, vR92, FvL94, KW96, Pou98, Ray98b, Aye96c]. designers [BDM90, Lea94]. Designing [Blu93, LHR99, Gri99a, KP98a, SAMS97]. designs [BF99, CDHW93, KRT97, LSJ97, YT90]. desk [PSTZ99]. desktop [RM99a]. DESMET [Kit98]. detecting [FW91, NA98]. detection [CGK97a, CW98b, LSJ97, TRC93, WRBM97]. detector [JD96]. determine [PP97]. deterministic [AL93]. Deutsch [Fin99]. developer [Gam97]. Developing [BEKM97, CCR90, Hen95a, ROMA92, Sca91, Tra97b, Dro97a, Doe97a]. Development [CHS90, CW98a, Dai95, DS94b, Doe97b, FCC94, Fin96a, Lei96a, Ray96, SFE97, AK99, AL95, BG90, BGV90, BCC^{+}99, Bec99, BSKH92, BK95a, CCCL90, Car93, Che94a, CS90, Con92, Dan91, DS94a, Dol91, DK99, DE91a, DE91b, Ebe97, FP98, GRW90, HK92, HS95, Kad92, KTC^{+}92, KT97, Kim98, KSZ91, Law98c, Loy90, Lyu91, MAM93, Mar97, MV96, MR92, Mrd90, OAFG98, PSTV95, RM95, Rin91a, RF90, SSP95, SC93a, SS90, SAMS97, SARL91, SVR97, Ste93, SC98, TTC95, Ter93a, Ter90, Tri93, Vaz93, WGHS99, Wyb90, XJG98, XZLY95, ZAD^{+}97, vM91, Müh96]. developments [Gol90, LS90a]. diagnosis [Kel96a]. diagnostic [Sar92a]. diagram [Guo97]. diagrams [DJJ96, Vaz94, Vaz96]. dialog [ND90]. Dichotomy [Dav95b]. Dictionary [O’L97a]. differences [HJN94]. Digital [PSTZ99]. Dijkstra [Ter93b]. dimensional [XJG98]. dimensions [DvK95]. Direct [Gri93, Smi91]. directed [MZ98]. direction [Tri91]. Directions [Ang99, THP93, FF97, GJK95, JG96, LA98]. Directory [Tra95c]. discipline [Bec99, Gar95b, OSWZ94]. discovery [JDL^{+}99]. discussion [Car93, Che92a, Tri92]. disseminating [Sea97]. Distributed [UCD99, AC95, AA98, BDG^{+}95, CK95, Clé90, DK99, JG96, Kem97, KSL90, Sha94]. distribution [Tak99, pP98]. Do [HB94, BO99, CS98, Kra97, Mai97]. doctorate [SJ96]. Document [Lai92, Bla97, LP97]. document-centered [Bla97]. documentation [Han98, HGGM98, Mar94, Mar98a]. documenting [HS90]. documents [Amo95, BBP99, Lai96]. DoD [Ove90, Pol90, SMDP96, Tra98b]. DoD-Std-2167A [Ove90]. Does [McL93, Vot93, Tra90, Zel99]. Domain [Bai96, CJPV90, Gom95, PD90, Sim96, Tra92, Tra94, Tra95e, CDP95, FH95, FLM95, Gac95, Gam97, HF98, Hen95a, LW96a, LM97, LF95, LZL99, LU95, MHF^{+}97, Rol94, Sim95, SMDP96, TTC95, THRE^{+}94, TCY93]. Domain-Specific [Tra95e, Tra94, LW96a, TTC95, THRE^{+}94, TCY93]. domains [Jar97]. dominator [Agr99]. Donald [O’L97a]. DOTSS [Gam97]. down [War96]. downcasting [PR99]. Draft [ELB^{+}98, ACM97]. driven [AG91, Cla90, DvL96, DS94b, LT93, LW99, Ram96].

Druffel [Lec99]. DSSA [Tra94, Tra95e]. DSSG [XZLY95]. Duality [Dav95b]. duals [MB96]. Duplication [CFI94]. Duration [YCZ98]. Durham [BB96]. during [Lut93]. Dynamic [Gor93, MK96, Bal99a, GS95, HLS^{+}99]. dynamically [FC98, MiS97]. dynamically-linked [FC98]. dynamics [Pre99a].

early [Bou93]. easy [HLS^{+}99]. EDCS [SS95, Tra96a, Wau97]. editing [BGV90, Kob90, WW98]. Edition [Ray97b]. editor [Lec96b]. editors [TWC92]. Edmund [Ray98a]. Educating [SJ96]. Education [Con99, Lud96, Sea97, Kel97a]. educational [RS97]. Edward [Doe97c, Doe97d, O’L97a]. Effective [AG98, DKW92, Gor93]. effectiveness [FI98, RW96]. effects [HS96d, SRLZ98, Tak99]. efficiency [CABN98, LH99b]. Efficient [Agr99, CGHS99, HMR93, LH99a, GBF99, KL96, NAC99]. effort [HS96a, Ves99]. Egret [Joh96b]. EIFFEL [Rin92, Fin96a]. elaboration [DvL96]. electronic [BBP99, BO99, LWM99]. Elements [JD96, DR97]. elevators [GAM95]. elicitation [Hol90, VPC98]. Embedded [Mar98a, Fie99, Mar94, THM99]. emphasized [Boe95]. empirical [FI98, Hen95b, MR92, TWC92, WRBM97, BK99]. Empowerment [Bor95]. enabling [HC99]. enactment [ACM90]. encapsulating [BG94, BM98]. ENCOMPASS [Ter90]. encountered [Kad92]. enforce [Gor93]. Eng [KJ97c, KP98a]. Engineer [Sar94, Joh96a]. Engineered [Sar94]. Engineering [Bai96, Con99, Doe96a, Doe96b, Doe96c, Doe97e, Doe97f, Doe97g, Doe97h, Doe98b, Doe98c, Doe99a, Doe99b, FD96, GM96, Hol95, How95a, How95b, How96, Lud96, Nor97, Pet96, Pfl95c, PSP95, PP96, Pou93, SD94, TL90, THP93, Tra95c, Tra95a, Ang99, Bag98b, Bar98, Boe96, BP92, Cas99, CD94, Che94b, CQW^{+}95, DKW92, Del90, DH98, ELB^{+}98, Fin92a, FLM95, Fug99, GJ94, Gar95b, Got98, Got99, Gri99a, Gri98, Hef93, Her99, HGGM98, Hol99, HK97, Hug91, JR94, JSZ97, JG96, Kai93, Kic96, Kit96c, Kit96a, Kit96b, KJ97a, KJ97b, Kit97, KP98c, KP98b, Kit98, Kog95, KT94, KRS98, dPL96, Lin92, LS90b, Lok92, LN99, Lyu91, Mah98a, Mai97, Mai95, MTO^{+}92, OM95, Par97, Par98a, Par98b, Pfl94, Pfl95d, Pfl95a, Pfl95b]. engineering [PD95, Que92, Rac97, Rac98, Rin93, Rol94, RGR92, SK96, Sar92b, Sco92, Sha98, Sol98, Spr91, Sta93, TC93, Tat96, TTB^{+}90, TCY93, Tri91, Tri92, War92, Woo93, Xia98, Yu91, Zuc94, vdBKV97, BK99, Sur97]. engineers [Bag98a, Bag98b, Mai97, Rac96, Uhl97]. enhancement [Mar98b]. Enhancing [CK93]. enough [GJJ^{+}95]. enriched [Kol98]. enterprise [LW99, SD94]. entity [DS94b, Kil91]. entity-relationship [DS94b, Kil91]. Environment [Ben96, CHS90, FCC94, TL90, Bla97, BF92, CD94, Che94a, DK99, DE91a, DE91b, GM97, Gru90, HS90, IWJ90, Jor90, Kad92, Kar98, KNN99, ML90, Mar97, MDE97, MTO^{+}92, OAFG98, OM95, RM95, ROMA92, Sha94, SS90, Ste99, SN90, Ter90, Vog93, Zuc94, Che92a]. environmental [PT97]. Environments [Dav96, Fin97, BSKH92, BP92, CB99, Clé90, Dro97b, Fav96, FHS^{+}92, GJ94, GI90, GCN92, GAO94, Gom95, Que92, SHO90, TC93, XDB99, You90]. Epistemology [Hol95]. Equivalence [LH99b]. error [DTF96, Gor93]. error-propagation [Gor93]. errors [CFI94, HF98, Lut93]. ESA [Sch92b]. Escaping [Boe99]. ESEC97 [Sch97]. ESEC97/FSE5 [Sch97]. essential [Kil94b]. Essentials [Pou95]. Estelle [CS90]. Estimating [Lai96, MI91, MZ98]. estimation [Ari93, BG90, CS97, Ves99]. ethics [Got99]. European [Mon97, FCC94, Rem97]. Evaluating [Kit96c, Kit96a, Kit96b, KJ97a, KJ97b, Kit97, KJ97c, KP98c, KP98a, KP98b, Kit98, SK96]. Evaluation [SH98, CR99b, Kit96c, Kit96a, Kit96b, Kit97, KJ97c, KSZ91, Tri90]. Event [CW98b, AD94, AG91, HS96c, Mar98b, RM99a, RW97]. Event-based [CW98b]. event-driven [AG91]. events [CA93]. Every [Per97, Vot93]. Evolution [CW98a, SPL96, SS90, WT91]. Evolutionary [SS95, vM91, Tra96a]. evolving [CFF93, SD94]. examining [BF92]. example [BCGS95, Tra95e]. Examples [Hol96, BFG96, Jia92]. exception [Iss91, RM99b]. exceptions [BM99, CGHS99]. Executable [DR97, Nes90, Ter90]. executive [Not96]. executives [SJ96]. exercise [Rya96]. existing [Sug95]. expect [Uhl97]. Experience [CS90, WG90, Apt94, Dai95, DBM97, FH95, HK97, JS90, Lam97a, LM97, LMD90]. Experienced [Doe97a]. Experienced-Based [Doe97a]. Experiences [Ebe97, Hei96, KGL98, Wyb90, Hen95a, Jar95, KSL90]. experiment [GM97, KL90, LSJ97, MZ98, Pfl95d, PSTV95, SH98, VPMM95]. Experimental [Pfl95d, Pfl95c, Pfl95a, Pfl95b, BF91, CR99b]. experiments [CS97, Pfl94]. expert [Dai95, Sch92b, WTW^{+}91]. explicit [MDE97]. Exploitation [Mil93]. Exploiting [Gac95, GAO94]. Explore [Kri95]. exploring [RM99a]. exponential [MI91]. exposure [Mus91b]. Expressing [PRM97]. extended [Day96, Par97]. Extending [KPP^{+}99]. extensible [Jor90, Kar98, KF98]. extension [XDH^{+}98]. extraction [MN95]. Extreme [Bec99]. Eykholt [O’L97a]. f [Bur96]. F3 [CD94]. FACADE [GLV99]. FACE [MDE97]. faceted [MAKGM97]. Facility [FCC94]. factor [RGR92]. factors [KJ97a, SK96, Sef94]. fail [Dro97a]. fails [WJ99]. failure [Bly97, Dow97]. failures [MI91, Wey95]. fair [GMK99]. Families [CW98a]. FAQ [Tra94]. fast [War96]. Fatal [Neu95a]. Fault [Fin96b, DPS96, FW91, Mus91b, TRC93]. fault-detecting [FW91]. Faults [HV93, CS98, DTF96, OH96]. favorite [DR96]. feasibility [Jan94]. Feasible [FB97b]. Feature [AG95, JD96, Kit97, KJ97c, PA96]. features [HS90, KJ97b]. February [THP93]. federal [Zuc94]. federation [EAD99]. feeding [RP98]. Fellow [Coo98a, Coo98b, CM98, Eic99b, Eic99a, Fin99, Lec99, Mes98]. Few [Bri95]. Fifth [GS96, GT93, Mon97]. Fifty [Rac97]. file [VH93]. Filter [DP98]. Filter-based [DP98]. filtration [RM99a]. filtration-control [RM99a]. Finding [Mah98a, Bra90, TCM98]. Fine [MAM93, SS90, MPR99]. Fine-grained [MAM93, MPR99]. finite [CK95]. finite-state [CK95]. Firesmith [O’L97a]. First [Add98, Gar95a, PSP95, Smi96, Tra92]. fixes [Kel96a]. flames [Not97]. flaw [TCM98]. flexibility [HHJ^{+}99]. flexible [Kad92, pP98]. flight [Mod93, WH99]. flow [Bal98, BGS97, Che92c, DC94, Gri93, Guo97, HR94, LH99b, MF96, NCO96, NA98, NCC99, RHR91, RM99b, RRL99, Ste99, SRLZ98, TRC93, Vaz94, Vaz96]. flow-aware [Ste99]. FOOlish [Hec90]. form [AS94, KKLK98]. Formal [ACLN95, AGI98, BCJ96, DvL96, Jac90, Kil94a, Lev92, NLS90, Pla90, Red90, Ros96, RvH91, AD94, Amo95, BF99, CDHW93, CLL96, Gar90, GPS96, Jia92, Kok99, KF98, OM95, PF95, Pfl94, TvKP90, Woo90, XJG98, ZZW98]. Formalised [RF90]. formalism [Che91a, Che93b, IWJ90, PY96]. formalisms [Che92c, FM95].

Formalization [KW95]. Formalizing [DBM97, Hua98a, BMS93, Sim95]. formatted [Dro99]. formula [Ada99]. Fortran [Pre99b]. Foundation [PSP95, PP96, JC95, Bar98]. Foundations [GM96, Ham94, LNS98, PW92, O’L96c]. fountain [Pil96]. Fourth [GM96, Sit97]. fragments [RRL99]. Frailey [Coo98a]. frame [MDE97]. Framework [Sha94, Sim96, BMS93, Bla97, BM99, BM98, CH95, CS95, CFF93, FHS^{+}92, HK92, Joh96b, LP97, MT97, OSWZ94, PF95, PR95, RW97, Son95, SC93b, TC97, XDB99, Que92]. frameworks [Gar90, GJJ^{+}95, Joh97, MHF^{+}97, MSB97, SSP95]. Frank [Ray98a]. frequently [Tra94]. FSE [Not97]. FSE4 [GM96]. FSE5 [Sch97]. full [CLL96]. fun [Not97]. Function [Cla90, GC96, KKLK98]. function-oriented [GC96]. Functional [Fin97, BCJ96, MMS96, OM91, SH98, VPC98]. functionality [Sha95a]. Fundamental [Bri95]. Further [FI98]. fusion [Nou98]. Future [THP93, LHR99, LA98, Tri91, Tri92]. fuzzy [JSZ97].

G [Eic99a, O’L97a, Tra95b]. gap

[Lam98, MNS95, Rac95c, Tra98b]. gateways [LW99]. gauntlet [Fug96]. general [LH99b, Vog93]. Generalization [Law98a, SR96]. generate [GH99]. Generating [DR96, MMS96, GCN92]. Generation [PY96, Ber93, GBR98, GMS98, HF98, Hua98b, Jar95, JH98, KL96, Kob90, Kor96, KAY98, MAH^{+}98b, Raj98, SO97, WBG^{+}99, War92]. generative [CE99, JK99]. generator [SRZ97, TC97, VH93, VB97]. generators [JPB97]. Generic [JSZ97, BN93, ML90]. GIDTS [KNN99]. Glass [Neu95c]. global [Agr99, AG95]. go [GJK95]. goal [DvL96]. goal-driven [DvL96]. Good [Edw96, Coo96]. Got [LWM99]. governed [MR90]. GQM [AK99]. GQM-based [AK99]. Grady [Coo98b]. grained [MAM93, MPR99, SS90]. grammar [CJPV90]. grammars [Dol91, Le 96, TSK90]. graph [Dol91, EHTE97, Le 96]. graph-grammars [Dol91]. Graphic [Che92c]. graphical [AB97, Bol90, CC90, DCH97, KNN99]. graphs [Agr99, HMR93]. Gries [Ter93b]. Gries/Dijkstra [Ter93b]. group [EW97, GT93, GRW90, LSJ97, LW96b, PT94, Sea97, Tak99, Tra92]. groupware [BBP99]. growth [MZ98]. GUI [BR94, BR95, OAFG98]. Guided [DHvL^{+}95]. Guideline [mWE91]. Guidelines [Aye95, Lei96b, Mar94].

Habermann [Not93]. Haim [Aye96b]. Hall [Doe97a, Ire97]. Handbook [Kel96c]. handling [Iss91, ZS99]. happen [NA98]. happier [McL93]. hard [XP91]. hard-real-time [XP91]. hardware [Sca91]. Harel [Wal92]. Harlan [Poo96]. harmful [PB94]. Haskell [Fin97]. Hatley [Tat95]. Hatley/Pirbhai [Tat95]. Hawk [WH99]. Hayes [Ola98]. HCI [Kol98, Sea97]. HCI-enriched [Kol98]. heart [GHJ98]. heart-beat [GHJ98]. hedging [BO99]. Heisenberg [Lap90]. help [Bal99b, BDM90, Loy93, PSTZ99]. Herbal [HC99]. Herbal-T [HC99]. here [GJK95]. heterogeneous [LA98, WP92]. heuristics [LM97]. hierarchical [AY98, Gir99, HW97, Hua98a, XDB99]. hierarchies [ST98]. hierarchy [AS95, PK97, RP98]. high [AL95, JZZ^{+}95, MNS95, Pou98]. high-level [AL95, JZZ^{+}95, MNS95]. Highlights [Tra98b]. history [Bar98, Mah98a, Rin91b]. HLA [AGI98]. holes [Mus91a]. Holistic [Ana98]. Holmes [O’L97b]. HOOD [ZZW98]. Horning [Mes98]. human [KJ97a, SK96]. Hybrid [GS95, Bha95, Vaz93]. hypercubes [Bec91, Mac91a]. hypermedia [FHS^{+}92]. hypermedia-based [FHS^{+}92]. HyperNet [Mar97]. Hypertext [Dro97b, Kil94b]. HyperWeb [FHS^{+}92].

Ian [Ola98]. IBM [Doe97a]. ICARUS [DHvL^{+}95]. ICECCS’97 [Nor97]. icon [Dol91]. ICSE [Ano99, Fug96, Ghe99, Par95, Tay97]. ICSR4 [Sit97]. ICSR’5 [PD98]. ideas [Not97]. Identification [Vaz94]. Identifying [FR99, KJ97b, Sea97]. IEEE [Ede93, Mam94, Nor97]. If [Sar94]. II [EHL^{+}94, GAM95, Hei96]. III [BHKW94]. illusion [Was95]. illustration [LF95]. Illustrative [Neu92, Neu94, Neu96a]. immaturity [Fin92b]. impact [Gor93, GGL^{+}97, Pil96]. implement [Gam97, Vaz93]. Implementation [BJ93, CP93, KM91, MLMK97, PD91, RDO98, Smi91]. Implementation-based [BJ93]. Implementing [AB97, YT90, BHKW94, Sef94]. implications [Dro97b]. implicit [GJND98]. important [RGR92]. impressions [Vil95]. improve [KRS98, LH99b, NCC99, Sco92]. Improved [WF93]. improvement [ADLK97, FP94, Han98, Hef93, Ves98]. Improving [Aye95, CC96, CABN98, Che93a, Fis93, KGL98]. in-the-large [MMS96]. inadequacies [MV91]. Including [Kos98]. inconsistent [Ves98]. incorrectly [Rya95]. increased [CP93]. Incremental [KT97, Kar98, KT96, Ter90]. incrementality [FM95]. indecision [Par99b]. independence [SN90]. independent [BDG^{+}95, LS94]. Index [Tra95c]. Industrial [Boe96, Agg97, DBM97, FH95, GPS96, LMD90]. industry [Bru96, Gri98, Uhl97]. Inertia [Shu93]. infeasibility [BF91]. infeasible [BGS97]. influence [KJ97a, SK96]. influential [HKMR98, Par95]. Information [Bor95, Kel97a, SGJ^{+}97, Ver96, AG95, BGS97, BDM90, Con92, Dov97, Dro96, Dro97a, Dro99, FF97, GS95, HR96, Hef93, Hen95b, MM96, NAC99, Pou97, Sar92a, SAMS97, TRC93, Aye96b]. infrastructure [AG95]. Inheritance [XZ95, AB97, BZ95, LXK98, MD93]. initial [Lam97a]. inject [CS98]. innovative [BA94]. insertion [CFI94]. insights [Apt94]. Inspection [FD95, Ada99, Bry99, MFR94, Vot93]. inspections [BW93, KM91, PSTV95]. Institutionalizing [Dav96, Fin97, Lei95, PBL99]. integer [BGL98]. integrate [Che92c]. Integrated [Doe97c, Doe97d, Pou93, BGV90, GI90, Gel94, Joe97, MFCS95, PW97, RM95, SVR97, Vog93]. Integrating [AD94, FF97, Kel97b, Kro98, ML90, WP92, ZZW98, ROMA92]. Integration [DFKM97, Dro99, KBZP98, AGI98, Arn95, BP92, BF92, GI90, GM97, HC99, Son95, SN90, Wyb91, YHR90]. intelligent [Nou98]. Inter [SPH99]. Inter-class [SPH99]. Interacting [Law98b]. interaction [VPC98]. interactions [AG95, PA96]. interactive [AD94, BM98, Che91a, PSTZ99]. interconnected [Bol90]. Interconnecting [FM95]. interconnection [MTO^{+}92]. interface [ACLN95, CCCL90, Fis93, PT97]. Interfaces [CCCL90, BL97, DCH97, TBK92]. intermediate [GLV99, Jac98]. intermediation [VW99]. International [CW98a, FvL94, Gar95a, KW96, Lud96, Nor97, PSP95, PP96, RI99, Som96, SFE97, Tra92, Tra95a, Tra95f, Wol97, PD98, Sit96, Sit97, KLS99, UCD99]. Internet [Ara95, Bru96, HC99, RW97]. Internet-scale [RW97]. Interoperability [SMG99, BKW96, BM97, HC99, MAH^{+}98b]. interpreters [Fro90]. interprocedural [HRS98, HRS95, RR95]. Intrinsic [Car96]. Introducing [GRW90]. Introduction [Fin97, Lei98, Lea94, PD90]. Invariant [ND90]. invariants [JH98]. Investigating [DJ98]. investigation [TWC92]. investment [Ber96, FP98]. Invitation [Jaz99, Sch97]. invited [Bec99, CE99]. invocation [GJND98]. involving [DTF96]. ISAW [Wol97]. ISAW-2 [Wol97]. ISBN [Kel96c]. ISEW [Smi96]. isn’t [CCR90]. ISO [Kel96d]. ISSTA [HKMR98]. issue [Aha99]. Issues [Bly97, CLL96, Del90, Kad92, SAMS97, Ves99, Zuc94, Kem97, Kit96b, Lyu91, Sha95a, Sta93, WBG^{+}99]. items [Gun96]. iterative [GMS98]. IV [HSWZ94, Mod93]. IWSEE3 [Lud96].

J [Coo98a]. James [Aye96b, Mes98]. Java [Bok99, CGHS99, Cor98, GGL^{+}97, NAC99, RM99b, SvG98, Tra97b, pP98]. Jay [Mes98]. Jeff [Kos95]. Jim [Ire97, Mes98, O’L97b]. John [Per97]. Jonathan [Lec96a]. Jowett [Lec96a]. JR [XZLY95]. JR-DSSG [XZLY95]. Jr. [Lei98]. JSD [YT90]. just [Sha95a].

kernel [BCJ96, SARL91]. key [ZW93]. kick [Tra96a]. Kidd [Kos95]. Killer [Neu95a]. Kilov [Aye96b]. Knight [KL90]. KNOW [Per97]. Knowledge [Ray97a, VW99, Hen95a, JDL^{+}99, MV96]. Krief [Aye97].

L [Ive97, Neu95c, Tra95c]. L. [Fin99]. L.0 [Nes90]. laboratory [LSJ97]. Language [SHO90, BGV90, GLV99, Jac98, LS94, LMD90, Mac91a, Nes90, Pfl94, RHR91, SO97, TWC92, VB97, WW98, WBG^{+}99]. language-based [BGV90, TWC92, WW98]. language-independent [LS94]. Languages [Bur96, Kim97, AB97, Ang99, De 90, Kic96, LHR99, MT97, Whi95, Wil90]. Larch [WG90]. Large [KLS99, ADLK97, ABB^{+}96, GCBM96, LS90a, LW93, MMS96, PSTV95, SNS99]. Large-Scale [KLS99]. Larry [Lec99]. law [MR90]. law-governed [MR90]. Laws [Bri95]. layer [BCJ96]. layered [Rya96]. lazy [Fro90]. LCLint [EGHT94]. Leanard [Tra95c]. leaner [Law98a]. learn [Ham98]. learned [BA94, GM97, LHR99, Tra95d]. learning [Ber93, Ell98, MSU98, Rac96]. legacy [BGW^{+}99, BM98, Mar98b, Pre99b, Zuc94]. Lehman [CM98]. less [BO99]. Lessons [GM97, BA94, Hei96, Kad92, LHR99, Tra95d]. Letter [Lec96b, Not96]. level [AL95, Gos94, JZZ^{+}95, MNS95, Nou96, SvK98]. levelled [Kok96]. Leveson [KL90]. libraries [BSST93, FKV95, HSF92, Hen95b]. Library [PSTZ99, BM97, Jar95, LW93]. license [Bag98a, Bag98b]. life [BG90, BF91, Car93, HSE93, Par99b, Sim95]. life-critical [BF91]. life-cycles [BG90]. lifecycle [KPP^{+}99]. Lightweight [MN95, BR95]. like [pP98]. limit [Bal98]. limitations [De 90]. Line [Add98, BGW^{+}99, SvK98, Tra98b]. Linear [HS96c]. lines [Tra98b]. linguistic [CJPV90, Kel96a]. link [Kai93]. linked [FC98]. linking [FC98]. links [Kil94b]. lists [Mar98b]. little [Gre97]. Liu [Ray98b]. Liveness [Fek95, CGK97b]. Load [AL93, YP96]. localization [DPS96]. Location [BDG^{+}95]. Location-independent [BDG^{+}95]. logic [AB96, DR96, MMS96, Nes90, Ter90]. logic-based [MMS96, Ter90]. logic-model [AB96]. logical [CH95, Tur93]. Lois [Lec96a]. Loomis [Pou95]. Lorenz [Kos95]. Lorge [Eic99b]. lossiness [CFI94]. LOTOS [Bol90, LMD90]. Low [GI90, JBVW94]. Low-cost [GI90, JBVW94]. LUSTRE [RHR91].

M [CM98, O’L97a, Ola98]. M.E.R.O.DE. [DS94b]. machine [AC95, RF90]. machines [AY98, Che91a, Che91b, Che93b, HW97, MD93]. Magee [Tra95c]. maintainability [GA96]. maintenance [Ede93, HLO^{+}92, Jør95, Mam94, PPRC99, RBDL97, Rin91b]. make [McL93, Sug95]. Making [BN92, De 90, FP98, Hua98b, MDE97, PD91, Tak99]. manage [Bal99b]. Management [Kel96c, Law98c, Apt94, AA98, Hef93, HHJ^{+}99, HM97, Joe97, KW95, KGL98, KRS98, LW99, ML90, Mar90, MR90, QJD90, SNS99, SJ96, SS90, Spr91, TSK90, TC93, Ver96, Zel95b, ZS99, vdHHHW97, Som96, Tra98a]. Manager [Per97, BR95, Mac91b, Mar91]. Managing [Gru90, SHO90, Lei96b]. manipulation [Dol91, NLS90]. Manny [CM98]. manual [GN92]. manufacturing [LP97]. many [Gre97]. MAP [Kel96b]. mapping [GC96]. mappings [Cas99]. Mark [Kos95]. market [BA94]. marriage [Par97]. Marvel [Bar92]. Mary [Pou95]. match [SC98]. matching [JC95, ZW93, ZW95]. mathematical [HS96b]. mathematics [Pla90]. matters [Kel95]. Maturity [Aye95, Bru96, Van95, BA94]. Maxick [Kos98]. may [NA98]. maybe [HB94]. me [Par98b]. meaning [Gri93]. meaning-preserving [Gri93]. means [Ber93, RHR91, SMG99]. measure [GTW93, MV91, MM96, SB98a, SB98b]. measurement [AK99, IGE97, Lot93, Roc94, Spr91, O’L96c]. measurement-based [Spr91]. measures [AS94, Bra90]. Measuring [Kok99, VPMM95]. mechanism [MiS97, RDO98, Shi91, pP98]. Mechanisms [BN93, SvK98]. medical [Bly97]. meeting [LWM99, Vot93]. megaprogramming [Tra91]. Meir [CM98]. Melding [PW95]. Mellor [Gos94, Lan93]. MELMAC [Gru90]. Memory [JPB97]. Mental [Edw96, Rya96]. menu [TWC92]. merging [LvO92]. messaging [BBP99]. meta [BdLLvdS90]. meta-modelling [BdLLvdS90]. metadata [DK99]. metadata-based [DK99]. Metamodeling [Kok93]. metamodels [PP97]. Metaphorical [Her99]. method [Bow90, BdLLvdS90, DS94b, Fie99, Gos94, GMS98, GBF99, Hua98a, Jia92, Kit96a, Kit96b, KGL98, Loy93, MS96, PT97, PD95, RGR92, Sch92b, Ste93, XJG98]. methodologies [Cla90, Con92, Ove90, PP97, Son95]. methodology [Ari93, BCC^{+}99, HSE93, Hol90, Jan94, JZZ^{+}95, KP98c, KSZ91, Mat97, MV96, Xia98, YH96]. methods [CLL96, DBDS93, FW91, Gel94, GPS96, Gom95, Jac96, Kil94a, Kit96c, Kit96a, Kit96b, KJ97a, KJ97b, Kit97, KJ97c, KP98c, KP98a, KP98b, Kit98, Lev92, Loy90, OM91, Pla90, Red90, Ros96, SK96, Sch92a, SC93a, Woo90]. Metric [LT93, Cla90, Kok99]. Metric-driven [LT93]. Metrics [FD96, Agg97, CS95, HS96a, HS96b, Kel97b, KBZP98, Roc94, Kos95, Doe97b]. MHP [NAC99]. Micro [Edw96]. Micro-Architecture [Edw96]. MIDAS [CCCL90]. middle [RP98]. middle-out [RP98]. middleware [AA98]. military [CCR90]. Mills [Poo96]. missing [Kai93]. ML [Ive97]. MMSD’96 [Müh96]. mobile [FC98, PRM97]. mobility [MPR99, PRM97]. Model [AY98, ABB^{+}96, Aye95, DCH97, GHJ98, WVF95, AG91, AB96, BL97, BEKM97, Boe99, Bru96, BFG96, CL91, CABN98, CGK97a, CYZ98, DS94a, DS94b, Den96, DP98, Fek95, Fie99, Fin92b, GH99, HK92, HJ99, IGE97, Iss91, JR94, JXXG98, KW95, Kol98, MPR99, MN95, MI91, OH96, Pad99, Pet91, Pil96, PD91, Rac95a, SAMS97, Tat95, TRC93, Tra91, Ver96, Wal92, WGHS99, XJG98, Zel95b]. model-based [Den96, Fek95, Fie99, Pet91]. model-checking [JXXG98]. model-driven [DS94b]. Modeling [Aye96b, Jar97, Sim96, AGI98, CGHS99, FH95, Gom95, JSW99, Kil91, Kro98, LXK98, LZL98, Mat97, Sim95, SC98, XDB99, ZW96]. modelling [BdLLvdS90, EHTE97, Sar92c]. Models [Doe98a, Edw96, FD96, Whi95, BG90, Car93, Cor96, Cor98, GC96, Hoa93, LU95, MHF^{+}97, MTO^{+}92, MNS95, Ves99, Zel94, Zel95a]. modern [Con92]. modification [HLS^{+}99, SRLZ98]. modification-side-effects [SRLZ98]. modifications [MFDM97]. modified [AS94]. Modula [Jor90]. Modula-3 [Jor90]. modular [CDHW93, FC98, HS95, NLS90]. modularity [FM95, VG94]. Module [SB98a, SB98b, mWE91, TTC95]. modules [WP92, Woi93]. Moksha [RM99a]. monitor [GHJ98, WH99]. monitoring [Zho90]. Montana [Kar98]. Morgan [Ray97b]. mortem [Gar97]. most [HKMR98, Par95]. Mostly [MS97a]. motion [Shu93]. Move [Lei96b]. Movement [Van95]. MUDs [CB99]. Multi [Nou96, ADLK97, BSKH92, CYZ98, IWJ90, PY96, RM99a, XGG98]. Multi-agent [Nou96, XGG98]. multi-formalism [IWJ90, PY96]. multi-site [ADLK97]. multi-threaded [CYZ98]. multi-user [BSKH92, RM99a]. Multibox [Dya94]. multicasting [CM96]. Multilingual [Hug91]. Multimedia [Müh96, BEKM97]. Multiple [PR99, SFE97, FLP99, Jar97, MR92]. multiple-view [MR92]. multipoint [Tak99]. multiprocessor [KSL90]. MUST [Per97]. mutant [UOH93]. Mutation [UOH93, DR90, WF93]. mutations [DTF96]. my [Zel99].

nail [MAKGM97]. name [KW95]. names [HM97]. naming [BDG^{+}95]. narrow [Tak99]. narrow-band [Tak99]. NASA [WTW^{+}91]. National [Bar98]. NATURE [Mai95]. navigation [PK97]. NDHORM [ZW96]. Need [Edw96, HB94, Sha98, Vot93]. needed [CCR90]. needs [Hen95b, Joe97, Lam98, PP97]. net [CC96, Doe96a, Doe96b, Doe96c, Doe97e, Doe97f, Doe97g, Doe97h, Doe98b, Doe98c, Doe99a, Doe99b, How95a, How95b, How96]. net-based [CC96]. nets [FGP93, JSZ97]. network [JBVW94]. Neumann [Tra95b, Mac91a]. newspaper [Ano99]. next [Boe95, Gar95b, Hua98b, SO97, War92]. next-generation [SO97]. Nico [Not93]. no [HB94, Sha98]. notation [MS96, ZZW98]. note [Bow90, LXK98, MD93, Rin93]. Notes [CW98a, Con99, McL91, McL92, vR92, Doe96a, Doe96b, Doe96c, Doe97e, Doe97f, Doe97g, Doe97h, Doe98b, Doe98c, Doe99a, Doe99b, How95a, How95b, How96]. notification [RW97]. notions [Not97]. Novel [Tra98a, Not97]. NSF [Ang99, BBB^{+}99, SGJ^{+}97]. number [MI91].

O [HSE93, Bur96, HSE93]. O-O-O [HSE93]. Object [Aye96b, Aye96c, Cas99, CHS90, DH98, Doe97a, Kos98, Law98b, Lei96b, O’L97a, Per96, Per97, Pou98, Sug95, AL95, BK95a, BK95b, BL97, CYZ98, Che93a, Che91b, Che91c, CS95, DS94b, GC96, Hec90, HSF92, HSE93, Hua98a, Lan93, Law98a, Lea94, LW93, Loy90, Mac91b, MS96, MORT96, MSB97, Mrd90, Nou96, Nou98, Rin91a, Rin92, Rin96, SH98, SPL96, SC93a, TSK90, TC93, TvKP90, Tom94, Vaz93, Vaz96, Wal92, WKC^{+}97, WKP^{+}98, WW98, Wyb90, XDH^{+}98, XJG98, XDB99, Doe97c, Doe97d, Doe97b, Doe98a, Kos95, O’L97b, Pou95, Ray97a, Ray98a, Sur97]. Object-Based [Kos98, BL97]. Object-Orientation [Sur97]. Object-Oriented [Aye96b, Aye96c, DH98, Doe97a, Kos98, Law98b, Pou98, CHS90, AL95, BK95b, CYZ98, Che93a, Che91b, Che91c, DS94b, GC96, Hec90, HSE93, Hua98a, Lan93, Law98a, Lea94, LW93, Loy90, Mac91b, MS96, MORT96, Mrd90, Rin91a, Rin92, Rin96, SH98, SC93a, Tom94, Wal92, WKC^{+}97, WKP^{+}98, WW98, Wyb90, XDH^{+}98, XJG98, XDB99, Doe97c, Doe97d, Doe97b, Kos95]. Objects [Ray98b, ACLN95, AA95, AA98, BMS93, Che93b, GJJ^{+}95, Sha95b, Aye97]. observation [RW97]. obstacles [VPC98]. obTIOS [Que92]. ODM [Sim95]. off [Gam97, MOT97, Tra96a, Was95]. off-the-shelf [Gam97, MOT97]. Oikos [ACM90]. On-line [SvK98]. one [Kim98, WWBJ98]. one-day [Kim98]. ongoing [CB99]. online [PSTZ99]. only [Bec99]. OO [Gri96, HS96d, ZW96]. OOA [Sch92b]. OOD [JS90]. OODPM [Dro96]. OOPSLA’94 [And95]. Open [Lec96a, MLMK97, Bla97, Kar98]. Operating [LA98]. Operation [LvO92]. Operation-based [LvO92]. operational [AB97, AKSR93, Woi93]. OPIUM [DE91a, DE91b]. opportunity [Ghe99]. Optimization [LGRM91]. optimized [JGS99]. optimizing [CD90]. option [MS97a]. Oracle8 [Pou98]. Oracles [DY94, DR96]. order [GPS96, NCC99]. Ordered [DR90]. ordinator [LW99]. organisational [LW99]. Organization [Ben96, Sim95, Gir99]. organizational [And95, Rem97, SC98]. organizations [Buc91, Ver96]. orientation [BK95a, Sur97]. Oriented [Aye96b, Aye96c, DH98, Doe97a, Kos98, Law98b, O’L97b, Pou98, Ray97a, Ray98a, AL95, BK95b, BM98, CHS90, CYZ98, Che93a, Che91b, Che91c, CS95, DS94b, EHTE97, GC96, Hec90, HSE93, Hua98a, Kok96, Lan93, Law98a, Lea94, LW93, LZL98, Loy90, Mac91b, Mai95, MS96, MORT96, Mrd90, RRG97, Ram96, Rin91a, Rin92, Rin96, SH98, SC93a, TvKP90, Tom94, Vaz93, Vaz96, Wal92, WKC^{+}97, WKP^{+}98, WW98, Wyb90, XDH^{+}98, XJG98, XDB99, Doe97c, Doe97d, Doe97b, Kos95]. OSI [BCJ96]. outline [TCY93]. outlines [Bal99b]. Overcoming [VPC98]. Overview [THRE^{+}94, Law97, Rin91b, Sit96, Ter90, Wah99].

P [Ray98a]. package [Car93]. packaging [HK97, Sha95a]. pages [Kel96c]. pairs [NA98]. Pan [BGV90]. panel [Boe96, DKLS96, Gar95b, GJJ^{+}95, HKMR98, Kic96, Lyu91]. paper [Bec99, CE99, RM95, Tri92]. papers [HKMR98, Kil99, Par95]. paradigm [Bal99b, KSZ91, Sha95b]. paradigmatic [AL95]. paradigms [SH98]. parallel [Iss91, JG96, NA98, Nes90, YSP98]. Parameterized [Che92b]. Parnas [Eic99b]. parsers [Dya94]. Part [Edw94, OSWZ94, Kit96c, Kit96a, Kit96b, KJ97a, KJ97b, Kit97, KJ97c, KP98c, KP98a, KP98b, Kit98, Pfl95a, Pfl95b, SK96, BHKW94, EHL^{+}94, HSWZ94, Pfl95d]. partial [DP98, GPS96, NCC99, SPH99]. partial-order [GPS96]. participation [Bru96]. Partition [PY93, Nta98]. Partnering [MK98]. Passing [Poo96]. patents [Aha99]. path [FB97b, MF96, YSP98]. path-wise [MF96]. PATHOS [AL95]. paths [BGS97]. Pattern [Bur96, Kim97, SvG98, RRG97, SRZ97]. Pattern-based [SvG98, SRZ97]. Patterns [Doe98a, FLM95, DvL96, HK97, Joh97, MSU98, MDE97, RDO98, SP98, Vil95]. Paulson [Ive97]. PCTE [Lec96a]. pedagogical [Tra95e]. people [CS98]. perceptions [Edw99]. Perestroika [Buc91]. Performance [Mus91a, Pou98, LSJ97, NCC99]. Performing [HR94]. personal [Bar98, Hum95]. perspective [Ell98, LWM99, Mac91b, Sar92b]. Perspectives [SFE97, Wyb91]. pertains [Rin91b]. perturbation [MM96]. Peter [Doe98a, Fin99, Tra95b]. Petri [CC96]. Phased [KM91]. phenomena [AD94]. Philippe [Aye97]. Physical [CR99a]. picking [Fug96]. pictures [Che91b, Che91c]. piece [WWBJ98]. Pirbhai [Tat95]. pit [Boe99]. Planning [Dro96, GCBM96, Kit97, Sar91, Sar93]. plans [AK99, Sef94]. play [BBK^{+}97]. plea [Kel95]. PLEIADES [TC93]. PLoP [Vil95]. PLoP-94 [Vil95]. plug [BBK^{+}97]. plug-and-play [BBK^{+}97]. PLUNGE [Ros97]. PM [Zho90]. pointer [MF96, ZRL96]. pointers [AG98, LH99b]. points [LH99a]. points-to [LH99a]. poised [Bag98b]. policies [GI90]. policy [AA98, Tra97c]. polite [BO99]. polylingual [BKW96]. Polymorphism [PB94]. position [RM95]. positive [Got99]. possible [Tom94]. post [Gar97]. power [Gel94]. powerful [De 90]. Practical [Pol90, Tat95, Buk91, CA93, Kit96b, Sar92b, Woo93, ZRL96]. practice [Bas97, DKLS96, Dro97a, Got99, MFCS95, Ros96, Tra98b]. practices [Rem97]. Precise [RR95, CGHS99]. predicates [BGL98]. predict [HS96a]. Predicting [Ham96, RW96]. predictions [Ebe97]. preference [Tri93]. Preliminary [Mar98b, Dai95, Rol94]. Prentice [Doe97a, Ire97]. prescriptive [Cla90]. presence [AG98, HS96d, LH99b]. presentation [GHM92, Mus91a]. preserving [Gri93, YHR90]. Press [Ive97, Tra95b]. primer [Rac96, Per96]. principal [FB97b]. Principles [FKV95, Rin91a, Roc94]. priorities [Boe96]. priority [GMK99]. probabilistic [Pad99]. Problem [Apt94, Fav96, Kel96a, RP98, RBDL97, SRLZ98]. Problems [Jac96, Bra90, DJ98, LW96a]. procedural [FR99]. procedure [LS90b, Zho90]. procedures [Kor96, VG94]. Process [Dav95b, Doe97b, Lam97b, Lot93, Mon97, SGJ^{+}97, AS95, ACM90, AC95, BN93, Bar92, Boe95, DS94a, Ell98, EAD99, Fin92b, Fis93, FP94, GJ94, Gos94, Han98, Hef93, Hei92, Hen95b, HGGM98, Hum95, JSW99, Joe97, Kir91, LF95, LHR99, PW97, PPRC99, Ram96, Rem97, Sco92, Sha95b, Sta93, SC98, SHO90, SO97, Ter93b, TCY93, Ves98, WW98, WBG^{+}99, Kri95, Aye95, Lei96a]. process-centered [Bar92, GJ94, SHO90]. process-integrated [PW97]. process-oriented [Ram96]. Processes [Bai96, Fra96, BR95, Bol90, BM99, CFF93, FF97, Gru90, JDL^{+}99, Kok96, KGL98, Sch92a, WGHS99]. ProcessWall [Hei92]. ProDAG [ROMA92]. produce [HB94]. Product [Add98, CW98a, Dav95b, BGW^{+}99, Gri99a, Tra98b, Was95]. Productive [GA96]. productivity [CP93]. profession [Got98, Got99, Gri98, Sha98]. professional [Bag98b, Buc91, Got99]. profile [AKSR93, Coo98a, Coo98b, CM98, Eic99b, Eic99a, Fin99, Lec99, Mes98]. profiles [Woi93]. profiling [RBDL97]. Program [BBB^{+}99, BB96, Bur96, Kim97, Law98b, Lei95, MRS98, SR96, ZRL96, Ang99, AG98, Egg90, Gri99c, GN92, Gri93, Ham98, HMR93, JR94, LSJ97, LH99a, LN99, Mar98a, McL91, MFDM97, OH96, RBDL97, RRL99, Shi91, SH96, Sol98, Spr91, Tay97, THRE^{+}94, TCM98, Tra95d, YHR90, You90, Zel99, vR92]. Programmer [Ive97]. programmers [Bou91]. Programming [Bri95, Fin97, RHR91, mWE91, Ana98, BMS93, Bec99, CYZ98, Che93a, Clé90, CE99, Fav96, Hei92, Iss91, Jor90, Kar98, Kic96, KNN99, Lai91, Mod92, Rin96, SH98, Ste99, WKC^{+}97, WKP^{+}98, WW98, Wil90, Ray97b]. programs [Bal99b, BJ93, Blu93, CC96, CGHS99, Cor98, DC94, FR99, Fro90, GPS96, HS96d, Kor96, MV91, NAC99, Red90, RM99b, Tom94, YSP98, Zho90]. progress [GMK99, PA96, Rac97]. Project [DHvL^{+}95, FCC94, Tra98a, ADLK97, BF92, Cas99, Hen95a, Kad92, KRS98, Mar91, Pil96, QJD90]. projects [Pad99, PV97, Pol90, Kel96e]. Prolog [BJ93, DE91a, DE91b, KNN99]. Promise [O’L96a]. Proof [FC98, CH95]. propagation [Gor93]. properties [CGK97b, DY94, DC94, ND90, WP92]. Property [FB97a]. Property-based [FB97a]. Proposed [Bri95, Han98, MV96, Rin91a]. proposition [Kol98]. PROTÉGÉ [GAM95]. PROTÉGÉ-II [GAM95]. Proteus [GHM92]. protocol [BCJ96, CM96]. protocols [NCO96]. Prototype [KKLK98, BG90]. prototypes [OM95]. Prototyping [Sur97, BL97, CC90, THM99, Zho90, Aye97]. Prover [WG90]. provers [TBK92]. Providing [AGM97]. PSP [DJ98]. PTR [Doe97a]. Public [Neu95e, Neu95f, CDP95, Neu90a, Neu90b, Neu90c, Neu91a, Neu91b, Neu92, Neu93, Neu94, Neu95d, Neu95g, Neu96a, Neu96b, Neu96c, Neu96d, Neu97a, Neu97b, Neu97c, Neu97d, Neu98a, Neu98b, Neu98c, Neu98d, Neu99a, Neu99b]. publishing [LP97]. purpose [Egg90, Vog93]. push [HJ99]. puzzle [WWBJ98].

Q [BR94]. Q-Sim [BR94]. Quality

[Bra90, Car96, FD96, HK97, Kel96d, O’L96b, PSP95, Smi91, AK99, Bro90, Buk91, DJ98, HS96a, Hef93, HB94, Jør95, Ket90, Kok96, KRS98, Kro98, Lai92, Lok92, Lot90, Sef94, SD94, VW99, PP96]. quantification [BF91]. Quantitative [Kel96c, BZ95, KP98c, KP98a, KP98b]. Query [Bal99b, TTB^{+}90]. querying [MSB97]. questionnaire [Jør95, KT94]. questions [Tra94]. queueing [BR94].

R [Che91a, Yu91]. R&D [Zuc94]. RAISE [Dan91]. random [Nta98]. rapid [OM95, Ray96]. rat [Mus91a]. ratio [Mus91b]. Rationale [Mus91b]. Re [DH98, Sar94, Ada99, dPL96, Yu91, Zuc94]. Re-Engineer [Sar94]. Re-engineering [DH98, dPL96, Yu91, Zuc94]. re-inspection [Ada99]. Reachability [PA96, CGK97a, CK93, CK95, CGK97b]. reaction [McL92]. reactive [Che91a]. read [War96]. readability [Han98]. reader [Kam91]. ready [Sha98]. Real [Kir91, TBK92, BF99, CL91, CC90, CA93, Cor96, DTF96, Gaf99, HRR91, JXXG98, KL96, KS91, XP91]. Real-time [Kir91, BF99, CL91, CC90, CA93, Cor96, Gaf99, HRR91, JXXG98, KL96, KS91]. reality [Pet96, Zel94, Zel95a]. really [HB94]. realtime [CPK93]. Reasoning [FXHZ98, GJND98, JSZ97]. Reconciling [SN90]. reconfiguration [WF99]. recovery [SvG98]. reduce [VG94]. Reduction [HW97, DBDS93]. Reengineering [BTS94, ST98, BA94, LT93, MRS98, PR95, SP98, Tom94, Bor95]. Reference [FCC94, BCGS95, WGHS99]. refinement [DvL96, Woo93]. refinements [FGP93]. Refining [BGS97, GS95]. reflections [Fug99]. reflexion [MNS95]. REFSQ [PSP95]. REFSQ’97 [DOP98]. REFSQ’98 [OP98]. regression [Bal98, KAY98, RW96, Wah99]. Related [Neu95e, Neu95f, Tra95b, Lut93, Neu90b, Neu90c, Neu91a, Neu91b, Neu92, Neu93, Neu94, Neu95d, Neu95g, Neu96a, Neu96b, Neu96c, Neu96d, Neu97a, Neu97b, Neu97c, Neu97d, Neu98a, Neu98b, Neu98c, Neu98d, Neu99a, Neu99b]. Relational [Car96, DJJ96, JSZ97, Wil90]. relations [SPL96]. relationship [DS94b, Kil91]. Relationships [Kri95]. relaxation [GMS98]. release [vdHHHW97]. relevant [GBF99]. reliability [AS95, BF91, HV93, Lyu91, MZ98, Wey95]. reliable [CM96, MS97b, Aye96c]. reluctance [Shu93]. remembered [Not93]. remote [Zho90]. removing [De 90]. renovation [vdBKV97]. Replace [Sej95]. replicated [SH98, WRBM97]. reply [KL90, McL92]. Report [Add98, DH98, Ede93, Fin98, Kim97, Kri95, Kwo97, Kwo98, Mon97, Rei97, SGJ^{+}97, Tri91, And95, BB96, CQW^{+}95, FCC94, Kim98, KSL90, Lam97a, LSJ97, LN99, PA96, SZ96, SFE97, Tat96, Tho98, Tra92, Wie92]. reports [EW97, Gir99, GT93, LW96b, PT94]. repositories [BDG^{+}95, Lec96a]. repository [Lin92, Mac91b].

Representation [Sim96, Her99].

representations [Bol90, Gri93, SPH99]. representative [MZ98]. Representing [MSB97]. Requirements [Fin92a, PSP95, PP96, Ram96, Amo95, AG91, AG93, AB96, CD94, CABN98, DvL96, FF97, GH99, HRR91, Hol90, JH98, JBVW94, Kai93, Kro98, LP97, Lut93, Mai95, MK95, OM95, PF95, Sha94, Tat95, VPC98, ZW96]. Requirements-driven [Ram96]. Research [BBB^{+}99, BK99, GJK95, Kri95, XGG98, Boe96, DE91a, DE91b, Edw94, FF97, Fug99, HKMR98, JG96, KSZ91, Lyu91, OSWZ94, Par98a, Par98b, Sta93, Xia98, ZAD^{+}97]. resilience [SS90]. resolutions [LW96a]. RESOLVE [BHKW94, EHL^{+}94, Edw94, HSWZ94, OSWZ94, SW94]. resources [Sea97]. response [AA95, Che91b, Che93b, GMR99, Mar98b, WTW^{+}91]. responses [Day96]. responsibility [SC98]. restructuring [GN92, Gri93, GCBM96]. results [dPL96]. resumption [Shi91]. retrieval [Dro99, MAKGM97, PW95]. reusability [Aik90, Ara95, Che93a, Che94a, GGL^{+}97, HC99, Lam98, PP97, VPMM95, Tra92]. reusability-based [Che94a]. reusable [ACLN95, CDP95, CCR90, Dai95, Dav95a, DR97, FKV95, Gar90, Hen95b, Lam97a, MSB97, PW95, Tra97b]. Reuse [Ben96, BZ95, Boe95, BM97, DvK95, Fra96, GAM95, Gri96, Lei95, Mai95, MOT97, PF95, Pou97, PBL99, Tra95f, ZAD^{+}97, BG90, Bas97, BK95a, Bha95, BK95b, CH95, CD94, Che92b, Edw99, FM95, Gac95, GJK95, GMR99, GJJ^{+}95, HR96, Hen95a, Jar95, Jar97, JC95, KRT97, Kog95, LW96a, Lam97b, Lam98, LW93, LU95, MFCS95, MK98, Mar98a, Mar98b, MS97a, Mug97, Par99a, Pet91, Pou93, PD98, PD91, PD95, RM95, Rin91b, Sha95a, SR96, Sit96, Sit97, SVR97, Tra90, WKP^{+}98, Was95, WT91, Whi95, XZLY95, Yu91, ZW93, EW97, GT93, LW96b, PT94]. Reuse-oriented [Mai95]. reuser [Lam98]. Reusing [Gol90]. Reveng [DKW92]. Reverse [KT94, vdBKV97, CQW^{+}95, DKW92, JR94, JSZ97, MTO^{+}92, OM95, Yu91]. reverse-engineering [DKW92, Yu91]. Review [Aye96a, Aye96b, Aye96c, Bor95, Doe97b, Fin96a, Fin96b, Hol96, Ive97, Kel96e, Kel96d, Kel96b, Kel96c, Kel97a, Lec96a, Lei96a, Lei96b, O’L96a, O’L96b, O’L96c, Per96, Ray96, Ray97a, Sur97, Tra97a, Doe97a, Ire97, Kos95, Neu95c, Tra95b, Tra95c]. Reviews [Aye95, Aye97, Doe97c, Doe97d, Doe98a, Kos98, Lei98, O’L97a, O’L97b, Ola98, Per97, Pou95, Pou98, Ray97b, Ray98a, Ray98b, Tra98a, LSJ97]. revision [MAM93]. revisited [PM96]. Revolutionary [Sej95]. ribosomes [GAM95]. right [ZAD^{+}97]. Rigorous [Dan91]. risk [KGL98]. Riskit [KGL98]. Risks [Neu90a, Neu90b, Neu90c, Neu91a, Neu91b, Neu93, Neu95d, Neu95e, Neu95f, Neu95g, Neu96b, Neu96c, Neu96d, Neu97a, Neu97b, Neu97c, Neu97d, Neu98a, Neu98b, Neu98c, Neu98d, Neu99a, Neu99b, Neu92, Neu94, Neu96a, Tra95b]. RIT [LN99]. Robert [Neu95c]. Robots [Gaf99]. Robust [CS97]. Role [RI99, Gar90, Kra97, Spr91]. Roles [PPRC99]. ROSATEA [RI99]. Rosetta [VB97]. Ross [Aye96b]. RPDE3 [OH90]. rules [Che94b, Jur98, Lan93]. Run [Che92a, Kel96e]. Run-Time [Che92a]. running [KP98a]. runtime [VG94]. Ryder [Eic99a].

S [Ola98, Pou95, Che91a]. S-R [Che91a]. Safety [Neu95b, WH99, Che92a, Lut93]. safety-related [Lut93]. Safeware [Neu95b]. Salesman [Lei95, Tra95d]. sampling [MV91, PY93]. satisfying [Hen95b, XP91]. save [Loy93]. SaveMe [BBP99]. scalability [Fav96]. Scalable [BSST93, BTS94, CA94]. Scale [KLS99, CLL96, PR95, PSTV95, RW97]. scenario [Hol90]. scenario-based [Hol90]. scenarios [DFKM97]. schedule [Pil96]. scheduling [BF99, Sar91]. schemata [UOH93]. Schlaer [Lan93]. Schloß [THP93]. Science [Bar98, Con99, AS94]. scientific [PD95]. scoring [KJ97b]. SCR [AB96]. SDE [Bar92]. seamless [BKW96]. Second [CQW^{+}95, Ray97b, WBG^{+}99, CW98a, PP96, Wol97]. Security [Kem97, Rei97]. SEE [TL90]. seem [BO99]. SEEs [Lot93]. SEI [Kra97, TTC95]. SEKE’97 [Kwo97, Kwo98]. Selected [GJ94]. selecting [Kit96a, Kit96b, RGR92]. selection [Bal98, FB97b, Tri90]. semantic [OH96]. Semantics [Kil99, AB96, CJPV90, XZ95, YHR90]. semantics-preserving [YHR90]. semi [Mar98a]. semi-automatic [Mar98a]. seminar [Kim98]. sensitive [WWBJ98]. sensitivity [SRLZ98]. Separate [HR96]. sequence [HS96c]. sequential [DFKM97]. serial [WF93]. seriously [Mai97]. server [Hei92]. service [BCJ96, Fek95, Que92, TTB^{+}90]. services [CL91]. session [BCJ96, Not97]. set [Pfl95d, Rin91a]. sets [MM96]. seven [Bou93]. Seventh [FvL94]. sharing [MS97a, VW99]. shelf [Gam97, MOT97]. Shlaer [Gos94]. short [Rin91b]. should [Che92c, Hum95]. shuttle [Mod93]. side [HS96d, SRLZ98]. side-effects [HS96d]. SIGCSE [Con99]. Signature [ZW93]. signatures [Pre99b]. signs [Bou93]. SIGSOFT [SFE97, Tra97c, Wie92]. SIGSOFT’96 [Gar97]. Sim [BR94]. simple [Pre99a, Ste93]. simulated [TCM98]. simulation [Aik90, SRZ97, VPMM95]. simulator [BR94]. simulators [JPB97]. site [ADLK97]. Sixth [Som96]. skills [Del90]. sleeve [HV93]. slices [GS95]. Slicing [Law98b, DPS96, FB97b, GS95, GBF99, HW97, Law98a, RHSR94, SH96]. Slow [War96]. small [PR95, Rya96]. Smalltalk [Ray96, Ray98b]. smart [GLV99, MK95]. Smooth [Lei96b]. social [Gri99a]. Soft [KSZ91]. Softool [Mar90]. Software [Add98, AS95, ACM90, Ara95, Aye95, Aye96c, BBB^{+}99, Ben96, Ber96, Bou93, BK99, Car96, Che91b, CW98a, Con99, Con92, Coo96, DTF96, DW97, Doe97a, Dov97, Edw96, EW97, FCC94, FvL94, FD95, FD96, Gar95a, Gar95b, GTP95, GM96, Got98, GT93, Gri98, GTW93, Han98, Han95, HGGM98, Hol95, Hol99, KLS99, Kel96b, Kel96c, Kos95, Kos98, KW96, Kri95, LW96b, Le 96, Lei95, LW93, Lok92, Lud96, Mar97, Mon97, Müh96, MNS95, Nou98, O’L96b, O’L96c, O’N97, Ola98, Par97, Per97, Pfl95c, PSP95, PP96, PT94, Pou97, PBL99, Ray96, Rec96, Rem97, RI99, Roc94, Sar92b, Som96, SFE97, Sur97, TTC95, TL90, Tho98, THP93, Tra92, Tra95e, Tra95a, Tra95f, Tra96a, VPMM95, Wol97, vdHHHW97, AK99, AAG93, AL95, AS94, Agg97, Aha99]. software [AC95, And95, ABB^{+}96, Ang99, Ari93, AB96, AL93, Bag98a, Bag98b, Bar98, BKW96, BSST93, Bec99, BSKH92, Bha95, BZ95, Boe96, Boe99, BA94, BBK^{+}97, BP92, BDG^{+}95, Bru96, BW93, Bry99, Buk91, BF91, CH95, Car93, Cas99, CS97, Che92b, Che93a, Che94a, Che94b, CS95, CFF93, CCR90, Cor96, DS94a, DPS96, Del90, DBM97, DK99, Dow97, Ebe97, Ede93, Ell98, ELB^{+}98, FP98, Fin92b, FLP99, FP94, Fug99, Gac95, GJK95, GJ94, GS96, GC96, Gol90, Gos94, Got99, GJJ^{+}95, Gru90, GGL^{+}97, Gun96, GRW90, HLO^{+}92, HV93, Ham94, HK92, HS90, Hef93, HS96b, Hen95b, HB94, Her99, HS95, HZG99, HK97, Hug91, Hum95, IGE97, JD96, JSW99, Jar97, Jaz95, JS90, JG96, JC95, JPB97, JZZ^{+}95, Joh96a, Jør95, Kad92, KTC^{+}92, Kem97, Kic96]. software [KPP^{+}99, Kit96c, Kit96a, Kit96b, KJ97a, KJ97b, Kit97, KP98c, KP98a, KP98b, Kit98, Kog95, KT96, KRS98, KF98, Lai92, Lai96, LSJ97, Law98c, dPL96, LZL99, Lin92, Lut93, LN99, Lyu91, MK96, MAM93, Mah98a, Mai97, MFCS95, Mam94, Mar94, Mar98a, MiS97, MFR94, McL92, MR92, Mil93, Mod93, MQ94, Mug97, MTO^{+}92, MRS98, MI91, NACO97, NLS90, Nor97, OM91, Pad99, PO96, Par98a, Par98b, PW92, Pet91, Pfl94, Pfl95d, Pfl95a, Pfl95b, PR95, PT97, PSTV95, Pou93, PD98, PD91, Rac96, Rac97, Rac98, RRG97, Ram96, RM95, RBDL97, Rin91b, Rin92, Rin93, RGR92, RF90, SK96, SS95, Sca91, Sch92a, SSP95, Sco92, SvG98, Sef94, Sha95a, Sha95b, Sha98, SJ96, SR96, Sit96, Sit97, SW94, SS90, Sol98, Spr91, SVR97]. software [Sta93, Ste93, SvK98, Sug95, TC93, Tat96, TTB^{+}90, Ter93a, THRE^{+}94, Ter90, TFW93, TCY93, Tra94, Tra96b, Tri91, Tri92, Uhl97, Ves98, War92, WT91, WF99, WP92, WJ99, WVF95, Woo90, Woo93, WRBM97, Wyb90, Xia98, XJG98, XZLY95, YH96, ZW95, Zuc94, ZR96, pP98, Aye96a, Doe96a, Doe96b, Doe96c, Doe97e, Doe97f, Doe97g, Doe97h, Doe98b, Doe98c, Doe99a, Doe99b, Fin96a, Fin96b, Hol96, How95a, How95b, How96, Kel96d, Neu95c, Tra95c]. solution [SVR97]. solving [GBR98]. Some [Fug99, McL91, McL92, YH96, vR92, CS98]. soundness [Bol90]. source [MAH^{+}98b, MN95, MNS95]. source-to-source [MAH^{+}98b]. Space [FCC94, DBDS93, KL96, Mod93, MV91, PY96]. spatial [MTO^{+}92]. SPC [Fin98]. special [Car93, Egg90, Not97]. special-purpose [Egg90]. specialized [VG94]. Specific [Tra95e, GCN92, LW96a, LZL99, TTC95, THRE^{+}94, TCY93, Tra94]. Specification [FvL94, JC95, KW96, PO96, THM99, ZW95, ACLN95, BCJ96, CRS96, CR99b, Den96, GTW93, IWJ90, Jac90, JZZ^{+}95, Kir91, Kok99, Kro98, LHR99, LS94, LMD90, Ole93, SC93b, Ter90, VG94, Woo90, ZZW98]. Specification-based [PO96, THM99, CRS96, CR99b, SC93b, VG94].

Specifications [Tra95c, AD94, Amo95, ABB^{+}96, CCCL90, CDHW93, Che92b, CPK93, DJJ96, DR96, EGHT94, Fek95, FGP93, GH99, JH98, Jia92, Kel96a, Mat97, MMS96, OM95, PF95, Rus97, TvKP90, Vaz96, Ray97b]. specified [CK95]. Specifying [HRR91, Woi93, EHL^{+}94]. Speeding [RHSR94]. spin [Was95]. spin-off [Was95]. spiral [BEKM97]. sponsored [Tra97c]. SQA [HB94]. SRI [SMDP96]. SSR [Jaz99]. SSR’95 [SZ96]. stall [Mus91a]. Stan [Tra95c]. Standard [Lec96a, AGI98, Buc91, Ede93, Mam94, Rin91a]. Standards [Tra95c, Tri93, Kil94a, RS97, Tri91, Tri92]. start [Tra90]. State [AG91, HS95, AY98, AL93, CABN98, CK95, DKLS96, DBDS93, Edw99, FGP93, Hei92, HW97, JH98, KL96, LXK98, MD93, MV91, PY96, Ros96]. State-based [AG91, CABN98]. state-of-the [Ros96]. state-of-the-art [DKLS96]. state-of-the-practice [DKLS96]. state-space [PY96]. statecharts [Wal92]. STATEMATE [TFW93]. statement [Fos91, Tra97c]. statements [NA98]. static [GS95, LT93, NACO97]. statically [Bok99]. statically-checked [Bok99]. statistical [TFW93]. status [AD94, FCC94, LN99]. Staugaard [Lei98]. STD [Ede93, Ove90]. step [Got99, ZRL96]. steps [Gar95b, LHR99]. Stepwise [KS91]. Stimulus [Che93b, Che91b]. Stimulus-response [Che93b, Che91b]. Storage [JBVW94]. straightforward [Ste93]. Strategies [Doe98a, Lei96b, Sim96, RW96]. strategy [Ber96, GC96, Rac95b]. stratified [PY93]. stream [TvKP90]. Structural [CRS96, CR99b, Rin96, HS96c, KT96, YP96]. structure [BTS94, FR99, Guo97, MK96]. Structured [Dol91, Mod93, TvKP90, Bow90, Jan94, Loy90, PW95, Vaz93]. structures [Jac96, Tur93]. Structuring [Lei98]. studies [FI98, Jør95, KP98b, Pfl94].

STUDIO [PT97]. study

[BZ95, DTF96, Fek95, Hen95b, KP98c, KP98a, Lam97b, MR92, O’N97, PW92, Ros97, SC93b, WVF95, WRBM97, XZLY95]. style [AAG93, Gaf99, GAO94, JD96, LZL99, MORT96, MOT97, TTC95]. styles [Den96, Le 96]. sub [Gor94]. sub-critical [Gor94]. Subjective [Car96]. sublation [XZ95]. subroutines [Pre99b]. Subsystem [Kos98]. Subsystems [Edw96]. Subtypes [Rus97]. SUCCEED [Per97]. Succeedings [FvL94, KW96, Wol97]. Success [Bor95, Sef94, Tra98b]. Successful [Kel96e, Par98a]. suite [LGRM91]. Summary [GTP95, Lud96, THP93, Tra95f, CQW^{+}95, DOP98, EW97, Gar95a, LM97, LW96b, OP98, PSP95, PP96, PT94, PD98, PBL99, SZ96, Smi96, SMDP96, Tat96, Tra96a]. SuperPascal [Han95]. Support [OH90, AK99, AGM97, AC95, BN93, BKW96, BG94, BF92, CR99b, EAD99, Gos94, GCBM96, HLS^{+}99, Jan94, Lai91, Lot93, LU95, MORT96, PP97, Pou93, SSP95, Ter93b, You90]. Supported [Kri95, Sch92b]. Supporting [Bar92, Gir99, Hen95b, KTC^{+}92, RM95, CGK97a, FM95, PK97, Rin92, WT91]. Surfing [Doe96a, Doe96b, Doe96c, Doe97e, Doe97f, Doe97g, Doe97h, Doe98b, Doe98c, Doe99a, Doe99b, How95a, How95b, How96]. surprising [Pre99a]. survey [Bru96, Bry99, FF97, OM91, Tri93]. SW [BG90, KJ97c]. switch [GHJ98]. symbolic [CABN98]. Symposium [Con99, GM96, Pou97, GS96]. synchronous [PO96, RHR91]. Synergy [JK99]. synopsis [OSWZ94]. System [Che91c, Neu95b, Nor97, Sej95, Ana98, Apt94, AG91, BGV90, BBP99, BK95b, Bly97, CABN98, CS90, Dai95, EHTE97, Gam97, Gor94, GHM92, HK92, Jac90, KSZ91, LZL99, Mar98b, MSU98, Mat97, Mrd90, ND90, PT97, Rya95, Swa96, TSK90, TC93, TTB^{+}90, VH93, Vaz93, Woo90, ZS99, Zho90, vdBKV97]. Systematic [GJJ^{+}95, Gri96, PD95, HS96d, Mug97, SARL91]. Systems [Doe97c, Doe97d, Fin97, Gri99a, Hol96, Neu95e, Neu95f, Nor97, Sar91, Sar92c, SP98, Tra97a, AD94, AGM97, BKW96, BL97, BDM90, BGL98, CCCL90, Che91a, CK95, Con92, CC90, CCR90, CA93, DY94, Dro96, Dro97a, Dro99, DP98, EAD99, Fie99, Gaf99, Gam97, Gar95b, GCBM96, HRR91, HJ99, HHJ^{+}99, JG96, JXXG98, Joh96a, KL96, Kir91, KSL90, KS91, Lai92, LHR99, LA98, LWM99, MV96, MR90, MMS96, MS97b, Neu90b, Neu90c, Neu91a, Neu91b, Neu92, Neu93, Neu94, Neu95d, Neu95g, Neu96a, Neu96b, Neu96c, Neu96d, Neu97a, Neu97b, Neu97c, Neu97d, Neu98a, Neu98b, Neu98c, Neu98d, Neu99a, Neu99b, NLS90, OAFG98, RHR91, Rec96, Rin93, RvH91, SNS99, SD94, SC98, Sug95, THM99, Ver96, WTW^{+}91, WVF95, XGG98, XP91, Gar95a, LNS98, SGJ^{+}97, UCD99].

T [HC99]. Tai [Off90]. take [Mai97]. Tale [Gri96, DeT91, PV97]. Taming [BDM90]. tar [Boe99]. target [Sca91]. Targeting [Lut93]. task [Lai91]. tasking [DBDS93]. taught [Par98b]. taxonomy [Bou91, BDM90, Lai91, LW96a]. TB [FGP93]. TCAS [Hei96]. Tcl [BR94]. Tcl/Tk [BR94]. teach [Mai97]. Teaching [Kel97a, JS90]. team [FXHZ98]. teams [FLM95]. Technical [Con99, Kit96a]. technique [BF92, CA93, Gor93, LH99b, RRG97]. Techniques [Lei98, Ber93, GBR98, HK92, MS97b, NCC99, PR99, WRBM97]. Technology [Bor95, Doe97a, Dov97, Lei96b, O’L96a, Dro99, Ell98, Fin98, Mil93, Neu92, Neu94, Neu96a, RS97, Sar92a, Wau97, Mon97, O’L97a, Per97]. Teknowledge [THRE^{+}94]. telecommunications [CL91]. telephone [GHJ98]. template [Lam97b, SC93b]. Temporal [ZS99, DY94, DR96, Nes90]. Terminology [TL90, Pet91]. terms [Pet91]. Test [Ber93, Raj98, SC93b, Tra96b, Bal98, BA94, Den96, DS97, FB97b, FI98, GBR98, GMS98, HF98, HS96a, Kob90, Kor96, KAY98, LGRM91, Mat92, Mil93, MAH^{+}98b, MMS96, MM96, OAFG98, Ram96, Vog93]. testability [VPMM95]. testable [WKC^{+}97, ZR96]. Testing [Kos98, UCD99, vM91, Agr99, AL93, BJ93, CRS96, CR99b, CLL96, DR90, FB97a, Fos91, FW91, HV93, Ham94, Ham96, Ham98, HR94, HS96d, KT96, MF96, MZ98, Nta98, OM91, PO96, PY93, Ros96, RW96, SC93b, TFW93, Wah99, Wey95, YP96, RI99, Kos98]. tests [GH99, WKP^{+}98]. Texas [Bag98a, Bag98b]. text [Dro99]. their [Blu93, KM91, LW96a, MB96]. them [Boe99]. theorem [TBK92]. theory [Bas97, BF99, Dro97a, Ham94, IGE97, KF98]. there [Sha98]. thing [Coo96]. Third [Tra95f, Nor97, Lud96]. threaded [CYZ98]. three [Yu91]. throughout [KPP^{+}99]. Time [Che92a, AGM97, Bec91, BF99, CL91, CC90, CA93, Cor96, Gaf99, Gre97, HRR91, JXXG98, KL96, Kir91, KS91, MMS96, WTW^{+}91, XP91]. time-critical [MMS96]. timing [AG93, XP91]. Tk [BR94]. Today [Zel99]. together [Gri98]. tolerance [Kel95, Fin96b]. Tolerating [BM99]. Tom [Tra98a]. too [War96]. Tool [GCBM96, AK99, Agg97, CD90, CC90, EGHT94, GI90, Gri93, Kir91, Kit96c, Kit96a, Kit96b, KJ97a, Ole93, PK97, Rin92, RGR92, SK96, Sch92b, Sha94, Sug95, Wyb91, YP96]. ToolCASE [Lin92]. Tools [Dav96, Fin97, BK95a, BA94, Day96, GRW90, HK92, Hua98b, KJ97b, Kit97, KJ97c, KP98c, KP98a, KP98b, Kit98, Law97, Lin92, LS90b, Mar91, PY96, PW97, Pol90, Ter93b, O’L96b]. total [Hef93, MI91]. Tour [DHvL^{+}95]. track [Nor97, ZAD^{+}97]. tracking [FXHZ98]. tradition [Rac98]. traditional [Con92, SH96]. Training [Con99, Del90]. transfer [Del90]. transformation [EHTE97, Guo97, RF90]. transformational [GCN92, Red90]. Transformations [BB96, YHR90]. transition [CL91, LS90b, Lei96b]. Transitioning [BGW^{+}99]. translators [MAH^{+}98b]. tree [MSU98]. trees [KBZP98]. Trip [Kim98, Wie92, Tra95c]. triplet [Mug97]. trustworthiness [MV91]. TTM15 [ADLK97]. turn [Tak99]. Tutorial [Kim97, YT90]. tutorials [Gri99b]. Two [Gri96, Ray97a, XJG98, Kok96, PV97, SC93a, WTW^{+}91]. Two-dimensional [XJG98]. type [BDM90, CR99a, Pre99b, SCB^{+}99]. typed [GLV99]. typedefed [BN92]. Types [Pfl95c, BN92, Bok99, BG94]. typing [MORT96]. typology [Kel97b].

ubiquity [RM99a]. Ulrich [Ola98]. ultra [MS97b]. ultra-reliable [MS97b]. UML [JSW99, Jur98]. Uncertainty [Lap90]. Unconstrained [MB96]. unconsummated [Par97]. understand [AAG93]. Understandability [RGR92, Lai96]. Understanding [Fis93, HJN94, WWBJ98, Kil94b, Kro98, Son95]. unified [BM99, SVR97, Zel95b]. unifying [KW95]. unit [BCJ96]. UNITY [PRM97]. University [Ive97, Mai97]. unreliable [CFI94]. update [Gri93, Tay97]. Urgent [Sta93]. Use [Gam97, Bok99, CB99, CLL96, Hum95, KPP^{+}99, KBZP98, MB96, Mar98b, Neu92, Neu94, Neu96a, RBDL97, SPH99]. Used [Lei95, Tra95d]. used-program [Tra95d]. Useful [mWE91]. User [Day96, BSKH92, BL97, CCCL90, CK95, DCH97, Fis93, RM99a, TBK92, VPC98]. user-analyst [VPC98]. user-interfaces [TBK92]. user-specified [CK95]. uses [MB96]. Using [AAG93, Dav96, DBDS93, Fin96a, Fin97, GH99, GPS96, Gos94, HS96a, HM97, JSW99, Lei98, MORT96, MM96, NCC99, PP97, SC98, VG94, Vaz93, Ves98, Wal92, Wey95, AS95, Agr99, AL93, BCJ96, BTS94, BBP99, BR94, BGS97, CGK97b, Dan91, Dro96, DCH97, EGHT94, GHJ98, GBR98, GS95, GMS98, HZG99, Hug91, Kok96, KGL98, Lam97b, LSJ97, MI91, NCO96, SPL96, SW94, ST98, SVR97, TTC95, Ter90, TvKP90, TCM98, UOH93, Woo93, Ire97].

V [Edw94, Mar91, Mod93]. Validating [LSJ97]. validation [CM96, GPS96, LGRM91, Nou96, OM95]. validity [HS96b]. value [O’N97]. variables [MF96]. variation [MM96]. VDM [HJN94, LS90a]. Verification [BF99, CGK97b, NCO96, BCJ96, BFG96, CM96, DS97, Egg90, FC98, Iss91, NCC99, Nou96, RvH91, Tat95]. Verified [You90]. Verifying [BGL98, DC94, HRR91, RHR91]. VeriSoft [GHJ98]. versa [Kic96]. version [ACM97, LW96a, Zel95b]. versus [Cla90, YH96]. vice [Kic96]. videoconferencing [Tak99]. view [Bar98, EHTE97, FLP99, MR92, Sar92a, Yu91]. view-oriented [EHTE97]. Viewpoint [Lam98]. Viewpoint-centred [Lam98]. Viewpoints [SFE97]. Virtual [CB99, LW99, AC95, BDG^{+}95, JBVW94, XDB99]. Visual [Che94b, Che91a, Che93b, MTO^{+}92, OAFG98, OM95, WW98]. vocabulary [MAKGM97]. votes [Bag98a]. vs [GN92, Kog95].

WACC’99 [Gri99b, Gri99c]. Wakeman [Lec96a]. warning [Bou93]. warp [Bec91]. Wasn’t [Sar94]. way [LM97]. weakness [SB98a, SB98b]. Web [Bla97, BM97, PW95]. weighted [AKSR93]. weird [CS98]. well [HB94]. west [PM96]. Where [Tra90, GJK95]. which [Rya95]. white [Mat92]. Who [Par98b]. whole [AG98, LH99a]. whole-program [AG98, LH99a]. Wide [BM97, PW95, BCC^{+}99, Bla97]. width [Tak99]. wild [Not97, PM96]. wild-west [PM96]. William [Ola98]. Williams [Per97]. Window [Ano99, Tra95a]. WinWin [BEKM97, BZ99]. Wirfs [MS96]. wise [MF96]. wish [Ole93, Ole93]. Wish-4 [Ole93]. WISR [PBL99]. WISR8 [EW97]. WISR’92 [GT93]. WISR’93 [PT94]. WISR’95 [LW96b]. within [Kob90, RM95]. without [GHJ98, Sca91]. won’t [Loy93]. Woodcock [Ire97]. WOOR’97 [DH98]. Work [Kri95, Blu93, CB99, MDE97, PD91, SMG99]. worked [Zel99]. Workflow [SGJ^{+}97, BCC^{+}99, HHJ^{+}99, LW99, SNS99, WGHS99, ZS99]. workflows [BM99, HLS^{+}99]. Working [Ive97, dPL96, Ray97a, Sea97, CCR90, CQW^{+}95, EW97, GT93, Gri98, Kob90, LW96b, PT94, Tra92]. workplace [CB99]. Workshop [Add98, BBB^{+}99, CW98a, EW97, FvL94, Gar95a, GTP95, GT93, KLS99, KSL90, KW96, Kri95, LW96b, LNS98, Lud96, Mon97, Müh96, OP98, PSP95, PP96, PT94, PBL99, RI99, SGJ^{+}97, Som96, SFE97, THP93, Tra92, UCD99, Wol97, Boe95, BA94, BK99, DOP98, Gri99c, Smi96, SMDP96, Tho98, Tra96a, Tra98b, And95, BB96, DH98, Kil99, Tho98]. workshops [Tra97c]. World [Ano99, BM97, PW95, Tra95a, Bla97]. World-Wide [Bla97]. writing [FKV95, Gir99]. wrong [Xia98]. WWAC [BZ99].

Year [Ola98, Pre99a, RBDL97, WWBJ98]. years [Rac97]. yes [HB94, Sha98]. Yesterday [Zel99]. Yourdon [Bow90, Doe97c, Doe97d].

Z [HJN94, Woo93, XDH^{+}98, Ire97].

References

Appelbe:1995:BOR

[AA95] Bill Appelbe and Gregory Abowd. Beyond objects: a response. ACM SIGSOFT Software Engineering Notes, 20(3): 45–48, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Sha95b].

Astley:1998:CCD

[AA98] Mark Astley and Gul A. Agha. Customization and composition of distributed objects: middleware abstractions for policy management. ACM SIGSOFT Software Engineering Notes, 23(6):1–9, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Abowd:1993:USU

[AAG93] Gregory Abowd, Robert Allen, and David Garlan. Using style to understand descriptions of software architecture. ACM SIGSOFT Software Engineering Notes, 18(5):9–20, December 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Atlee:1996:LMS

[AB96] Joanne M. Atlee and Michael A. Buckley. A logic-model semantics for SCR software requirements. ACM SIGSOFT Software Engineering Notes, 21 (3):280–292, May 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agarwal:1997:IIO

[AB97] Rakesh Agarwal and Giorgio Bruno. Implementing inheritance in operational graphical languages. ACM SIGSOFT Software Engineering Notes, 22 (4):50–55, July 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anderson:1996:MCL

[ABB^{+}96] Richard J. Anderson, Paul Beame, Steve Burns, William Chan, Francesmary Modugno, David Notkin, and Jon D. Reese. Model checking large software specifications. ACM SIGSOFT Software Engineering Notes, 21(6): 156–166, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ambriola:1995:DVM

[AC95] Vincenzo Ambriola and Giovanni A. Cignoni. A distributed virtual machine to support software process. ACM SIGSOFT Software Engineering Notes, 20 (1):85–89, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alencar:1995:FSR

[ACLN95] P. S. C. Alencar, D. D. Cowan, C. J. P. Lucena, and L. C. M. Nova. Formal specification of reusable interface objects. ACM SIGSOFT Software Engineering Notes, 20(SI):88–96, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ambriola:1990:SPE

[ACM90] V. Ambriola, P. Ciancarini, and C. Montangero. Software process enactment in oikos. ACM SIGSOFT Software Engineering Notes, 15(6):183–192, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ethics:1997:DSE

[ACM97] ACM / IEEE-CS Task Force on Software Engineering Ethics and Practice. Draft software engineering code of ethics, version 2.1. ACM SIGSOFT Software Engineering Notes, 22 (4):3–6, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Abowd:1994:ISE

[AD94] Gregory D. Abowd and Alan J. Dix. Integrating status and event phenomena in formal specifications of interactive systems. ACM SIGSOFT Software Engineering Notes, 19(5): 44–52, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Adams:1999:FRI

[Ada99] Tom Adams. A formula for the re-inspection decision. ACM SIGSOFT Software Engineering Notes, 24(3):80, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Addy:1998:RFA

[Add98] Edward A. Addy. Report from the First Annual Workshop on Software Architectures in Product Line Acquisitions. ACM SIGSOFT Software Engineering Notes, 23(3):32–39, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Allen:1997:TLM

[ADLK97] Graham Allen, Leighton Davies, Göran Lindmark, and Even-André Karlsson. TTM15 — a large multi-site improvement project. ACM SIGSOFT Software Engineering Notes, 22(6): 111–124, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Atlee:1991:SBM

[AG91] Joanne Atlee and John Gannon. State-based model checking of event-driven system requirements. ACM SIGSOFT Software Engineering Notes, 16 (5):16–28, December 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Atlee:1993:ATR

[AG93] Joanne M. Atlee and John Gannon. Analyzing timing requirements. ACM SIGSOFT Software Engineering Notes, 18 (3):117–127, July 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aho:1995:FIG

[AG95] Alfred V. Aho and Nancy D. Griffeth. Feature interactions in the global information infrastructure. ACM SIGSOFT Software Engineering Notes, 20 (4):2–4, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Atkinson:1998:EWP

[AG98] Darren C. Atkinson and William G. Griswold. Effective whole-program analysis in the presence of pointers. ACM SIGSOFT Software Engineering Notes, 23(6):46–55, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:1997:TCI

[Agg97] K. K. Aggarwal. A tool for the collection of industrial software metrics data. ACM SIGSOFT Software Engineering Notes, 22 (2):54–57, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Allen:1998:FMA

[AGI98] Robert J. Allen, David Garlan, and James Ivers. Formal modeling and analysis of the HLA component integration standard. ACM SIGSOFT Software Engineering Notes, 23 (6):70–79, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alborghetti:1997:PAS

[AGM97] Andrea Alborghetti, Angelo Gargantini, and Angelo Morzenti. Providing automated support to deductive analysis of time critical systems. ACM SIG-SOFT Software Engineering Notes, 22(6):211–226, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agrawal:1999:ECT

[Agr99] Hira Agrawal. Efficient coverage testing using global dominator graphs. ACM SIGSOFT Software Engineering Notes, 24(5):11–20, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aharonian:1999:SPI

[Aha99] Greg Aharonian. 17,500 software patents to issue in 1998. ACM SIGSOFT Software Engineering Notes, 24(3):58–62, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aiken:1990:ABS

[Aik90] M. W. Aiken. AI based simulation in reusability. ACM SIGSOFT Software Engineering Notes, 15(5):23–27, October 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Abib:1999:GBT

[AK99] Janaina C. Abib and Tereza G. Kirner. A GQM-based tool to support the development of software quality measurement plans. ACM SIGSOFT Software Engineering Notes, 24(4): 75–80, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:1993:TWO

[AKSR93] E. K. Aggarwal, M. Pavan Kumar, Vinay Santurkar, and Radha Ratnaparkhi. Towards a weighted operational profile. ACM SIGSOFT Software Engineering Notes, 18(1): 52–53, January 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Avritzer:1993:LTS

[AL93] Alberto Avritzer and Brian Larson. Load testing software using deterministic state testing. ACM SIGSOFT Software Engineering Notes, 18(3): 82–88, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Agarwal:1995:PPA

[AL95] Rakesh Agarwal and Patricia Lago. PATHOS — a paradigmatic approach to high-level object-oriented software development. ACM SIGSOFT Software Engineering Notes, 20(2): 36–41, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Amoroso:1995:CFS

[Amo95] Edward G. Amoroso. Creating formal specifications from requirements documents. ACM SIGSOFT Software Engineering Notes, 20(1):67–70, January 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anane:1998:HDP

[Ana98] R. Anane. Holistic design of a programming system. ACM SIGSOFT Software Engineering Notes, 23(1):72–76, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anderson:1995:BOC

[And95]  Bruce Anderson.  Building

organizational competence in software architecture: OOP-SLA’94 Workshop report. ACM SIGSOFT Software Engineering Notes, 20(2):25–28, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anger:1999:DNS

[Ang99] Frank D. Anger. Directions for the NSF software engineering and languages program. ACM SIGSOFT Software Engineering Notes, 24(4):50–52, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:1999:ICD

[Ano99] Anonymous. ICSE ’99 conference daily newspaper: Window on the world. ACM SIGSOFT Software Engineering Notes, 24 (4):53–74, July 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Apte:1994:PMS

[Apt94] Kishor Apte. Problem management system: experience and insights. ACM SIGSOFT Software Engineering Notes, 19(1): 53–54, January 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arango:1995:SRI

[Ara95] Guillermo Arango. Software reusability and the Internet. ACM SIGSOFT Software Engineering Notes, 20(SI):22–23, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arifoglu:1993:MSC

[Ari93] Ali Arifoglu. A methodology for software cost estimation. ACM SIGSOFT Software Engineering Notes, 18(2):96–105, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arnold:1995:CIB

[Arn95] John E. Arnold. Control integration: a briefly annotated bibliography. ACM SIGSOFT Software Engineering Notes, 20 (2):62–67, April 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:1994:MFS

[AS94] K. K. Aggarwal and Yogesh Singh. A modified form of software science measures. ACM SIGSOFT Software Engineering Notes, 19(3):48–50, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aggarwal:1995:SRA

[AS95] K. K. Aggarwal and Yogesh Singh. Software reliability apportionment using the analytic hierarchy process. ACM SIGSOFT Software Engineering Notes, 20(5):56–61, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alur:1998:MCH

[AY98] Rajeev Alur and Mihalis Yannakakis. Model checking of hierarchical state machines. ACM SIGSOFT Software Engineering Notes, 23(6):175– 188, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ayers:1995:BRC

[Aye95]  Michael Ayers.  Book re-

views: The Capability Maturity Model Guidelines for Improving the Software Process. ACM SIGSOFT Software Engineering Notes, 20(5):88–89, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ayers:1996:BRB

[Aye96a] Michael Ayers. Book review: Bringing Design to Software. ACM SIGSOFT Software Engineering Notes, 21(4):95–96, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ayers:1996:BRI

[Aye96b]  Michael Ayers.  Book re-

view: Information Modeling — An Object-Oriented Approach by Haim Kilov and James Ross. ACM SIGSOFT Software Engineering Notes, 21(2): 91–92, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ayers:1996:BRR

[Aye96c] Michael Ayers. Book review: Reliable Object-Oriented Software — Applying Analysis and Design. ACM SIGSOFT Software Engineering Notes, 21(4): 94–95, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ayers:1997:BRP

[Aye97]  Michael Ayers.  Book re-

views: Prototyping with Objects, Philippe Krief. ACM SIGSOFT Software Engineering Notes, 22(1):105–106, January 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bourque:1994:ISR

[BA94] Pierre Bourque and Alain Abran. An innovative software reengineering tools workshop — a test of market maturity and lessons learned. ACM SIGSOFT Software Engineering Notes, 19(3):30–34, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bagert:1998:TBV

[Bag98a] Donald J. Bagert. Texas board votes to license software engineers. ACM SIGSOFT Software Engineering Notes, 23(5): 7, September 1, 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bagert:1998:TPL

[Bag98b] Donald J. Bagert. Texas poised to license professional engineers in software engineering. ACM SIGSOFT Software Engineering Notes, 23(3):8–10, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bailin:1996:DPE

[Bai96] Sid Bailin. Domain processes and engineering. ACM SIG-SOFT Software Engineering Notes, 21(1):39–42, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ball:1998:LCF

[Bal98] Thomas Ball. On the limit of control flow analysis for regression test selection. ACM SIG-SOFT Software Engineering Notes, 23(2):134–142, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ball:1999:CDA

[Bal99a] Thoms Ball. The concept of dynamic analysis. ACM SIG-SOFT Software Engineering Notes, 24(6):216–234, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balmas:1999:QON

[Bal99b] Françoise Balmas. Query by outlines: a new paradigm to help manage programs. ACM SIGSOFT Software Engineering Notes, 24(5):86–94, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barghouti:1992:SCM

[Bar92] Naser S. Barghouti. Supporting cooperation in the Marvel process-centered SDE. ACM SIGSOFT Software Engineering Notes, 17(5):21–31, December 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barnes:1998:HSE

[Bar98] Bruce H. Barnes. A history of software engineering at the National Science Foundation (a personal view). ACM SIGSOFT Software Engineering Notes, 23(3):68–69, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bassett:1997:TPA

[Bas97] Paul G. Bassett. The theory and practice of adaptive reuse. ACM SIGSOFT Software Engineering Notes, 22 (3):2–9, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bull:1996:RDP

[BB96] Tim Bull and Keith Bennett.

A report on the Durham Program Transformations Workshop. ACM SIGSOFT Software Engineering Notes, 21(4): 51–53, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Basili:1999:NWS

[BBB^{+}99] Victor R. Basili, Laszlo Belady, Barry Boehm, Frederick Brooks, James Browne, Richard DeMillo, Stuart I. Feldman, Cordell Green, Butler Lampson, Duncan Lawrie, Nancy Leveson, Nancy Lynch, Mark Weiser, and Jeannette Wing. NSF Workshop on a Software Research Program for the 21st Century. ACM SIGSOFT Software Engineering Notes, 24(3):37–44, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bronsard:1997:TSP

[BBK^{+}97] François Bronsard, Douglas Bryan, W. Kozaczynski, Edy S. Liongosari, Jim Q. Ning, Ásgeir Ólafsson, and John W. Wetterstrand. Toward software plug-and-play. ACM SIGSOFT Software Engineering Notes, 22 (3):19–29, May 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berchtold:1999:SSA

[BBP99] Stefan Berchtold, Alexandros Biliris, and Euthimios Panagos. SaveMe: a system for archiving electronic documents using messaging groupware. ACM SIGSOFT Software Engineering Notes, 24(2): 167–176, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baresi:1999:WWD

[BCC^{+}99] L. Baresi, F. Casati, S. Castano, M. G. Fugini, I. Mirbel, and B. Pernici. WIDE workflow development methodology. ACM SIGSOFT Software Engineering Notes, 24(2): 19–28, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Batory:1995:CRA

[BCGS95] Don Batory, Lou Coglianese, Mark Goodwin, and Steve Shafer. Creating reference architectures: an example from avionics. ACM SIGSOFT Software Engineering Notes, 20 (SI):27–37, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barjaktarovic:1996:FSV

[BCJ96] Milica Barjaktarovic, Shiu-Kai Chin, and Kamal Jabbour. Formal specification and verification of the kernel functional unit of the OSI session layer protocol and service using CCS. ACM SIGSOFT Software Engineering Notes, 21(3):270–279, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Browne:1995:LIN

[BDG^{+}95] Shirley Browne, Jack Dongarra, Stan Green, Keith Moore, Theresa Pepin, Tom Rowan, and Reed Wade. Location-independent naming for virtual distributed software repositories. ACM SIGSOFT Software Engineering Notes, 20 (SI):179–185, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brinkkemper:1990:DMC

[BdLLvdS90] S. Brinkkemper, M. de Lange, R. Looman, and F. H. G. C. van der Steen. On the derivation of method companionship by meta-modelling. ACM SIGSOFT Software Engineering Notes, 15(1):49–58, January 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Britton:1990:TAD

[BDM90] Carol Britton, Jill Doake, and Richard Mitchell. Taming the abstract data type: a taxonomy to help information systems designers. ACM SIGSOFT Software Engineering Notes, 15(1): 36–41, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beckman:1991:DHT

[Bec91] Brian Beckman. Debugging for hypercubes in time warp. ACM SIGSOFT Software Engineering Notes, 16(2):57, April 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beck:1999:EPD

[Bec99] Kent Beck. Extreme programming: a discipline of software development (invited paper) (abstract only). ACM SIGSOFT Software Engineering Notes, 24(6):1, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boehm:1997:DMA

[BEKM97] Barry Boehm, Alex Egyed, Julie Kwan, and Ray Madachy. Developing multimedia applications with the WinWin spiral model. ACM SIGSOFT Software Engineering Notes, 22(6): 20–39, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benner:1996:OSR

[Ben96] Kevin Benner. “The Organization” (or software reuse in a business environment). ACM SIGSOFT Software Engineering Notes, 21(1):35–39, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bergadano:1993:TCG

[Ber93] F. Bergadano. Test case generation by means of learning techniques. ACM SIGSOFT Software Engineering Notes, 18(5): 149–162, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:1996:SIS

[Ber96] Lawrence Bernstein. Software investment strategy. ACM SIGSOFT Software Engineering Notes, 21(2):47–53, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Butler:1991:IEQ

[BF91] Ricky W. Butler and George B. Finelli. The infeasibility of experimental quantification of life-critical software reliability. ACM SIGSOFT Software Engineering Notes, 16(5):66–76, December 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brown:1992:ATE

[BF92] Alan W. Brown and Peter H. Feiler. An analysis technique for examining integration in a project support environment. ACM SIGSOFT Software Engineering Notes, 17(5): 139–148, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Braberman:1999:VRT

[BF99] Victor A. Braberman and Miguel Felder. Verification of real-time designs: combining scheduling theory with automatic formal verification. ACM SIGSOFT Software Engineering Notes, 24(6):494– 510, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bultan:1996:CVM

[BFG96] Tevfik Bultan, Jeffrey Fischer, and Richard Gerber. Compositional verification by model checking for counter-examples. ACM SIGSOFT Software Engineering Notes, 21(3):224–238, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balda:1990:CEM

[BG90] D. Balda and D. A. Gustafson.

Cost estimation models for reuse and prototype SW development life-cycles. ACM SIGSOFT Software Engineering Notes, 15(3):42–50, July 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowdidge:1994:ASE

[BG94] Robert W. Bowdidge and William G. Griswold. Automated support for encapsulating abstract data types. ACM SIGSOFT Software Engineering Notes, 19(5):97–110, December 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bultan:1998:VSI

[BGL98] Tevfik Bultan, Richard Gerber, and Christopher League. Verifying systems with integer constraints and Boolean predicates: a composite approach. ACM SIGSOFT Software Engineering Notes, 23(2): 113–123, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bodik:1997:RDF

[BGS97] Rastislav Bodík, Rajiv Gupta, and Mary Lou Soffa. Refining data flow information using infeasible paths. ACM SIGSOFT Software Engineering Notes, 22 (6):361–377, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ballance:1990:PLB

[BGV90] Robert A. Ballance, Susan L. Graham, and Michael L. Van De Vanter. The Pan languagebased editing system for integrated development. ACM SIGSOFT Software Engineering Notes, 15(6):77–93, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bayer:1999:TLA

[BGW^{+}99] Joachim Bayer, Jean-François Girard, Martin Würthner, Jean-Marc DeBaud, and Martin Apel. Transitioning legacy assets to a product line architecture. ACM SIGSOFT Software Engineering Notes, 24(6): 446–463, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhansali:1995:HAS

[Bha95] Sanjay Bhansali. A hybrid approach to software reuse. ACM SIGSOFT Software Engineering Notes, 20(SI):215– 218, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bucci:1994:PII

[BHKW94] Paolo Bucci, Joseph E. Hollingsworth, Joan Krone, and Bruce W. Weide. Part III: implementing components in RE-SOLVE. ACM SIGSOFT Software Engineering Notes, 19(4): 40–51, October 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Belli:1993:IBA

[BJ93] B. Belli and O. Jack. Implementationbased analysis and testing of Prolog programs. ACM SIG-SOFT Software Engineering Notes, 18(3):70–80, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

BenGhezala:1995:RAB

[BK95a] Henda Hadjami Ben Ghezala and Farouk Kamoun. A reuse approach based on object orientation: its contributions in the development of CASE tools. ACM SIGSOFT Software Engineering Notes, 20(SI):53–62, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bieman:1995:CRO

[BK95b] James M. Bieman and Byung-Kyoo Kang. Cohesion and reuse in an object-oriented system. ACM SIGSOFT Software Engineering Notes, 20(SI):259– 262, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brilliant:1999:ERS

[BK99] Susan S. Brilliant and John C. Knight. Empirical Research in Software Engineering: a workshop. ACM SIGSOFT Software Engineering Notes, 24(3): 44–52, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barrett:1996:ASS

[BKW96] Daniel J. Barrett, Alan Kaplan, and Jack C. Wileden. Automated support for seamless interoperability in polylingual software systems. ACM SIG-SOFT Software Engineering Notes, 21(6):147–155, November 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bing:1997:OBM

[BL97] Mao Bing and Xie Li. An object-based model for prototyping user interfaces of cooperative systems. ACM SIGSOFT Software Engineering Notes, 22(2):72–76, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blanqui:1997:DCA

[Bla97] Frédéric Blanqui. A documentcentered approach for an open CASE environment framework connected with the World-Wide Web. ACM SIGSOFT Software Engineering Notes, 22 (2):58–63, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blum:1993:DPC

[Blu93] Manuel Blum. Designing programs to check their work (abstract). ACM SIGSOFT Software Engineering Notes, 18 (3):1, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blyth:1997:IAM

[Bly97] Andrew Blyth. Issues arising from medical system’s failure. ACM SIGSOFT Software Engineering Notes, 22(2): 85–86, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Browne:1997:RLI

[BM97] Shirley V. Browne and James W.

Moore. Reuse library interoperability and the world wide Web. ACM SIGSOFT Software Engineering Notes, 22(3):182–189, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Butkus:1998:FEC

[BM98] Al Butkus and Barbara L.

Mitchell. A framework for encapsulating card-oriented, interactive, legacy applications. ACM SIGSOFT Software Engineering Notes, 23(1): 44–46, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Borgida:1999:TEW

[BM99] Alex Borgida and Takahiro Murata. Tolerating exceptions in workflows: a unified framework for data and processes. ACM SIGSOFT Software Engineering Notes, 24(2):59–68, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bahsoun:1993:FPF

[BMS93] Jean Paul Bahsoun, Stephan Merz, and Corinne Servieres. A framework for programming and formalizing concurrent objects. ACM SIGSOFT Software Engineering Notes, 18(5): 126–137, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Natan:1992:MAT

[BN92] Ron Ben-Natan. Making all types typedefed. ACM SIG-SOFT Software Engineering Notes, 17(2):72–73, April 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Balzer:1993:MGP

[BN93] Robert Balzer and K. Narayanaswamy. Mechanisms for generic process support. ACM SIGSOFT Software Engineering Notes, 18(5): 21–32, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brennan:1999:WDE

[BO99] Susan E. Brennan and Justina O.

Ohaeri. Why do electronic conversations seem less polite? the costs and benefits of hedging. ACM SIGSOFT Software Engineering Notes, 24(2): 227–235, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boehm:1995:REN

[Boe95] Barry Boehm. Reuse emphasized at next process workshop. ACM SIGSOFT Software Engineering Notes, 20(5): 17, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boehm:1996:IPS

[Boe96] Barry Boehm. Industrial priorities for software engineering research (panel). ACM SIGSOFT Software Engineering Notes, 21 (6):2, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boehm:1999:EST

[Boe99] Barry Boehm. Escaping the software tar pit: model clashes and how to avoid them. ACM SIGSOFT Software Engineering Notes, 24(1):36–48, January 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bokowski:1999:CSC

[Bok99] Boris Bokowski. CoffeeStrainer: statically-checked constraints on the definition and use of types in Java. ACM SIGSOFT Software Engineering Notes, 24 (6):355–374, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bolognesi:1990:SGR

[Bol90] Tommaso Bolognesi. On the soundness of graphical representations of interconnected processes in LOTOS. ACM SIGSOFT Software Engineering Notes, 15(4):1–7, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Borstler:1995:BRR

[Bor95] Jurgen Borstler. Book review: Reengineering Information Technology Success Through Empowerment. ACM SIGSOFT Software Engineering Notes, 20(5):89, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boundy:1991:TP

[Bou91] David Boundy. A taxonomy of programmers. ACM SIGSOFT Software Engineering Notes, 16 (4):23–30, October 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boundy:1993:SCS

[Bou93] David Boundy. Software cancer: the seven early warning signs. ACM SIGSOFT Software Engineering Notes, 18 (2):19, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowles:1990:NYS

[Bow90] Adrion J. Bowles. A note on the yourdon structured method. ACM SIGSOFT Software Engineering Notes, 15(2):27, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brown:1992:ABI

[BP92] Alan W. Brown and Maria H.

Penedo. An annotated bibliography on integration in software engineering environments. ACM SIGSOFT Software Engineering Notes, 17(3):47–55, July 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Billard:1994:QSG

[BR94] Edward A. Billard and Alice E.

Riedmiller. Q-Sim: a GUI for a queueing simulator using Tcl/Tk. ACM SIGSOFT Software Engineering Notes, 19 (4):82–85, October 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Billard:1995:GML

[BR95] Edward A. Billard and Alice E. Riedmiller. A GUI for a manager of lightweight processes. ACM SIGSOFT Software Engineering Notes, 20(5): 48–50, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brandt:1990:QMD

[Bra90] Dennis L. Brandt. Quality measures in design: finding problems before coding. ACM SIGSOFT Software Engineering Notes, 15(1):68–72, January 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Britchef:1995:FPF

[Bri95] Bob Britchef. A few (proposed) fundamental laws of programming. ACM SIGSOFT Software Engineering Notes, 20(2): 19–20, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brown:1990:CCQ

[Bro90] Bradley J. Brown. Correctness is not congruent with quality. ACM SIGSOFT Software Engineering Notes, 15(2):28, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bruce:1996:IAM

[Bru96]  Michael Bruce.  Internet

agility/maturity model survey for the software industry call for participation. ACM SIG-SOFT Software Engineering Notes, 21(1):75–76, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brykczynski:1999:SSI

[Bry99] Bill Brykczynski. A survey of software inspection checklists. ACM SIGSOFT Software Engineering Notes, 24(1):82, January 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Shaul:1992:AMU

[BSKH92] Israel Z. Ben-Shaul, Gail E.

Kaiser, and George T. Heineman. An architecture for multiuser software development environments. ACM SIGSOFT Software Engineering Notes, 17 (5):149–158, December 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Batory:1993:SSL

[BSST93] Don Batory, Vivek Singhal, Marty Sirkin, and Jeff Thomas. Scalable software libraries. ACM SIGSOFT Software Engineering Notes, 18(5):191– 199, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Batory:1994:RCA

[BTS94] Don Batory, Jeff Thomas, and Marty Sirkin. Reengineering a complex application using a scalable data structure compiler. ACM SIGSOFT Software Engineering Notes, 19(5):

111–120, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Buckley:1991:PAS

[Buc91] Fletcher J. Buckley. Perestroika, anyone? A standard for democracy in professional organizations. ACM SIGSOFT Software Engineering Notes, 16 (3):70–72, July 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bukovsky:1991:PAS

[Buk91] Nikolay S. Bukovsky. A practical approach to software quality assurance. ACM SIGSOFT Software Engineering Notes, 16 (1):68–72, January 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Burstler:1996:PLF

[Bur96] Jurgen Burstler. Pattern languages o f program design. ACM SIGSOFT Software Engineering Notes, 21(1):88–89, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brykczynski:1993:ABS

[BW93] Bill Brykczynski and David A. Wheeler. An annotated bibliography on software inspections. ACM SIGSOFT Software Engineering Notes, 18(1): 81–88, January 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bieman:1995:RTI

[BZ95] James M. Bieman and Josephine Xia Zhao. Reuse through inheritance: a quantitative study of C++ software. ACM SIG-SOFT Software Engineering Notes, 20(SI):47–52, 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bose:1999:WWA

[BZ99] Prasanta Bose and Xiaoqing Zhou. WWAC: WinWin abstraction based decision coordination. ACM SIGSOFT Software Engineering Notes, 24(2): 127–136, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Corbett:1993:PTB

[CA93] J. Corbett and G. Avrunin. A practical technique for bounding the time between events in concurrent real-time systems. ACM SIGSOFT Software Engineering Notes, 18(3):110–116, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Corbett:1994:TSC

[CA94] James C. Corbett and George S.

Avrunin. Towards scalable compositional analysis. ACM SIGSOFT Software Engineering Notes, 19(5):53–61, December 1994. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Chan:1998:IES

[CABN98] William Chan, Richard J. Anderson, Paul Beame, and David Notkin. Improving efficiency of symbolic model checking for state-based system requirements. ACM SIGSOFT Software Engineering Notes, 23(2): 102–112, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carmel:1993:DSC

[Car93] Erran Carmel. A discussion of special characteristics for software package development life cycle models. ACM SIGSOFT Software Engineering Notes, 18 (2):23–24, April 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carey:1996:SQI

[Car96] Dick Carey. Is software quality intrinsic, subjective, or relational? ACM SIGSOFT Software Engineering Notes, 21(1): 74–75, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Casey:1999:OMS

[Cas99] Richard M. Casey. Object mappings in a software engineering project. ACM SIGSOFT Software Engineering Notes, 24 (1):55–59, January 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Churchill:1999:VEW

[CB99] Elizabeth F. Churchill and Sara Bly. Virtual environments at work: ongoing use of MUDs in the workplace. ACM SIGSOFT Software Engineering Notes, 24 (2):99–108, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Coomber:1990:GTP

[CC90] C. J. Coomber and R. E.

Childs. A graphical tool for the prototyping of real-time systems. ACM SIGSOFT Software Engineering Notes, 15(2): 70–82, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chamillard:1996:IAP

[CC96] A. T. Chamillard and Lori A.

Clarke. Improving the accuracy of Petri net-based analysis of concurrent programs. ACM SIGSOFT Software Engineering Notes, 21(3):24–38, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cabral:1990:ISM

[CCCL90] Regina H. B. Cabral, Ivan M.

Campos, Donald D. Cowan, and Carlos J. P. Lucena. Interfaces as specifications in the MIDAS user interface development systems. ACM SIGSOFT Software Engineering Notes, 15 (2):55–69, April 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Coomer:1990:DRS

[CCR90] T. N. Coomer, Jr., J. R. Comer, and D. J. Rodjak. Developing reusable software for military systems, why it is needed, why it isn’t working. ACM SIGSOFT Software Engineering Notes, 15(3):33–38, July 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Caron:1990:BTD

[CD90] Jacqueline M. Caron and Peter A. Darnell. Bugfind: a tool for debugging optimizing compilers. ACM SIGSOFT Software Engineering Notes, 15(1): 64–65, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Castano:1994:FRE

[CD94] Silvana Castano and Valeria De Antonellis. The f3 reuse environment for requirements engineering. ACM SIGSOFT Software Engineering Notes, 19 (3):62–65, July 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carrington:1993:DMD

[CDHW93] David Carrington, David Duke, Ian Hayes, and Jim Welsh. Deriving modular designs from formal specifications. ACM SIGSOFT Software Engineering Notes, 18(5):89–98, December 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Castano:1995:BRC

[CDP95] S. Castano, V. De Antonellis, and B. Pernici. Building reusable components in the public administration domain. ACM SIGSOFT Software Engineering Notes, 20(SI):81–87, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Czarnecki:1999:CGP

[CE99] Krzysztof Czarnecki and Ulrich W. Eisenecker. Components and generative programming (invited paper). ACM SIGSOFT Software Engineering Notes, 24(6):2–19, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Conradi:1993:CFE

[CFF93] Reidar Conradi, Christer Fernström, and Alfonso Fuggetta. A conceptual framework for evolving software processes. ACM SIGSOFT Software Engineering Notes, 18(4):26– 35, October 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cece:1994:DIL

[CFI94] Gérard Cécé, Alain Finkel, and S. Purushothaman Iyer. Duplication, insertion and lossiness errors in unreliable communication channels. ACM SIGSOFT Software Engineering Notes, 19 (5):35–43, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Choi:1999:EPM

[CGHS99] Jong-Deok Choi, David Grove, Michael Hind, and Vivek Sarkar. Efficient and precise modeling of exceptions for the analysis of Java programs. ACM SIGSOFT Software Engineering Notes, 24(5): 21–31, September 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:1997:CDM

[CGK97a] Yih-Farn R. Chen, Emden R. Gansner, and Eleftherios Koutsofios. A C++ data model supporting reachability analysis and dead code detection. ACM SIGSOFT Software Engineering Notes, 22(6): 414–431, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheung:1997:VLP

[CGK97b] Shing Chi Cheung, Dimitra Giannakopoulou, and Jeff Kramer. Verification of liveness properties using compositional reachability analysis. ACM SIGSOFT Software Engineering Notes, 22(6):227– 243, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Caplan:1995:LFS

[CH95] Joshua E. Caplan and Mehdi T.

Harandi. A logical framework for software proof reuse. ACM SIGSOFT Software Engineering Notes, 20(SI):106– 113, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cherry:1991:RMV

[Che91a] George W. Cherry. S-R machines: a visual formalism for reactive and interactive systems. ACM SIGSOFT Software Engineering Notes, 16(3): 52–55, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cherry:1991:SCOa

[Che91b]  George W. Cherry.  Soft-

ware construction by objectoriented pictures: stimulusresponse machines. ACM SIGSOFT Software Engineering Notes, 16(2):52–56, April 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cherry:1991:SCOb

[Che91c] George W. Cherry. System construction with object-oriented pictures. ACM SIGSOFT Software Engineering Notes, 16(4): 42–52, October 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chelini:1992:DAR

[Che92a] James V. Chelini. A discussion on the Ada Run-Time Environment in safety critical applications. ACM SIGSOFT Software Engineering Notes, 17(4): 24–27, October 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheng:1992:PSS

[Che92b] Jingwen Cheng. Parameterized specifications for software reuse. ACM SIGSOFT Software Engineering Notes, 17(4): 53–59, October 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cherry:1992:GFS

[Che92c] George W. Cherry. Graphic formalisms should integrate communication, control, and data flow. ACM SIGSOFT Software Engineering Notes, 17(2): 64–69, April 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheng:1993:ISR

[Che93a] Jingwen Cheng. Improving the software reusability in objectoriented programming. ACM SIGSOFT Software Engineering Notes, 18(4):70–74, October 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cherry:1993:SRM

[Che93b] George W. Cherry. Stimulusresponse machines: a new visual formalism for describing classes and objects. ACM SIGSOFT Software Engineering Notes, 18(2):86–95, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheng:1994:RBS

[Che94a] Jingwen Cheng. A reusabilitybased software development environment. ACM SIGSOFT Software Engineering Notes, 19 (2):57–62, April 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cherry:1994:VSE

[Che94b] George W. Cherry. Visual software engineering with rules. ACM SIGSOFT Software Engineering Notes, 19(3):26–29, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chan:1990:COO

[CHS90] M. L. Chan and B. Henderson-Sellers. Corporate Objectoriented Development Environment (CODE). ACM SIG-SOFT Software Engineering Notes, 15(1):42–43, January 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carasik:1990:DDG

[CJPV90] R. P. Carasik, S. M. Johnson, D. A. Patterson, and G. A. Von Glahn. Domain description grammar: application of linguistic semantics. ACM SIGSOFT Software Engineering Notes, 15(5):28–43, October 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheung:1993:ECR

[CK93] S. C. Cheung and J. Kramer. Enhancing compositional reachability analysis with context constraints. ACM SIGSOFT Software Engineering Notes, 18 (5):115–125, December 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheung:1995:CRA

[CK95] S. C. Cheung and J. Kramer. Compositional reachability analysis of finite-state distributed systems with user-specified constraints. ACM SIGSOFT Software Engineering Notes, 20(4):140–150, October 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cameron:1991:RTT

[CL91] E. Jane Cameron and Yow-Jian Lin. A real-time transition model for analyzing behavioral compatibility of telecommunications services. ACM SIG-SOFT Software Engineering Notes, 16(5):101–111, December 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clark:1990:FVD

[Cla90] Jon D. Clark. Function versus data-driven methodologies: a prescriptive metric. ACM SIGSOFT Software Engineering Notes, 15(2):26, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clement:1990:DAP

[Clé90] Dominique Clément. A distributed architecture for programming environments. ACM SIGSOFT Software Engineering Notes, 15(6):11–21, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Crowley:1996:IFS

[CLL96] J. L. Crowley, J. F. Leathrum, and K. A. Liburdy. Issues in the full scale use of formal methods for automated testing. ACM SIGSOFT Software Engineering Notes, 21(3):71–78, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Callahan:1996:AVV

[CM96] John R. Callahan and Todd L.

Montgomery. An approach to verification and validation of a reliable multicasting protocol. ACM SIGSOFT Software Engineering Notes, 21(3):187–194, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:1998:AFPc

[CM98] Greg Cooper and Arnulf Mester. ACM Fellow profile: Meir M (Manny) Lehman. ACM SIGSOFT Software Engineering Notes, 23(5):11, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Connors:1992:SDM

[Con92]  Danny T. Connors.  Soft-

ware development methodologies and traditional and modern information systems. ACM SIGSOFT Software Engineering Notes, 17(2):43–49, April 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Conn:1999:NCS

[Con99] Richard Conn. Notes from the 12th conference on software engineering education and training (CSEET) and 13th SIGCSE technical symposium on computer science education (CSE). ACM SIGSOFT Software Engineering Notes, 24(4):30–32, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cook:1996:SDG

[Coo96] Michael Cook. Software design is a good thing. ACM SIG-SOFT Software Engineering Notes, 21(5):56–57, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:1998:AFPa

[Coo98a] Greg Cooper. ACM Fellow profile: Dennis J. Frailey. ACM SIGSOFT Software Engineering Notes, 23(4):9–11, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:1998:AFPb

[Coo98b] Greg Cooper. ACM Fellow profile: Grady Booch. ACM SIGSOFT Software Engineering Notes, 23(5):8–10, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Corbett:1996:CAM

[Cor96] James C. Corbett. Constructing abstract models of concurrent real-time software. ACM SIGSOFT Software Engineering Notes, 21(3):250–260, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Corbett:1998:CCM

[Cor98] James C. Corbett. Constructing compact models of concurrent Java programs. ACM SIGSOFT Software Engineering Notes, 23(2):1–10, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cai:1993:TIP

[CP93] Jiazhen Cai and Robert Paige. Towards increased productivity of algorithm implementation. ACM SIGSOFT Software Engineering Notes, 18(5): 71–78, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Coen-Porisini:1993:CAR

[CPK93] Alberto Coen-Porisini and Richard A. Kemmerer. The composability of ASTRAL realtime specifications. ACM SIGSOFT Software Engineering Notes, 18(3):128–138, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cross:1995:SWC

[CQW^{+}95] James H. Cross II, Alex Quilici, Linda Wills, Philip Newcomb, and Elliot Chikofsky. Second working conference on reverse engineering summary report. ACM SIGSOFT Software Engineering Notes, 20(5): 23–26, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chandra:1999:PTC

[CR99a] Satish Chandra and Thomas Reps. Physical type checking for C. ACM SIGSOFT Software Engineering Notes, 24(5): 66–75, September 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chang:1999:SSB

[CR99b] Juei Chang and Debra J.

Richardson. Structural specificationbased testing: automated support and experimental evaluation. ACM SIGSOFT Software Engineering Notes, 24(6): 285–302, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chang:1996:SSB

[CRS96] Juei Chang, Debra J. Richardson, and Sriram Sankar. Structural specification-based testing with ADL. ACM SIGSOFT Software Engineering Notes, 21 (3):62–70, May 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chung:1990:EED

[CS90] Anthony Chung and Deepinder Sidhu. Experience with an Estelle development system. ACM SIGSOFT Software Engineering Notes, 15(4): 8–17, September 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Churcher:1995:TCF

[CS95] Neville I. Churcher and Martin J. Shepperd. Towards a conceptual framework for object oriented software metrics. ACM SIGSOFT Software Engineering Notes, 20(2):69–75, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:1997:RES

[CS97] Yeh-Ling Chen and Arnold J. Stromberg. Robust estimation in software experiments. ACM SIGSOFT Software Engineering Notes, 22(4):60–64, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carreira:1998:WDS

[CS98] João Carreira and João Gabriel Silva. Why do some (weird) people inject faults? ACM SIGSOFT Software Engineering Notes, 23(1):42–43, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Clements:1998:NSI

[CW98a] Paul C. Clements and Nelson Weiderman. Notes on the Second International Workshop on Development and Evolution of Software Architectures for Product Families. ACM SIGSOFT Software Engineering Notes, 23(3):39–43, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cook:1998:EBD

[CW98b] Jonathan E. Cook and Alexander L. Wolf. Event-based detection of concurrency. ACM SIGSOFT Software Engineering Notes, 23(6):35–45, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:1998:MTO

[CYZ98] Jiajun Chen, Xiaodong Yuan, and Guolian Zhengp. A multi-threaded object-oriented programming model. ACM SIGSOFT Software Engineering Notes, 23(3):83–86, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dai:1995:DRE

[Dai95]  W. Dai.  Development of

reusable expert system components: preliminary experience. ACM SIGSOFT Software Engineering Notes, 20(SI):238– 246, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dandanell:1991:RDU

[Dan91] Bent Dandanell. Rigorous development using RAISE. ACM SIGSOFT Software Engineering Notes, 16(5):29–43, December 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:1995:ARC

[Dav95a] Margaret J. Davis. Adaptable, reusable code. ACM SIGSOFT Software Engineering Notes, 20(SI):38–46, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:1995:PPD

[Dav95b] Margaret J. Davis. Process and product: Dichotomy or duality? ACM SIGSOFT Software Engineering Notes, 20 (2):17–18, April 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:1996:BIU

[Dav96] Margaret J. Davis. Barriers to institutionalizing using current tools and environments. ACM SIGSOFT Software Engineering Notes, 21(1):52–54, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Day:1996:URC

[Day96] Donald L. Day. User responses to constraints in computerized design tools (extended abstract). ACM SIGSOFT Software Engineering Notes, 21(5): 47–50, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Duri:1993:USS

[DBDS93] S. Duri, U. Buy, R. Devarapalli, and S. M. Shatz. Using state space reduction methods for deadlock analysis in Ada tasking. ACM SIGSOFT Software Engineering Notes, 18 (3):51–60, July 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dini:1997:FSA

[DBM97] Petre Dini, Amina Belkhelladi, and Walcélio L. Melo. Formalizing software architectures: an industrial experience. ACM SIGSOFT Software Engineering Notes, 22(6): 527–529, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dwyer:1994:DFA

[DC94] Matthew B. Dwyer and Lori A.

Clarke. Data flow analysis for verifying properties of concurrent programs. ACM SIGSOFT Software Engineering Notes, 19(5):62–75, December 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dwyer:1997:MCG

[DCH97] Matthew B. Dwyer, Vicki Carr, and Laura Hines. Model checking graphical user interfaces using abstractions. ACM SIG-SOFT Software Engineering Notes, 22(6):244–261, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DeMan:1990:MLM

[De 90] Jozef De Man. Making languages more powerful by removing limitations. ACM SIG-SOFT Software Engineering Notes, 15(4):18–24, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ducasse:1991:ODEa

[DE91a] Mireille Ducassé and Anna-Maria Emde. OPIUM: a debugging environment for Prolog development and debugging research. ACM SIGSOFT Software Engineering Notes, 16(1): 54–59, January 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ducasse:1991:ODEb

[DE91b] Mireille Ducassé and Anna-Maria Emde. OPIUM: a debugging environment for Prolog development and debugging research. ACM SIGSOFT Software Engineering Notes, 16 (2):67–72, April 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Delgado:1990:ITS

[Del90] Joseph F. Delgado. Issues in transfer of skills in software engineering training. ACM SIGSOFT Software Engineering Notes, 15(2):23–25, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Denney:1996:CMB

[Den96] Richard Denney. A comparison of the model-based & algebraic styles of specification as a basis for test specification. ACM SIGSOFT Software Engineering Notes, 21(5): 60–64, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DeTreville:1991:CT

[DeT91] John DeTreville. A cautionary tale. ACM SIGSOFT Software Engineering Notes, 16(2): 19–22, April 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Desharnais:1997:ISS

[DFKM97] Jules Desharnais, Marc Frappier, Ridha Khédri, and Ali Mili. Integration of sequential scenarios. ACM SIGSOFT Software Engineering Notes, 22 (6):310–326, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Demeyer:1998:RWO

[DH98] Serge Demeyer and Harald Hall. Report: Workshop on Object-Oriented Reengineering (WOOR’97). ACM SIGSOFT Software Engineering Notes, 23(1):28–29, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubois:1995:GTT

[DHvL^{+}95] E. Dubois, J. Hagelstein, A. van Lamsweerde, F. Orejas, J. Souquieres, and P. Wodon. A guided tour through the ICARUS project. ACM SIG-SOFT Software Engineering Notes, 20(2):28–33, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Disney:1998:IDQ

[DJ98] Anne M. Disney and Philip M. Johnson. Investigating data quality problems in the PSP. ACM SIGSOFT Software Engineering Notes, 23(6):143–152, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Damon:1996:CRS

[DJJ96] Craig A. Damon, Daniel Jackson, and Somesh Jha. Checking relational specifications with binary decision diagrams. ACM SIGSOFT Software Engineering Notes, 21(6):70–80, November 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dossick:1999:CMB

[DK99] Stephen E. Dossick and Gail E. Kaiser. CHIME: a metadatabased distributed software development environment. ACM SIGSOFT Software Engineering Notes, 24(6):464–475, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Denney:1996:WSA

[DKLS96] Richard Denney, Dick Kemmerer, Nancy Leveson, and Alberto Savoia. Why state-of-the-art is not state-of-the-practice (panel). ACM SIGSOFT Software Engineering Notes, 21 (3):293, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Debest:1992:RCE

[DKW92] Xavier A. Debest, Rüdiger Knoop, and Jürgen Wagner. Reveng: a cost-effective approach to reverse-engineering. ACM SIGSOFT Software Engineering Notes, 17(4):60– 67, October 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1996:SNSa

[Doe96a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 21 (2):7–10, March 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1996:SNSb

[Doe96b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 21 (4):6–9, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1996:SNSc

[Doe96c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 21(5): 6–10, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1997:BRD

[Doe97a]  Mark Doernhoefer.  Book

review: Developing Object-Oriented Software An Experienced-Based Approach by IBM Object-Oriented Technology Center (PTR Prentice Hall, 1997). ACM SIGSOFT Software Engineering Notes, 22 (4):113–114, July 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1997:BROc

[Doe97b]  Mark Doernhoefer.  Book

review: Object-Oriented Development Process and Metrics. ACM SIGSOFT Software Engineering Notes, 22(2): 100, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1997:BROa

[Doe97c] Mark Doernhoefer. Book reviews: Object-Oriented Systems Design: an Integrated Approach, Edward Yourdon. ACM SIGSOFT Software Engineering Notes, 22(1):106–107, January 1997. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1997:BROb

[Doe97d] Mark Doernhoefer. Book reviews: Object-Oriented Systems Design: an Integrated Approach, Edward Yourdon. ACM SIGSOFT Software Engineering Notes, 22(1):106–107, January 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1997:SNSa

[Doe97e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 22 (1):8–12, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1997:SNSb

[Doe97f] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 22 (2):8–15, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1997:SNSc

[Doe97g] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 22 (4):13–20, July 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1997:SNSd

[Doe97h] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 22(5): 5–9, September 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1998:BRO

[Doe98a] Mark Doernhoefer. Book reviews: Object Models Strategies, Patterns, and Applications, by Peter Coad. ACM SIGSOFT Software Engineering Notes, 23(3):125, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1998:SNSa

[Doe98b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 23(1): 16–23, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1998:SNSb

[Doe98c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 23 (3):13–19, May 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1999:SNSa

[Doe99a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 24 (3):15–24, May 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:1999:SNSb

[Doe99b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 24 (4):15–25, July 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dolado:1991:SDG

[Dol91] José Javier Dolado. Structured development of graphgrammars for icon manipulation. ACM SIGSOFT Software Engineering Notes, 16(3): 46–51, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubois:1998:RWS

[DOP98] Eric Dubois, Andreas L.

Opdahl, and Klaus Pohl. REFSQ’97 workshop summary. ACM SIGSOFT Software Engineering Notes, 23(1): 30–34, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dovenbarger:1997:STC

[Dov97] Dana Dovenbarger. Software technology conference information. ACM SIGSOFT Software Engineering Notes, 22 (2):6, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dowson:1997:ASF

[Dow97] Mark Dowson. The ariane 5 software failure. ACM SIGSOFT Software Engineering Notes, 22(2):84, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dwyer:1998:FBM

[DP98] Matthew B. Dwyer and Corina S. Pasareanu. Filter-based model checking of partial systems. ACM SIGSOFT Software Engineering Notes, 23(6): 189–202, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leite:1996:WRS

[dPL96] Julio Cesar Sampaio do Prado Leite. Working results on software reengineering. ACM SIGSOFT Software Engineering Notes, 21 (2):39–44, March 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DeMillo:1996:CSS

[DPS96] Richard A. DeMillo, Hsin Pan, and Eugene H. Spafford. Critical slicing for software fault localization. ACM SIGSOFT Software Engineering Notes, 21 (3):121–134, May 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Duncan:1990:OMT

[DR90] I. M. M. Duncan and D. J. Robson. Ordered mutation testing. ACM SIGSOFT Software Engineering Notes, 15(2):29–30, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dillon:1996:GOY

[DR96] L. K. Dillon and Y. S. Ramakrishna. Generating oracles from your favorite temporal logic specifications. ACM SIGSOFT Software Engineering Notes, 21 (6):106–117, November 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ducasse:1997:ECT

[DR97] Stéphane Ducasse and Tamar Richner. Executable connectors: towards reusable design elements. ACM SIGSOFT Software Engineering Notes, 22 (6):483–499, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Drori:1996:PDI

[Dro96] Offer Drori. Planning and design of information systems using OODPM. ACM SIGSOFT Software Engineering Notes, 21(5):95–98, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Drori:1997:TPH

[Dro97a] Offer Drori. From theory to practice or how not to fail in developing information systems. ACM SIGSOFT Software Engineering Notes, 22(1): 85–87, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Drori:1997:HIC

[Dro97b] Offer Drori. Hypertext implications for CASE environments. ACM SIGSOFT Software Engineering Notes, 22(4):35–38, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Drori:1999:ITR

[Dro99] Offer Drori. Integration of text retrieval technology into formatted (conventional) information systems. ACM SIGSOFT Software Engineering Notes, 24 (1):78–80, January 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:1994:CPM

[DS94a] Alan M. Davis and Pradip Sitaram. A concurrent process model of software development. ACM SIGSOFT Software Engineering Notes, 19(2):38–51, April 1994. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Dedene:1994:MRM

[DS94b] G. Dedene and M. Snoeck.

M.E.R.O.DE.: a model-driven entity-relationship object-oriented development method. ACM SIGSOFT Software Engineering Notes, 19(3):51–61, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Devanbu:1997:CVT

[DS97] Prem Devanbu and Stuart G.

Stubblebine. Cryptographic verification of test coverage claims. ACM SIGSOFT Software Engineering Notes, 22(6): 395–413, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Daran:1996:SEA

[DTF96] Murial Daran and Pascale Thévenod-Fosse. Software error analysis: a real case study involving real faults and mutations. ACM SIGSOFT Software Engineering Notes, 21 (3):158–171, May 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dusink:1995:RD

[DvK95] Liesbeth Dusink and Jan van Katwijk. Reuse dimensions. ACM SIGSOFT Software Engineering Notes, 20(SI):137– 149, 1995. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Darimont:1996:FRP

[DvL96] Robert Darimont and Axel van Lamsweerde. Formal refinement patterns for goaldriven requirements elaboration. ACM SIGSOFT Software Engineering Notes, 21(6): 179–190, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Davis:1997:SAC

[DW97] Margaret J. Davis and Roger B. Williams. Software architecture characterization. ACM SIGSOFT Software Engineering Notes, 22(3):30–38, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dillon:1994:OCT

[DY94] Laura K. Dillon and Qing Yu. Oracles for checking temporal properties of concurrent systems. ACM SIGSOFT Software Engineering Notes, 19(5): 140–153, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dyadkin:1994:MP

[Dya94] Lev J. Dyadkin. Multibox parsers. ACM SIGSOFT Software Engineering Notes, 19(3): 23–25, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Estublier:1999:BFP

[EAD99] Jacky Estublier, Mahfound Amiour, and Samir Dami. Building a federation of process support systems. ACM SIG-SOFT Software Engineering Notes, 24(2):197–206, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ebert:1997:ECP

[Ebe97]  Christof Ebert.  Experi-

ences with criticality predictions in software development. ACM SIGSOFT Software Engineering Notes, 22(6):278– 293, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edelstein:1993:RIS

[Ede93] D. Vera Edelstein. Report on the IEEE STD 1219–1993 — standard for software maintenance. ACM SIGSOFT Software Engineering Notes, 18(4): 94–95, October 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edwards:1994:PVA

[Edw94] Stephen H. Edwards. Part v: annotated bibliography of RESOLVE research. ACM SIGSOFT Software Engineering Notes, 19(4):64–67, October 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edwards:1996:MAS

[Edw96]  Steve Edwards.  Micro-

architecture of software components and the need for good mental models of software subsystems. ACM SIGSOFT Software Engineering Notes, 21(1): 46–50, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edwards:1999:SRP

[Edw99] Stephen H. Edwards. The state of reuse: perceptions of the reuse community. ACM SIGSOFT Software Engineering Notes, 24(3):32–36, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eggert:1990:TSP

[Egg90] Paul Eggert. Toward specialpurpose program verification. ACM SIGSOFT Software Engineering Notes, 15(4):25–29, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Evans:1994:LTU

[EGHT94] David Evans, John Guttag, James Horning, and Yang Meng Tan. LCLint: a tool for using specifications to check code. ACM SIGSOFT Software Engineering Notes, 19 (5):87–96, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edwards:1994:PIS

[EHL^{+}94] Stephen H. Edwards, Wayne D.

Heym, Timothy J. Long, Murali Sitaraman, and Bruce W. Weide. Part II: specifying components in RESOLVE. ACM SIGSOFT Software Engineering Notes, 19(4):29–39, October 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Engels:1997:VOA

[EHTE97] Gregor Engels, Reiko Heckel, Gabi Taentzer, and Hartmut Ehrig. A view-oriented approach to system modelling based on graph transformation. ACM SIGSOFT Software Engineering Notes, 22(6): 327–343, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eickelmann:1999:AFPb

[Eic99a] Nancy Eickelmann. ACM Fellow profile: Barbara G. Ryder. ACM SIGSOFT Software Engineering Notes, 24(4):13–14, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eickelmann:1999:AFPa

[Eic99b] Nancy Eickelmann. ACM Fellow profile: David Lorge Parnas. ACM SIGSOFT Software Engineering Notes, 24(3): 10–14, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Engel:1998:DAC

[ELB^{+}98] Gerald L. Engel, Richard J. LeBlanc, Bruce Ho Barnes, Martin Lo Griss, Tony Wasserman, and Laurie Werth. Draft accreditation criteria for software engineering. ACM SIG-SOFT Software Engineering Notes, 23(3):11–12, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ellmer:1998:LPS

[Ell98] Ernst Ellmer. A learning perspective on software process technology. ACM SIGSOFT Software Engineering Notes, 23 (4):65–69, July 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Edwards:1997:WAW

[EW97] Stephen H. Edwards and Bruce W. Weide. WISR8: 8th Annual Workshop on Software Reuse: summary and working group reports. ACM SIGSOFT Software Engineering Notes, 22(5):17–32, September 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Favaro:1996:SPC

[Fav96] John Favaro. On the scalability problem in COTSbased programming environments. ACM SIGSOFT Software Engineering Notes, 21(5): 43–46, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fink:1997:PBT

[FB97a] George Fink and Matt Bishop.

Property-based testing: a new approach to testing for assurance. ACM SIGSOFT Software Engineering Notes, 22(4): 74–80, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Forgacs:1997:FTP

[FB97b] István Forgács and Antonia Bertolino. Feasible test path selection by principal slicing. ACM SIGSOFT Software Engineering Notes, 22(6):378–394, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fong:1998:PLA

[FC98] Philip W. L. Fong and Robert D. Cameron. Proof linking: an architecture for modular verification of dynamicallylinked mobile code. ACM SIG-SOFT Software Engineering Notes, 23(6):222–230, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Favaro:1994:ESS

[FCC94] John Favaro, Yves Coene, and Marco Casucci. The European space software development environment reference facility project: a status report.

ACM SIGSOFT Software Engineering Notes, 19(2):68–71, April 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkbine:1995:SI

[FD95] Ronald B. Finkbine and Ph. D. Software inspection. ACM SIGSOFT Software Engineering Notes, 20(5):90, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkbine:1996:MMS

[FD96] Ronald B. Finkbine and Ph. D. Metrics and models in software quality engineering. ACM SIGSOFT Software Engineering Notes, 21(1):89, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fekete:1995:LCM

[Fek95] Alan Fekete. Liveness conditions in model-based service specifications: a case study. ACM SIGSOFT Software Engineering Notes, 20(4): 62–71, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Francalanci:1997:IIR

[FF97] C. Francalanci and A. Fuggetta. Integrating information requirements along processes: a survey and research directions. ACM SIGSOFT Software Engineering Notes, 22(1):

68–74, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Felder:1993:ARS

[FGP93] Miguel Felder, Carlo Ghezzi, and Mauro Pezzè. Analyzing refinements of state based specifications: the case of TB nets. ACM SIGSOFT Software Engineering Notes, 18(3):28–39, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

France:1995:ADA

[FH95] Robert B. France and Thomas B.

Horton. Applying domain analysis and modeling: an industrial experience. ACM SIG-SOFT Software Engineering Notes, 20(SI):206–214, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ferrans:1992:HFH

[FHS^{+}92] James C. Ferrans, David W.

Hurst, Michael A. Sennett, Burton M. Covnot, Wenguang Ji, Peter Kajka, and Wei Ouyang. HyperWeb: a framework for hypermedia-based environments. ACM SIGSOFT Software Engineering Notes, 17 (5):1–10, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frankl:1998:FES

[FI98] Phyllis G. Frankl and Oleg Iakounenko. Further empirical studies of test effectiveness. ACM SIGSOFT Software Engineering Notes, 23(6): 153–162, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fierz:1999:CMC

[Fie99] Hugo Fierz. The CIP method: component- and model-based construction of embedded systems. ACM SIGSOFT Software Engineering Notes, 24(6): 375–392, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkelsteiin:1992:RE

[Fin92a]  Anthony Finkelsteiin.  Re-

quirements engineering. ACM SIGSOFT Software Engineering Notes, 17(1):45–47, January 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkelsteiin:1992:SPI

[Fin92b]  Anthony Finkelsteiin.  A

software process immaturity model. ACM SIGSOFT Software Engineering Notes, 17(4): 22–23, October 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkbine:1996:BRSa

[Fin96a] Ronald B. Finkbine. Book review: Software Development Using Eiffel. ACM SIGSOFT Software Engineering Notes, 21 (4):96, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkbine:1996:BRSb

[Fin96b] Ronald B. Finkbine. Book review: Software Fault Tolerance. ACM SIGSOFT Software Engineering Notes, 21(4):96, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkbine:1997:BIU

[Fin97] Robert B. Finkbine. Barriers to institutionalizing using current tools and environments: an introduction to functional programming systems using Haskell. ACM SIGSOFT Software Engineering Notes, 22 (2):99, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finneran:1998:RTC

[Fin98] Lisa Finneran. Report: technology challenges at SPC. ACM SIGSOFT Software Engineering Notes, 23(1):35– 37, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Finkbine:1999:AFP

[Fin99] Ronald B. Finkbine. ACM Fellow profile: L. Peter Deutsch.

ACM SIGSOFT Software Engineering Notes, 24(1):21, January 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fisch:1993:UIU

[Fis93] Eric A. Fisch. Understanding and improving the user interface design process. ACM SIGSOFT Software Engineering Notes, 18(2):27–29, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fowler:1995:PWR

[FKV95] Glenn S. Fowler, David G. Korn, and Kiem-Phong Vo. Principles for writing reusable libraries. ACM SIGSOFT Software Engineering Notes, 20 (SI):150–159, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fraser:1995:PTD

[FLM95] Steven Fraser, Deborah Leishman, and Robert McLellan. Patterns, teams and domain engineering. ACM SIGSOFT Software Engineering Notes, 20 (SI):222–224, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fradet:1999:CCM

[FLP99] Pascal Fradet, Daniel Le Métayer, and Michaël Périn. Consistency checking for multiple view software architectures. ACM SIGSOFT Software Engineering Notes, 24(6):

410–428, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fiadeiro:1995:IFS

[FM95] José Luiz Fiadeiro and Tom Maibaum. Interconnecting formalisms: supporting modularity, reuse and incrementality. ACM SIGSOFT Software Engineering Notes, 20(4): 72–80, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Foster:1991:AST

[Fos91] Kenneth A. Foster. Arithmetic statement testing. ACM SIGSOFT Software Engineering Notes, 16(3):74–77, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fuggetta:1994:ABS

[FP94] Alfonso Fuggetta and Gian Pietro Picco. An annotated bibliography on software process improvement. ACM SIGSOFT Software Engineering Notes, 19 (3):66–68, July 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Favaro:1998:MSD

[FP98] John Favaro and Shari Lawrence Pfleeger. Making software development investment decisions. ACM SIGSOFT Software Engineering Notes, 23(5):

69–74, September 1, 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Field:1999:IPS

[FR99] John Field and G. Ramalingam. Identifying procedural structure in Cobol programs. ACM SIGSOFT Software Engineering Notes, 24(5): 1–10, September 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frakes:1996:RP

[Fra96] Bill Frakes. The reuse of processes. ACM SIGSOFT Software Engineering Notes, 21(1): 46, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frost:1990:CPC

[Fro90] Richard A. Frost. Constructing programs in a calculus of lazy interpreters. ACM SIGSOFT Software Engineering Notes, 15(4):30–41, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fuggetta:1996:IPG

[Fug96] Alfonso Fuggetta. ICSE 97: picking up the gauntlet. ACM SIGSOFT Software Engineering Notes, 21(5):5–6, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fuggetta:1999:SRS

[Fug99] Alfonso Fuggetta. Some reflections on software engineering research. ACM SIGSOFT Software Engineering Notes, 24 (1):74–77, January 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feather:1994:SSI

[FvL94] Martin S. Feather and Axel van Lamsweerde. Succeedings of the Seventh International Workshop on Software Specification and Design. ACM SIGSOFT Software Engineering Notes, 19(3):18–22, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Frankl:1991:AFD

[FW91] Phyllis G. Frankl and Elaine J.

Weyuker. Assessing the faultdetecting ability of testing methods. ACM SIGSOFT Software Engineering Notes, 16(5): 77–91, December 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fan:1998:RAT

[FXHZ98] Xiaocong Fan, Dianxiang Xu, Jianmin Hou, and Guoliang Zheng. Reasoning about team tracking. ACM SIGSOFT Software Engineering Notes, 23(3): 79–82, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garcia:1996:PM

[GA96] Manuel J. Barranco García and Juan Carlos Granja Alvarez. Productive maintainability. ACM SIGSOFT Software Engineering Notes, 21(2): 89–91, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gacek:1995:EDA

[Gac95] Cristina Gacek. Exploiting domain architectures in software reuse. ACM SIGSOFT Software Engineering Notes, 20 (SI):229–232, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gafni:1999:RRT

[Gaf99]  Vered Gafni.  Robots: a

real-time systems architectural style. ACM SIGSOFT Software Engineering Notes, 24(6): 57–74, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gennari:1995:RPI

[GAM95] John H. Gennari, Russ B. Altman, and Mark A. Musen. Reuse with PROTÉGÉ-II: from elevators to ribosomes. ACM SIGSOFT Software Engineering Notes, 20(SI):72–80, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gambhir:1997:UDA

[Gam97] S. S. Gulu Gambhir. Use of domain analysis to implement the developer off-the-shelf systems (DOTSS) system acquisition approach. ACM SIGSOFT Software Engineering Notes, 22 (2):48–53, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1994:ESA

[GAO94] David Garlan, Robert Allen, and John Ockerbloom. Exploiting style in architectural design environments. ACM SIGSOFT Software Engineering Notes, 19 (5):175–188, December 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1990:RFR

[Gar90] David Garlan. The role of formal reusable frameworks. ACM SIGSOFT Software Engineering Notes, 15(4):42–44, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1995:FIW

[Gar95a] David Garlan. First International Workshop on Architectures for Software Systems workshop summary. ACM SIGSOFT Software Engineering Notes, 20(3):84–89, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1995:SAP

[Gar95b] David Garlan. Software architecture (panel): next steps towards an engineering discipline for software systems design. ACM SIGSOFT Software Engineering Notes, 20(4):5, October 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1997:SPM

[Gar97] David Garlan. SIGSOFT’96 post mortem. ACM SIGSOFT Software Engineering Notes, 22 (1):23, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gyimothy:1999:ERS

[GBF99] Tibor Gyimóthy, Árpád Beszédes, and Istán Forgács. An efficient relevant slicing method for debugging. ACM SIGSOFT Software Engineering Notes, 24(6): 303–321, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gotlieb:1998:ATD

[GBR98] Arnaud Gotlieb, Bernard Botella, and Michel Rueher. Automatic test data generation using constraint solving techniques. ACM SIGSOFT Software Engineering Notes, 23 (2):53–62, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

George:1996:SMF

[GC96] Joseph George and Bradley D. Carter. A strategy for mapping from function-oriented software models to object-oriented software models. ACM SIGSOFT Software Engineering Notes, 21 (2):56–63, March 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griswold:1996:TSP

[GCBM96] William G. Griswold, Morison I. Chen, Robert W. Bowdidge, and J. David Morgenthaler. Tool support for planning the restructuring of data abstractions in large systems. ACM SIGSOFT Software Engineering Notes, 21(6): 33–45, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1992:TAG

[GCN92] David Garlan, Linxi Cai, and Robert L. Nord. A transformational approach to generating application-specific environments. ACM SIGSOFT Software Engineering Notes, 17 (5):68–77, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gelperin:1994:PIM

[Gel94] David Gelperin. The power of integrated methods. ACM SIGSOFT Software Engineering Notes, 19(4):77–78, October 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guerrieri:1997:IJS

[GGL^{+}97] Ernesto Guerrieri, Martin Griss, Doug Lea, Prashant Sridharan, and Will Tracz. The impact of Java on software reusability. ACM SIGSOFT Software Engineering Notes, 22 (3):210–211, May 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gargantini:1999:UMC

[GH99] Angelo Gargantini and Constance Heitmeyer. Using model checking to generate tests from requirements specifications. ACM SIGSOFT Software Engineering Notes, 24(6): 146–162, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ghezzi:1999:IO

[Ghe99] Carlo Ghezzi. ICSE 2000 opportunity. ACM SIGSOFT Software Engineering Notes, 24 (3):8, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Godefroid:1998:MCM

[GHJ98] Patrice Godefroid, Robert S. Hanmer, and Lalita Jategaonkar Jagadeesan. Model checking without a model: an analysis of the heart-beat monitor of a telephone switch using VeriSoft. ACM SIGSOFT Software Engineering Notes, 23 (2):124–133, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Graham:1992:PPS

[GHM92] Susan L. Graham, Michael A.

Harrison, and Ethan V. Munson. The Proteus presentation system. ACM SIGSOFT Software Engineering Notes, 17(5): 130–138, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1990:LCA

[GI90] David Garlan and Ehsan Ilias.

Low-cost, adaptable tool integration policies for integrated environments. ACM SIGSOFT Software Engineering Notes, 15 (6):1–10, December 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Girgensohn:1999:SWR

[Gir99] Andreas Girgensohn. Supporting the writing of reports in a hierarchical organization. ACM SIGSOFT Software Engineering Notes, 24(2): 147–156, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garg:1994:SAB

[GJ94] Pankaj K. Garg and Mehdi Jazayeri. Selected, annotated bibliography on processcentered software engineering environments. ACM SIGSOFT Software Engineering Notes, 19 (2):18–21, April 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griss:1995:SSR

[GJJ^{+}95] Martin Griss, Ivar Jacobson, Chris Jette, Bob Kessler, and Doug Lea. Systematic software reuse (panel): objects and frameworks are not enough. ACM SIGSOFT Software Engineering Notes, 20(SI):17–20, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gall:1995:RDS

[GJK95] Harald Gall, Mehdi Jazayeri, and René Klösch. Research directions in software reuse: where to go from here? ACM SIGSOFT Software Engineering Notes, 20(SI):225– 228, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1998:RAI

[GJND98] D. Garlan, S. Jha, D. Notkin, and J. Dingel. Reasoning about implicit invocation. ACM SIG-SOFT Software Engineering Notes, 23(6):209–221, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grimaud:1999:FTI

[GLV99] Gilles Grimaud, Jean-Louis Lanet, and Jean-Jacques Vandewalle. FACADE: a typed intermediate language dedicated to smart cards. ACM SIG-SOFT Software Engineering Notes, 24(6):476–493, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1996:FSF

[GM96] David Garlan and Mark Morieoni. The fourth symposium on the foundations of software engineering (FSE4). ACM SIGSOFT Software Engineering Notes, 21(5):4, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gintell:1997:LLA

[GM97] John W. Gintell and Gérard Memmi. Lessons learned with ACME an environment integration experiment. ACM SIGSOFT Software Engineering Notes, 22(2):77–81, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Giannakopoulou:1999:CPA

[GMK99] Dimitra Giannakopoulou, Jeff Magee, and Jeff Kramer. Checking progress with action priority: is it fair? ACM SIG-SOFT Software Engineering Notes, 24(6):511–527, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gotterbarn:1999:CRR

[GMR99] Don Gotterbarn, Keith Miller, and Simon Rogerson. On code reuse: a response. ACM SIGSOFT Software Engineering Notes, 24(3):4–6, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Par99a].

Gupta:1998:ATD

[GMS98] Neelam Gupta, Aditya P. Mathur, and Mary Lou Soffa. Automated test data generation using an iterative relaxation method. ACM SIGSOFT Software Engineering Notes, 23 (6):231–244, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griswold:1992:CAV

[GN92] William G. Griswold and David Notkin. Computer-aided vs. manual program restructuring. ACM SIGSOFT Software Engineering Notes, 17(1): 33–41, January 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goldberg:1990:RSD

[Gol90] Allen Goldberg. Reusing software developments. ACM SIG-SOFT Software Engineering Notes, 15(6):107–119, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gomaa:1995:DMM

[Gom95] Hassan Gomaa. Domain modeling methods and environments. ACM SIGSOFT Software Engineering Notes, 20 (SI):256–258, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goradia:1993:DIA

[Gor93] Tarak Goradia. Dynamic impact analysis: a cost-effective technique to enforce errorpropagation. ACM SIGSOFT Software Engineering Notes, 18 (3):171–181, July 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gorecki:1994:YCS

[Gor94] Andrzej Tomasz Gorecki. Is your computer system subcritical? ACM SIGSOFT Software Engineering Notes, 19 (2):65–67, April 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gossain:1994:USM

[Gos94] Sanjiv Gossain. Using the Shlaer–Mellor method to support a level 4 software process. ACM SIGSOFT Software Engineering Notes, 19(3):69–74, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gotterbarn:1998:SEP

[Got98] Don Gotterbarn. Software engineering as a profession.

ACM SIGSOFT Software Engineering Notes, 23(6):205–206, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gotterbarn:1999:PST

[Got99] Donald Gotterbarn. A positive step toward a profession: the software engineering code of ethics and professional practice. ACM SIGSOFT Software Engineering Notes, 24(1): 9–14, January 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Godefroid:1996:UPO

[GPS96] Patrice Godefroid, Doron Peled, and Mark Staskauskas. Using partial-order methods in the formal validation of industrial concurrent programs. ACM SIGSOFT Software Engineering Notes, 21(3):261–269, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Greenwood:1997:MAL

[Gre97] Garrison W. Greenwood. So many algorithms. so little time. ACM SIGSOFT Software Engineering Notes, 22 (2):92–93, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griswold:1993:DUD

[Gri93] William G. Griswold. Direct update of data flow representations for a meaningpreserving program restructuring tool. ACM SIGSOFT Software Engineering Notes, 18(5): 42–55, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griss:1996:SOR

[Gri96] Martin L. Griss. Systematic OO reuse — a tale of two cultures. ACM SIGSOFT Software Engineering Notes, 21(1): 50–52, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griss:1998:SEP

[Gri98] Martin Griss. Software engineering as a profession: industry and academia working together. ACM SIGSOFT Software Engineering Notes, 23(6): 203–204, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grinter:1999:SAP

[Gri99a] Rebecca E. Grinter. Systems architecture: product designing and social engineering. ACM SIGSOFT Software Engineering Notes, 24(2):11–18, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grinter:1999:WT

[Gri99b] Rebecca E. Grinter. WACC’99 tutorials. ACM SIGSOFT Software Engineering Notes, 24(2): 239–240, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grinter:1999:WWP

[Gri99c] Rebecca E. Grinter. WACC’99 workshop program. ACM SIG-SOFT Software Engineering Notes, 24(2):237–238, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gruhn:1990:MSP

[Gru90] Volker Gruhn. Managing software processes in the environment MELMAC. ACM SIG-SOFT Software Engineering Notes, 15(6):193–205, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gyorkos:1990:ICT

[GRW90] J. Györkös, I. Rozman, and T. Welzer. Introducing CASE tools into the software development group. ACM SIGSOFT Software Engineering Notes, 15 (3):39–41, July 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gupta:1995:HSA

[GS95] Rajiv Gupta and Mary Lou Soffa. Hybrid slicing: an approach for refining static slices using dynamic information. ACM SIGSOFT Software Engineering Notes, 20(4):

29–40, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garg:1996:FCS

[GS96] Pankaj K. Garg and Sriram Sankar. Fifth California software symposium. ACM SIG-SOFT Software Engineering Notes, 21(1):68–70, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Griss:1993:WFA

[GT93] Martin Griss and Will Tracz.

WISR’92: Fifth Annual Workshop in Software Reuse: working group reports. ACM SIGSOFT Software Engineering Notes, 18(2):74–85, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garlan:1995:SDW

[GTP95] David Garlan, Walter Tichy, and Frances Paulisch. Summary of the Dagstuhl Workshop on Software Architecture. ACM SIGSOFT Software Engineering Notes, 20(3):63–83, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gustafson:1993:SMS

[GTW93] David A. Gustafson, Joo T.

Tan, and Perla Weaver. Software measure specification. ACM SIGSOFT Software Engineering Notes, 18(5):163– 168, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gunter:1996:ADB

[Gun96] Carl A. Gunter. Abstracting dependencies between software configuration items. ACM SIG-SOFT Software Engineering Notes, 21(6):167–178, November 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guo:1997:ATD

[Guo97] Minyi Guo. Automatic transformation from data flow diagram to structure chart. ACM SIGSOFT Software Engineering Notes, 22(4):44–49, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:1994:FST

[Ham94] Dick Hamlet. Foundations of software testing: dependability theory. ACM SIGSOFT Software Engineering Notes, 19(5): 128–139, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:1996:PDT

[Ham96] Dick Hamlet. Predicting dependability by testing. ACM SIGSOFT Software Engineering Notes, 21(3):84–91, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:1998:WCW

[Ham98]  Dick Hamlet.  What can

we learn by testing a program? ACM SIGSOFT Software Engineering Notes, 23(2): 50–52, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hansen:1995:SS

[Han95] Brinch Hansen. The SuperPascal software. ACM SIGSOFT Software Engineering Notes, 20 (5):91, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Haneef:1998:SDR

[Han98] Nuzhat J. Haneef. Software documentation and readability: a proposed process improvement. ACM SIGSOFT Software Engineering Notes, 23 (3):75–77, May 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henry:1994:DWR

[HB94] Joel Henry and Bob Blasewitz.

Do we really need SQA to produce quality software?: no! well maybe. it depends. yes! ACM SIGSOFT Software Engineering Notes, 19(2):63–64, April 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hilerio:1999:HEI

[HC99] Israel Hilerio and Weidong Chen. Herbal-T, enabling integration, interoperability, and reusability of Internet components. ACM SIGSOFT Software Engineering Notes, 24(2): 49–58, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hecht:1990:COO

[Hec90] Alan Hecht. Cute objectoriented acronyms considered FOOlish. ACM SIGSOFT Software Engineering Notes, 15(1): 48, January 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hefley:1993:CCA

[Hef93] William E. Hefley. The Cobbler’s children: applying total quality management to business process improvement, information engineering and software engineering. ACM SIG-SOFT Software Engineering Notes, 18(4):19–25, October 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heimbigner:1992:PPS

[Hei92] Dennis Heimbigner. The ProcessWall: a process state server approach to process programming. ACM SIGSOFT Software Engineering Notes, 17(5): 159–168, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heimdahl:1996:ELA

[Hei96] Mats P. E. Heimdahl. Experiences and lessons from the analysis of TCAS II. ACM SIGSOFT Software Engineering Notes, 21(3):79–83, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henninger:1995:DDK

[Hen95a] Scott Henninger. Developing domain knowledge through the reuse of project experiences. ACM SIGSOFT Software Engineering Notes, 20(SI):186– 195, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henninger:1995:SPS

[Hen95b] Scott Henninger. Supporting the process of satisfying information needs with reusable software libraries: an empirical study. ACM SIGSOFT Software Engineering Notes, 20 (SI):267–270, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Herbsleb:1999:MRC

[Her99] James D. Herbsleb. Metaphorical representation in collaborative software engineering. ACM SIGSOFT Software Engineering Notes, 24(2):117– 126, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hajnal:1998:ATD

[HF98] Ákos Hajnal and István Forgács. An applicable test data generation algorithm for domain errors. ACM SIGSOFT Software Engineering Notes, 23 (2):63–72, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hilera:1998:SDE

[HGGM98] José R. Hilera, León A. González, José A. Guitiérrez, and J. M. Martinez. Software documentation as an engineering process. ACM SIGSOFT Software Engineering Notes, 23 (5):61–64, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heinl:1999:CAF

[HHJ^{+}99] Petra Heinl, Stefan Horn, Stefan Jablonski, Jens Neeb, Katrin Stein, and Michael Teschke. A comprehensive approach to flexibility in workflow management systems. ACM SIGSOFT Software Engineering Notes, 24(2):79–88, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hauswirth:1999:CCM

[HJ99] Manfred Hauswirth and Mehdi Jazayeri. A component and communication model for push systems. ACM SIGSOFT Software Engineering Notes, 24(6):

20–38, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hayes:1994:UDB

[HJN94] I. J. Hayes, C. B. Jones, and J. E. Nicholls. Understanding the differences between VDM and z. ACM SIGSOFT Software Engineering Notes, 19(3): 75–81, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hangensen:1992:CSS

[HK92] Torben Mejlvang Hangensen and Bent Bruun Kristensen. Consistency in software system development: framework, model, techniques & tools. ACM SIGSOFT Software Engineering Notes, 17(5):58–67, December 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Houdek:1997:QPA

[HK97] Frank Houdek and Hubert Kempter. Quality patterns — an approach to packaging software engineering experience. ACM SIGSOFT Software Engineering Notes, 22(3):81–88, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:1998:MIP

[HKMR98] Richard Hamlet, Richard Kemmerer, Edward F. Miller, and Debra J. Richardson. The most influential papers from the ISSTA research community (panel). ACM SIGSOFT Software Engineering Notes, 23 (2):49, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hagemeister:1992:ABS

[HLO^{+}92] J. Hagemeister, B. Lowther, P. Oman, X. Yu, and W. Zhu. An annotated bibliography on software maintenance. ACM SIGSOFT Software Engineering Notes, 17(2):79–84, April 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hull:1999:DWS

[HLS^{+}99] Richard Hull, François Llirbat, Eric Siman, Jianwen Su, Guozhu Dong, Bharat Kumar, and Gang Zhou. Declarative workflows that support easy modification and dynamic browsing. ACM SIGSOFT Software Engineering Notes, 24 (2):69–78, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hollingsworth:1997:UCD

[HM97] Jeffrey K. Hollingsworth and Ethan L. Miller. Using contentderived names for configuration management. ACM SIGSOFT Software Engineering Notes, 22 (3):104–109, May 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrold:1993:ECP

[HMR93] Mary Jean Harrold, Brian Malloy, and Gregg Rothermel. Efficient construction of program dependence graphs. ACM SIGSOFT Software Engineering Notes, 18(3):160–170, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hoare:1993:AM

[Hoa93] C. A. R. Hoare. Algebra and models. ACM SIGSOFT Software Engineering Notes, 18(5): 1–8, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Holbrook:1990:SBM

[Hol90] H. Holbrook III. A scenariobased methodology for conducting requirements elicitation. ACM SIGSOFT Software Engineering Notes, 15(1): 95–104, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Holloway:1995:SEE

[Hol95] C. Michael Holloway. Software engineering and epistemology. ACM SIGSOFT Software Engineering Notes, 20(2):20–21, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hollenbach:1996:BRS

[Hol96] Frank Hollenbach. Book review: Software Systems Construction with Examples in Ada. ACM SIGSOFT Software Engineering Notes, 21(4):97, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Holmes:1999:SEW

[Hol99] Neville Holmes. Software engineering: to be or what to be? ACM SIGSOFT Software Engineering Notes, 24(3):81–83, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Howell:1995:SNSa

[How95a] Chuck Howell. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 20(1): 7–15, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Howell:1995:SNSb

[How95b] Chuck Howell. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 20 (3):2–7, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Howell:1996:SNS

[How96] Chuck Howell. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 21(1): 6–10, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrold:1994:PDF

[HR94] Mary Jean Harrold and Gregg Rothermel. Performing data flow testing on classes. ACM SIGSOFT Software Engineering Notes, 19(5):154–163, December 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrold:1996:SCA

[HR96] Mary Jean Harrold and Gregg Rothermel. Separate computation of alias information for reuse. ACM SIGSOFT Software Engineering Notes, 21(3): 107–120, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hansen:1991:SVR

[HRR91] Kirsten M. Hansen, Anders P.

Ravn, and Hans Rischel. Specifying and verifying requirements of real-time systems. ACM SIGSOFT Software Engineering Notes, 16(5):44–54, December 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Horwitz:1995:DID

[HRS95] Susan Horwitz, Thomas Reps, and Mooly Sagiv. Demand interprocedural dataflow analysis. ACM SIGSOFT Software Engineering Notes, 20(4): 104–115, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrold:1998:CIC

[HRS98] Mary Jean Harrold, Gregg Rothermel, and Saurabh Sinha. Computation of interprocedural control dependence. ACM SIGSOFT Software Engineering Notes, 23(2):11–20, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hart:1990:EDS

[HS90] C. Frederick Hart and John J. Shilling. An environment for documenting software features. ACM SIGSOFT Software Engineering Notes, 15(6): 120–132, December 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hoffman:1995:SAM

[HS95] Daniel Hoffman and Paul Strooper. State abstraction and modular software development. ACM SIGSOFT Software Engineering Notes, 20(4): 53–61, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrison:1996:UTC

[HS96a] R. Harrison and L. G. Samaraweera. Using test case metrics to predict code quality and effort. ACM SIGSOFT Software Engineering Notes, 21(5):

78–88, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson-Sellers:1996:MVS

[HS96b]  B. Henderson-Sellers.  The

mathematical validity of software metrics. ACM SIGSOFT Software Engineering Notes, 21(5):89–94, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Howden:1996:LSE

[HS96c] W. E. Howden and G. M.

Shi. Linear and structural event sequence analysis. ACM SIGSOFT Software Engineering Notes, 21(3):98–106, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hughes:1996:DSA

[HS96d] Merlin Hughes and David Stotts. Daistish: systematic algebraic testing for OO programs in the presence of sideeffects. ACM SIGSOFT Software Engineering Notes, 21(3): 53–61, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson-Sellers:1993:MOO

[HSE93] B. Henderson-Sellers and J. M.

Edwards. The O-O-O methodology for the object-oriented life cycle. ACM SIGSOFT Software Engineering Notes, 18(4):

54–60, October 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Henderson-Sellers:1992:CCO

[HSF92] B. Henderson-Sellers and C. Freeman. Cataloguing and classification for object libraries. ACM SIGSOFT Software Engineering Notes, 17(1):62– 64, January 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hollingsworth:1994:PIR

[HSWZ94] Joseph E. Hollingsworth, Sethu Sreerama, Bruce W. Weide, and Sergey Zhupanov. Part IV: RESOLVE components in Ada and C++. ACM SIGSOFT Software Engineering Notes, 19 (4):52–63, October 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huang:1998:FHO

[Hua98a] Riri Huang. Formalizing hierarchical object-oriented design method. ACM SIGSOFT Software Engineering Notes, 23(5): 82–88, September 1, 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huang:1998:MAC

[Hua98b] Riri Huang. Making active CASE tools — toward the next generation CASE tools. ACM SIGSOFT Software Engineering Notes, 23(1):47–50, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hughes:1991:MSE

[Hug91] David K. Hughes. Multilingual software engineering using Ada and C. ACM SIGSOFT Software Engineering Notes, 16 (4):55–59, October 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Humphrey:1995:WSY

[Hum95] Watts S. Humphrey. Why should you use a personal software process? ACM SIGSOFT Software Engineering Notes, 20 (3):33–36, July 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:1993:FSA

[HV93] Dick Hamlet and Jeff Voas.

Faults on its sleeve: amplifying software reliability testing. ACM SIGSOFT Software Engineering Notes, 18(3):89–98, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Heimdahl:1997:RSH

[HW97] Mats P. E. Heimdahl and Michael W. Whalen. Reduction and slicing of hierarchical state machines. ACM SIGSOFT Software Engineering Notes, 22 (6):450–467, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hong:1999:CAS

[HZG99] Liu Hong, Lin Zongkai, and Zeng Guangzhou. A construction approach for software agents using components. ACM SIGSOFT Software Engineering Notes, 24(3):76–79, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Idri:1997:TAC

[IGE97] A. Idri, B. Griech, and A. El Iraki. Towards an adaptation of the COCOMO cost model to the software measurement theory. ACM SIGSOFT Software Engineering Notes, 22(6): 525–526, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ires:1997:BRU

[Ire97] Barry L. Ires. Book review: Using Z by Jim Woodcock and Jim Davies (Prentice Hall, 1996). ACM SIGSOFT Software Engineering Notes, 22(4): 114–115, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Issarny:1991:EHM

[Iss91] Valérie Issarny. An exception handling model for parallel programming and its verification. ACM SIGSOFT Software Engineering Notes, 16(5):

92–100, December 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ives:1997:BRM

[Ive97]  Barry L. Ives.  Book re-

view: ML for the Working Programmer by L. C. Paulson (Cambridge University Press, 1996). ACM SIGSOFT Software Engineering Notes, 22 (4):114, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ipser:1990:MFS

[IWJ90] Edward A. Ipser, David S.

Wile, and Dean Jacobs. A multi-formalism specification environment. ACM SIGSOFT Software Engineering Notes, 15(6):94–106, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jacky:1990:FSC

[Jac90] Jonathan Jacky. Formal specification for a clinical cyclotron control system. ACM SIG-SOFT Software Engineering Notes, 15(4):45–54, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:1993:AAA

[Jac93]  Daniel Jackson.  Abstract

analysis with aspect. ACM SIGSOFT Software Engineering Notes, 18(3):19–27, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:1996:PMS

[Jac96] Michael Jackson. Problems, methods, and structures (abstract). ACM SIGSOFT Software Engineering Notes, 21(6): 1, November 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:1998:IDL

[Jac98] Daniel Jackson. An intermediate design language and its analysis. ACM SIGSOFT Software Engineering Notes, 23(6): 121–130, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jankowski:1994:FCS

[Jan94]  David J. Jankowski.  The

feasibility of CASE structured analysis methodology support. ACM SIGSOFT Software Engineering Notes, 19(2):72–82, April 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jarzabek:1995:RLE

[Jar95] Stan Jarzabek. From reuse library experiences to application generation architectures. ACM SIGSOFT Software Engineering Notes, 20(SI):114– 122, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jarzabek:1997:MMD

[Jar97] Stan Jarzabek. Modeling multiple domains in software reuse. ACM SIGSOFT Software Engineering Notes, 22(3):65–74, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jazaa:1995:TBS

[Jaz95] Adid Jazaa. Toward better software automation. ACM SIGSOFT Software Engineering Notes, 20(1):79–84, January 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jazayeri:1999:IS

[Jaz99] Mehdi Jazayeri. Invitation to SSR. ACM SIGSOFT Software Engineering Notes, 24 (3):7, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Juell:1994:SNR

[JBVW94] Paul Juell, Daniel Brekke, Ronald Vetter, and John Wasson. Storage and network requirements of a low-cost computer-based virtual classroom. ACM SIGSOFT Software Engineering Notes, 19(3): 41–44, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jeng:1995:SMS

[JC95] Jun-Jang Jeng and Betty H. C.

Cheng. Specification matching for software reuse: a foundation. ACM SIGSOFT Software Engineering Notes, 20(SI):97– 105, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:1996:ESA

[JD96] Daniel Jackson and Craig A. Damon. Elements of style: analyzing a software design feature with a counterexample detector. ACM SIGSOFT Software Engineering Notes, 21(3): 239–249, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jensen:1999:CAA

[JDL^{+}99] David Jensen, Yulin Dong, Barbara Staudt Legner, Eric K. McCall, Leon J. Osterweil, Stanley M. Sutton, Jr., and Alexander Wise. Coordinating agent activities in knowledge discovery processes. ACM SIG-SOFT Software Engineering Notes, 24(2):137–146, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jelly:1996:CRD

[JG96] Innes Jelly and Ian Gorton. Current research directions in software engineering for parallel and distributed systems. ACM SIGSOFT Software Engineering Notes, 21(4):44–46, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jaramillo:1999:CCA

[JGS99] Clara Jaramillo, Rajiv Gupta, and Mary Lou Soffa. Comparison checking: an approach to avoid debugging of optimized code. ACM SIGSOFT Software Engineering Notes, 24(6): 268–284, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jeffords:1998:AGS

[JH98] Ralph Jeffords and Constance Heitmeyer. Automatic generation of state invariants from requirements specifications. ACM SIGSOFT Software Engineering Notes, 23(6): 56–69, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jian:1992:MAF

[Jia92] Lu Jian. A method of acquiring formal specifications from examples. ACM SIGSOFT Software Engineering Notes, 17(1): 52–56, January 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jarzabek:1999:SBC

[JK99] Stan Jarzabek and Peter Knauber. Synergy between component-based and generative approaches. ACM SIG-SOFT Software Engineering Notes, 24(6):429–445, November 1999. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Joeris:1997:CMN

[Joe97] Gregor Joeris. Change management needs integrated process and configuration management. ACM SIGSOFT Software Engineering Notes, 22(6): 125–141, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:1996:SES

[Joh96a] D. M. Johnson. The systems engineer and the software crisis. ACM SIGSOFT Software Engineering Notes, 21 (2):64–73, March 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:1996:EFA

[Joh96b] Philip Johnson. Egret: a framework for advanced CSCW applications. ACM SIGSOFT Software Engineering Notes, 21(5):70–71, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:1997:CFP

[Joh97] Ralph E. Johnson. Components, frameworks, patterns. ACM SIGSOFT Software Engineering Notes, 22(3):10–17, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jordan:1990:EPE

[Jor90] Mick Jordan. An extensible programming environment for modula-3. ACM SIGSOFT Software Engineering Notes, 15 (6):66–76, December 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jorgensen:1995:QQB

[Jør95] Magne Jørgensen. The quality of questionnaire based software maintenance studies. ACM SIGSOFT Software Engineering Notes, 20(1):71–73, January 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jimenez-Perez:1997:MSS

[JPB97] Guillermo Jiménez-Pérez and Don Batory. Memory simulators and software generators. ACM SIGSOFT Software Engineering Notes, 22(3):136–145, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jackson:1994:NMP

[JR94] Daniel Jackson and Eugene J.

Rollins. A new model of program dependences for reverse engineering. ACM SIGSOFT Software Engineering Notes, 19 (5):2–10, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jean:1990:ETO

[JS90] C. Jean and A. Strohmeier. An experience in teaching OOD for Ada software. ACM SIGSOFT Software Engineering Notes, 15 (5):44–49, October 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jager:1999:UUS

[JSW99] Dirk Jäger, Ansgar Schleicher, and Bernhard Westfechtel. Using UML for software process modeling. ACM SIGSOFT Software Engineering Notes, 24 (6):91–108, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jahnke:1997:GFR

[JSZ97] Jens H. Jahnke, Wilhelm Schäfer, and Albert Zündorf. Generic fuzzy reasoning nets as a basis for reverse engineering relational database applications. ACM SIGSOFT Software Engineering Notes, 22(6): 193–210, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Juric:1998:UR

[Jur98] Radmila Juric. The UML rules. ACM SIGSOFT Software Engineering Notes, 23(1): 92–97, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jianmin:1998:CMC

[JXXG98] Hou Jianmin, Li Xuandong, Fan Xiaocon, and Zheng Guoliang. Compositional modelchecking for real-time systems. ACM SIGSOFT Software Engineering Notes, 23(1): 99–ff., January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jing:1995:MHL

[JZZ^{+}95] Ying Jing, He Zhijun, Wu Zhaohui, Li Jiangyun, Fan Weicheng, and Xu Zhaohui. A methodology for high-level software specification construction. ACM SIGSOFT Software Engineering Notes, 20(2):48–54, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kadia:1992:IEB

[Kad92]  R. Kadia.  Issues encoun-

tered in building a flexible software development environment: lessons from the Arcadia project. ACM SIGSOFT Software Engineering Notes, 17(5): 169–180, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaindl:1993:MLR

[Kai93] Hermann Kaindl. The missing link in requirements engineering. ACM SIGSOFT Software Engineering Notes, 18 (2):30–39, April 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kamat:1991:CR

[Kam91]  Vikas K. Kamat.  Code

reader. ACM SIGSOFT Software Engineering Notes, 16(1): 60–61, January 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Karasick:1998:AMO

[Kar98] Michael Karasick. The architecture of Montana: an open and extensible programming environment with an incremental C++ compiler. ACM SIG-SOFT Software Engineering Notes, 23(6):131–142, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Korel:1998:ART

[KAY98] Bogdan Korel and Ali M. Al-Yami. Automated regression test generation. ACM SIG-SOFT Software Engineering Notes, 23(2):143–152, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1998:ICM

[KBZP98] Peter Kokol, Janez Brest, Milan Zorman, and Vili Podgorelec. Integration of complexity metrics with the use of decision trees. ACM SIGSOFT Software Engineering Notes, 23 (1):56–58, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kelsey:1995:PTM

[Kel95] Robert Bruce Kelsey. “A plea for tolerance in matters epistemological… ”. ACM SIGSOFT Software Engineering Notes, 20 (5):39, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kelsey:1996:BFC

[Kel96a] Robert Bruce Kelsey. Bad fixes, change specifications, and linguistic constraints on problem diagnosis. ACM SIGSOFT Software Engineering Notes, 21 (2):74–78, March 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kelsey:1996:BRM

[Kel96b] Robert Bruce Kelsey. Book review: A MAP For Software Acquisition. ACM SIGSOFT Software Engineering Notes, 21 (4):99, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kelsey:1996:BRQ

[Kel96c] Robert Bruce Kelsey. Book review: A Quantitative Approach to Software Management: The ami Handbook [ISBN 0-201-87746-5, 179 pages, $24.69]. ACM SIGSOFT Software Engineering Notes, 21(4):98–99, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kelsey:1996:BRI

[Kel96d] Robert Bruce Kelsey. Book review: An ISO 9000 Approach To Building Quality Software. ACM SIGSOFT Software Engineering Notes, 21(4):97–98, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kelsey:1996:BRH

[Kel96e] Robert Bruce Kelsey. Book review: How To Run Successful Projects. ACM SIGSOFT Software Engineering Notes, 21 (4):99, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kelsey:1997:BRE

[Kel97a] Robert Bruce Kelsey. Book review: Education for an Information Age: Teaching in the Computerized Classroom. ACM SIGSOFT Software Engineering Notes, 22(2):99, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kelsey:1997:IDT

[Kel97b] Robert Bruce Kelsey. Integrating a defect typology with containment metrics. ACM SIGSOFT Software Engineering Notes, 22(2):64–67, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kemmerer:1997:SID

[Kem97] Richard A. Kemmerer. Security issues in distributed software. ACM SIGSOFT Software Engineering Notes, 22(6): 52–59, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ketchum:1990:CCQ

[Ket90] D. W. Ketchum. Correctness is not congruent with quality. ACM SIGSOFT Software Engineering Notes, 15(5):18, October 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishnamurthi:1998:TFT

[KF98] Shriram Krishnamurthi and Matthias Felleisen. Toward a formal theory of extensible software. ACM SIGSOFT Software Engineering Notes, 23(6): 88–98, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kontio:1998:EIR

[KGL98] Jyrki Kontio, Gerhard Getto, and Dieter Landes. Experiences in improving risk management processes using the concepts of the riskit method. ACM SIGSOFT Software Engineering Notes, 23(6):163–174, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kiczales:1996:WCP

[Kic96] Gregor Kiczales. What can programming languages contribute to software engineering, and vice versa? (panel). ACM SIGSOFT Software Engineering Notes, 21(6):2, November 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kilov:1991:CCE

[Kil91]  Haim Kilov.  Conventional

and convenient in entityrelationship modeling. ACM SIGSOFT Software Engineering Notes, 16(2):31–32, April 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kilov:1994:FMS

[Kil94a] Haim Kilov. Formal methods and standards. ACM SIGSOFT Software Engineering Notes, 19(3):40, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kilov:1994:UHL

[Kil94b] Haim Kilov. On understanding hypertext: are links essential? ACM SIGSOFT Software Engineering Notes, 19(1):30, January 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kilov:1999:BSW

[Kil99] Haim Kilov. Behavioral Semantics Workshop (call for papers).

ACM SIGSOFT Software Engineering Notes, 24(4):3, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kim:1997:RBA

[Kim97] Hyoseob Kim. Report on Bruce Anderson’s tutorial on pattern languages of program design. ACM SIGSOFT Software Engineering Notes, 22(4):34, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kim:1998:TRO

[Kim98] Hyoseob Kim. Trip report to the one-day seminar on componentware development. ACM SIGSOFT Software Engineering Notes, 23(5):50–51, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kirner:1991:RTS

[Kir91] Tereza G. Kirner. Real-time systems specification: a process and a tool. ACM SIGSOFT Software Engineering Notes, 16 (1):62–67, January 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1996:ESEb

[Kit96a] Barbara Ann Kitchenham.

Evaluating software engineering methods and tool — part 2: selecting an appropriate evaluation method — technical criteria. ACM SIGSOFT Software Engineering Notes, 21 (2):11–15, March 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1996:ESEc

[Kit96b] Barbara Ann Kitchenham. Evaluating software engineering methods and tool — part 3: selecting an appropriate evaluation method — practical issues. ACM SIGSOFT Software Engineering Notes, 21(4):9–12, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1996:ESEa

[Kit96c] Barbara Ann Kitchenham. Evaluating software engineering methods and tool part 1: The evaluation context and evaluation methods. ACM SIGSOFT Software Engineering Notes, 21(1):11–14, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1997:ESEc

[Kit97] Barbara Ann Kitchenham. Evaluating software engineering methods and tools, part 7: planning feature analysis evaluation. ACM SIGSOFT Software Engineering Notes, 22 (4):21–24, July 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1998:ESEd

[Kit98] Barbara Ann Kitchenham.

Evaluating software engineering methods and tools part 12: evaluating DESMET. ACM SIGSOFT Software Engineering Notes, 23(5):21–24, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1997:ESEa

[KJ97a] Barbara Ann Kitchenham and Lindsay Jones. Evaluating software engineering methods and tool part 5: the influence of human factors. ACM SIGSOFT Software Engineering Notes, 22 (1):13–15, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1997:ESEb

[KJ97b] Barbara Ann Kitchenham and Lindsay Jones. Evaluating software engineering methods and tools part 6: identifying and scoring features. ACM SIGSOFT Software Engineering Notes, 22(2):16–18, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1997:ESEd

[KJ97c] Barbara Ann Kitchenham and Lindsay Jones. Evaluating SW eng. methods and tools, part 8: analysing a feature analysis evaluation. ACM SIGSOFT Software Engineering Notes, 22(5):10–12, September 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kang:1998:PFB

[KKLK98] Kyo C. Kang, Gerard J. Kim, Ji Y. Lee, and Hye J. Kim. Prototype = function + behavior + form. ACM SIGSOFT Software Engineering Notes, 23 (4):44–49, July 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Knight:1990:RCK

[KL90] John C. Knight and Nancy G. Leveson. A reply to the criticisms of the Knight & Leveson experiment. ACM SIGSOFT Software Engineering Notes, 15 (1):24–35, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kang:1996:ESS

[KL96] Inhye Kang and Insup Lee. An efficient state space generation for analysis of real-time systems. ACM SIGSOFT Software Engineering Notes, 21(3):4–13, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Keller:1999:IWL

[KLS99] Rudolf K. Keller, Bruno Laguë, and Reinhard Schauer. International Workshop on Large-Scale Software Composition.

ACM SIGSOFT Software Engineering Notes, 24(1):49– 54, January 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Knight:1991:PIT

[KM91] John C. Knight and E. Ann Myers. Phased inspections and their implementation. ACM SIGSOFT Software Engineering Notes, 16(3):29–35, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokai:1999:GGP

[KNN99] Gabriella Kókai, Jörg Nilson, and Christian Niss. GIDTS: a graphical programming environment for Prolog. ACM SIG-SOFT Software Engineering Notes, 24(5):95–104, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kobialka:1990:CEG

[Kob90] Hans-Ulrich Kobialka. Configuration editing, generation and test within working contexts. ACM SIGSOFT Software Engineering Notes, 15(6): 173–182, December 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kogut:1995:DRC

[Kog95] Paul Kogut. Design reuse:

chemical engineering vs. software engineering. ACM SIG-SOFT Software Engineering Notes, 20(5):73–77, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1993:MHW

[Kok93] Peter Kokol. Metamodeling: how, why and what? ACM SIGSOFT Software Engineering Notes, 18(2):25–26, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1996:TQO

[Kok96] Peter Kokol. Toward quality oriented IS design processes using two levelled design. ACM SIGSOFT Software Engineering Notes, 21(2):53–55, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1999:MFS

[Kok99]  Peter Kokol.  Measuring

formal specification with αmetric. ACM SIGSOFT Software Engineering Notes, 24(1): 80–81, January 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kolski:1998:CAA

[Kol98] Christophe Kolski. A “call for answers” around the proposition of an HCI-enriched model. ACM SIGSOFT Software Engineering Notes, 23(3):93–96, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Korel:1996:ATD

[Kor96] Bogdan Korel. Automated test data generation for programs with procedures. ACM SIGSOFT Software Engineering Notes, 21(3):209–215, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kostecki:1995:BRO

[Kos95] John A. Kostecki. Book review: Object-Oriented Software Metrics by Mark Lorenz and Jeff Kidd. ACM SIGSOFT Software Engineering Notes, 20 (1):91–93, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kostecki:1998:BRC

[Kos98] John A. Kostecki. Book reviews: The Craft of Software Testing: Subsystem Testing Including Object-Based and Object-Oriented Testing, by Brian Maxick. ACM SIGSOFT Software Engineering Notes, 23 (3):129–130, May 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1998:ESEb

[KP98a] Barbara Ann Kitchenham and Lesley M. Pickard. Evaluating software eng. methods and tools part 10: designing and running a quantitative case study. ACM SIGSOFT Software Engineering Notes, 23 (3):20–22, May 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1998:ESEc

[KP98b] Barbara Ann Kitchenham and Lesley M. Pickard. Evaluating software engineering methods and tools, part 11: analysing quantitative case studies. ACM SIGSOFT Software Engineering Notes, 23(4):18–20, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kitchenham:1998:ESEa

[KP98c] Barbara Ann Kitchenham and Lesley M. Pickard. Evaluating software engineering methods and tools: part 9: quantitative case study methodology. ACM SIGSOFT Software Engineering Notes, 23(1):24–26, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kirner:1999:EUC

[KPP^{+}99] D. Kirner, R. Porter, P. Punniamoorthy, M. Schuh, D. Shoup, S. Tindall, and D. Umphress. Extending use cases throughout the software lifecycle. ACM SIGSOFT Software Engineering Notes, 24(3):66–68, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krasner:1997:CRS

[Kra97] Herb Krasner. Clarifying the role of the SEI: they do not certify assessors or assessments. ACM SIGSOFT Software Engineering Notes, 22(4):7, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishnamurthy:1995:CWE

[Kri95] Balachander Krishnamurthy.

CSCW 94 Workshop to Explore Relationships between Research in Computer Supported Cooperative Work and Software Process: Workshop report. ACM SIGSOFT Software Engineering Notes, 20 (2):34–35, April 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krogstie:1998:IUQ

[Kro98] John Krogstie. Integrating the understanding of quality in requirements specification and conceptual modeling. ACM SIGSOFT Software Engineering Notes, 23(1):86–91, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krasna:1998:HIQ

[KRS98] Marjan Krašna, Ivan Rozman, and Bruno Stiglic. How to improve the quality of software engineering project management. ACM SIGSOFT Software Engineering Notes, 23(3):

120–125, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Karhinen:1997:CDR

[KRT97] Anssi Karhinen, Alexander Ran, and Tapio Tallgren. Configuring designs for reuse. ACM SIGSOFT Software Engineering Notes, 22(3):199–208, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kurki-Suonio:1991:SDR

[KS91] Reino Kurki-Suonio. Stepwise design of real-time systems. ACM SIGSOFT Software Engineering Notes, 16(5): 120–131, December 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koelbel:1990:WRE

[KSL90] Chuck Koelbel, Gene Spafford, and George Leach. Workshop report: experiences with building distributed and multiprocessor systems. ACM SIGSOFT Software Engineering Notes, 15(2):39–41, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kokol:1991:SSM

[KSZ91] Peter Kokol, Bruno Stiglic, and Viljem Zumer. Soft system methodology and is research: development of a new is design paradigm evaluation approach. ACM SIGSOFT Software Engineering Notes, 16(4): 64–66, October 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kontogiannis:1994:REQ

[KT94] Kostas A. Kontogiannis and Scott R. Tilley. Reverse engineering questionnaire. ACM SIGSOFT Software Engineering Notes, 19(1):31–38, January 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koppol:1996:IAS

[KT96] Pramod V. Koppol and Kuo-Chung Tai. An incremental approach to structural testing of concurrent software. ACM SIGSOFT Software Engineering Notes, 21(3):14–23, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Karlsson:1997:IDA

[KT97] Even-André Karlsson and Lars Taxen. Incremental development for AXE 10. ACM SIG-SOFT Software Engineering Notes, 22(6):519–520, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaplan:1992:SCS

[KTC^{+}92] Simon M. Kaplan, William J.

Tolone, Alan M. Carroll, Douglas P. Bogia, and Celsina Bignoli. Supporting collaborative software development with ConversationBuilder. ACM SIGSOFT Software Engineering Notes, 17(5):11–20, December 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaplan:1995:FAU

[KW95] Alan Kaplan and Jack C. Wileden. Formalization and application of a unifying model for name management. ACM SIGSOFT Software Engineering Notes, 20(4):161–172, October 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kramer:1996:SIW

[KW96] Jeff Kramer and Alexander L. Wolf. Succeedings of the 8th International Workshop on Software Specification and Design. ACM SIGSOFT Software Engineering Notes, 21(5): 21–35, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kwon:1997:RS

[Kwo97] Oh Cheon Kwon. Report from SEKE’97. ACM SIGSOFT Software Engineering Notes, 22(5):15–16, September 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kwon:1998:RS

[Kwo98] Oh Cheon Kwon. Report from SEKE’97. ACM SIGSOFT Software Engineering Notes, 23 (1):27–28, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:1998:OSA

[LA98] Li Li and Dharma P. Agrawal.

Operating systems architecture future directions for heterogeneous communication systems. ACM SIGSOFT Software Engineering Notes, 23(4):35–37, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lai:1991:ATT

[Lai91] Robert Chi Tau Lai. Ada task taxonomy support for concurrent programming. ACM SIGSOFT Software Engineering Notes, 16(1):73–91, January 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Laitinen:1992:DCS

[Lai92] Kari Laitinen. Document classification for software quality systems. ACM SIGSOFT Software Engineering Notes, 17(4): 32–39, October 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Laitinen:1996:EUS

[Lai96] Kari Laitinen. Estimating understandability of software documents. ACM SIGSOFT Software Engineering Notes, 21(4): 81–92, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lam:1997:CRA

[Lam97a] W. Lam. Creating reusable architectures: initial experience report. ACM SIGSOFT Software Engineering Notes, 22 (4):39–43, July 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lam:1997:PRU

[Lam97b] W. Lam. Process reuse using a template approach: a case-study from avionics. ACM SIGSOFT Software Engineering Notes, 22(2):35–38, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lam:1998:VCR

[Lam98] W. Lam. Viewpoint-centred reuse: bridging the gap between reusability and the needs of the reuser. ACM SIGSOFT Software Engineering Notes, 23(1):100–103, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lang:1993:SMO

[Lan93]  Neil Lang.  Schlaer–Mellor

object-oriented analysis rules. ACM SIGSOFT Software Engineering Notes, 18(1):54– 58, January 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Laplante:1990:HU

[Lap90] P. Laplante. Heisenberg uncertainty. ACM SIGSOFT Software Engineering Notes, 15(5): 21–22, October 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:1997:ODT

[Law97] Rob Law. An overview of debugging tools. ACM SIGSOFT Software Engineering Notes, 22 (2):43–47, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:1998:GLO

[Law98a]  Rob Law.  Generalization

of leaner object-oriented slicing. ACM SIGSOFT Software Engineering Notes, 23(1): 59–61, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Law:1998:IOO

[Law98b] Rob Law. Interacting with object-oriented program slicing. ACM SIGSOFT Software Engineering Notes, 23(1): 66–67, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Laws:1998:MCS

[Law98c] John Laws. Management of complexity in software development. ACM SIGSOFT Software Engineering Notes, 23(4):

50–52, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

LeMetayer:1996:SAS

[Le 96] Daniel Le Métayer. Software architecture styles as graph grammars. ACM SIGSOFT Software Engineering Notes, 21 (6):15–23, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lea:1994:CAI

[Lea94] Doug Lea. Christopher Alexander: an introduction for objectoriented designers. ACM SIG-SOFT Software Engineering Notes, 19(1):39–46, January 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leciston:1996:BRP

[Lec96a] David John Leciston. Book review: PCTE: The Standard for Open Repositories by Lois Wakeman and Jonathan Jowett. ACM SIGSOFT Software Engineering Notes, 21(2): 92, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leciston:1996:LE

[Lec96b] David John Leciston. Letter to the editor. ACM SIGSOFT Software Engineering Notes, 21 (1):2–5, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leciston:1999:AFP

[Lec99] David John Leciston. ACM Fellow profile: Larry E. Druffel. ACM SIGSOFT Software Engineering Notes, 24(3):9–10, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leisner:1995:CUP

[Lei95] Marry Leisner. Confessions of a used program salesman: Institutionalizing software reuse. ACM SIGSOFT Software Engineering Notes, 20 (5):90, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leisner:1996:BRD

[Lei96a]  Marty Leisner.  Book re-

view: Debugging the Development Process. ACM SIGSOFT Software Engineering Notes, 21 (4):99–100, July 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leisner:1996:BRM

[Lei96b] Marty Leisner. Book review:

Managing Your Move to Object Technology: Guidelines and Strategies for a Smooth Transition. ACM SIGSOFT Software Engineering Notes, 21(4):100, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leisner:1998:BRS

[Lei98] Marty Leisner. Book reviews: Structuring Techniques: an Introduction Using C++, Andrew C. Staugaard, Jr. ACM SIGSOFT Software Engineering Notes, 23(3):126, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leveson:1992:FM

[Lev92] Nancy Leveson. Formal methods. ACM SIGSOFT Software Engineering Notes, 17(1): 49–50, January 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leach:1995:IDA

[LF95] Ronald J. Leach and Terrence L. Fuller. An illustration of the domain analysis process. ACM SIGSOFT Software Engineering Notes, 20(5): 78–82, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lindquist:1991:OVT

[LGRM91] Timothy E. Lindquist, Kurt M. Gutzmann, David L. Remkes, and Gary McKee. Optimization of validation test suite coverage. ACM SIGSOFT Software Engineering Notes, 16(3): 87–92, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liang:1999:EPA

[LH99a] Donglin Liang and Mary Jean Harrold. Efficient pointsto analysis for whole-program analysis. ACM SIGSOFT Software Engineering Notes, 24(6): 199–215, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liang:1999:EAG

[LH99b] Donglin Liang and Mary Jean Harrold. Equivalence analysis: a general technique to improve the efficiency of dataflow analyses in the presence of pointers. ACM SIGSOFT Software Engineering Notes, 24(5): 39–46, September 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leveson:1999:DSL

[LHR99] Nancy G. Leveson, Mats P. E.

Heimdahl, and Jon Damon Reese. Designing specification languages for process control systems: lessons learned and steps to the future. ACM SIG-SOFT Software Engineering Notes, 24(6):127–145, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Linos:1992:TRC

[Lin92] Panagiotis Linos. ToolCASE:

a repository of computer-aided software engineering tools. ACM SIGSOFT Software Engineering Notes, 17(2):74–78, April 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lam:1997:SDA

[LM97] W. Lam and J. A. McDermid. A summary of domain analysis experience by way of heuristics. ACM SIGSOFT Software Engineering Notes, 22(3):54–64, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Logrippo:1990:ASL

[LMD90] Luigi Logrippo, Tim Melanchuk, and Robert J. Du Wors. The algebraic specification language LOTOS: an industrial experience. ACM SIGSOFT Software Engineering Notes, 15(4): 59–66, September 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lutz:1999:RBP

[LN99] Michael J. Lutz and J. Fernando Naveda. RIT’s BS program in software engineering: a status report. ACM SIGSOFT Software Engineering Notes, 24 (3):53–55, May 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Leavens:1998:WFC

[LNS98] Gary T. Leavens, Oscar Nierstrasz, and Murali Sitaraman. 1997 Workshop on Foundations of Component-Based Systems. ACM SIGSOFT Software Engineering Notes, 23(1): 38–41, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loka:1992:SEQ

[Lok92] Raghavendra Rao Loka. Software engineering: quality assurance. ACM SIGSOFT Software Engineering Notes, 17(3): 34–38, July 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lott:1990:CCQ

[Lot90] C. M. Lott. Correctness is congruent with quality. ACM SIGSOFT Software Engineering Notes, 15(5):19–20, October 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lott:1993:PMS

[Lot93] Christopher M. Lott. Process and measurement support in SEEs. ACM SIGSOFT Software Engineering Notes, 18(4): 83–93, October 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loy:1990:COO

[Loy90]  Patrick H. Loy.  A com-

parison of object-oriented and structured development methods. ACM SIGSOFT Software Engineering Notes, 15(1): 44–48, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Loy:1993:MWS

[Loy93] Patrick Loy. The method won’t save you: (but it can help). ACM SIGSOFT Software Engineering Notes, 18(1): 30–34, January 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

LeBissonnais:1997:MDR

[LP97] Marc Le Bissonnais and François Prunet. From manufacturing document requirements to customized authoring and automated publishing framework. ACM SIGSOFT Software Engineering Notes, 22 (2):94–98, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ledru:1990:AVL

[LS90a] Yves Ledru and Pierre-Yves Schobbens. Applying VDM to large developments. ACM SIG-SOFT Software Engineering Notes, 15(4):55–58, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lockman:1990:PTT

[LS90b] Abe Lockman and John Salasin. A procedure and tools for transition engineering. ACM SIGSOFT Software Engineering Notes, 15(6):

157–172, December 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Levy:1994:LIA

[LS94] N. Lévy and G. Smith.

A language-independent approach to specification construction. ACM SIGSOFT Software Engineering Notes, 19 (5):76–86, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Land:1997:VDD

[LSJ97] Lesley Pek Wee Land, Chris Sauer, and Ross Jeffery. Validating the defect detection performance advantage of group designs for software reviews: report of a laboratory experiment using program code. ACM SIGSOFT Software Engineering Notes, 22(6):294–309, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Levine:1993:MDR

[LT93] David L. Levine and Richard N.

Taylor. Metric-driven reengineering for static concurrency analysis. ACM SIGSOFT Software Engineering Notes, 18(3): 40–50, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lung:1995:ACD

[LU95] Chung-Horng Lung and Joseph E.

Urban. An approach to the classification of domain models in support of analogical reuse. ACM SIGSOFT Software Engineering Notes, 20(SI):169– 178, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ludewig:1996:STI

[Lud96] Jochen Ludewig. Summary of the Third International Workshop on Software Engineering Education (IWSEE3). ACM SIGSOFT Software Engineering Notes, 21(4):47–51, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lutz:1993:TSR

[Lut93]  Robyn R. Lutz.  Target-

ing safety-related errors during software requirements analysis. ACM SIGSOFT Software Engineering Notes, 18(5): 99–106, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lippe:1992:OBM

[LvO92] Ernst Lippe and Norbert van Oosterom. Operation-based merging. ACM SIGSOFT Software Engineering Notes, 17(5): 78–87, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liao:1993:SRB

[LW93] Hsian-Chou Liao and Feng-Jian Wang. Software reuse based on a large object-oriented library. ACM SIGSOFT Software Engineering Notes, 18(1): 74–80, January 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lam:1996:TDS

[LW96a] Wing Lam and Ben Whittle. A taxonomy of domainspecific reuse problems and their resolutions — version 1.0. ACM SIGSOFT Software Engineering Notes, 21(5):72–77, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Latour:1996:WAW

[LW96b] Larry Latour and Kevin Wentzel. WISR’95: 7th Annual Workshop on Software Reuse summary and working group reports. ACM SIGSOFT Software Engineering Notes, 21(1): 35–54, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ludwig:1999:VEC

[LW99] Heiko Ludwig and Keith Whittingham. Virtual enterprise co-ordinator — agreementdriven gateways for crossorganisational workflow management. ACM SIGSOFT Software Engineering Notes, 24 (2):29–38, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:1999:GCN

[LWM99] Du Li, Zhenghao Wang, and Richard R. Muntz. “Got COCA?” a new perspective in building electronic meeting systems. ACM SIGSOFT Software Engineering Notes, 24(2): 89–98, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:1998:NSM

[LXK98] Jonathan Lee, Nien-Lin Xue, and Tai-Lin Kuei. A note on state modeling through inheritance. ACM SIGSOFT Software Engineering Notes, 23(1): 104–110, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lyu:1991:PRD

[Lyu91] Michael Lyu. PANEL: research and development issues in software reliability engineering. ACM SIGSOFT Software Engineering Notes, 16(2): 23–30, April 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:1998:AOM

[LZL98] Hong Liu, G. Zeng, and Zongkai Lin. An agent-oriented modeling approach. ACM SIG-SOFT Software Engineering Notes, 23(3):87–92, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:1999:DSS

[LZL99] Baojian Li, Guangzhou Zeng, and Zongkai Lin. A domain specific software architecture style for CSCD system. ACM SIGSOFT Software Engineering Notes, 24(1):59–64, January 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

MacAuslan:1991:CDH

[Mac91a] Joel MacAuslan. Comments on debugging hypercubes in a von Neumann language. ACM SIGSOFT Software Engineering Notes, 16(3):73, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maciaszek:1991:ACR

[Mac91b] Leszek A. Maciaszek. AD/ cycle repository manager from object-oriented perspective. ACM SIGSOFT Software Engineering Notes, 16(1):50– 53, January 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mahony:1998:FHS

[Mah98a] Michael S. Mahony. Finding a history for software engineering. ACM SIGSOFT Software Engineering Notes, 23 (6):87, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Molloy:1998:AIT

[MAH^{+}98b] Mark Molloy, Kristy Andrews, James Herren, David Cutler, and Paul Del Vigna. Automatic interoperability test generation for source-to-source translators. ACM SIGSOFT Software Engineering Notes, 23 (2):93–101, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maiden:1995:ROR

[Mai95] N. A. M. Maiden. Reuseoriented requirements engineering in NATURE. ACM SIG-SOFT Software Engineering Notes, 20(3):90–93, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maibaum:1997:WWT

[Mai97] Tom S. E. Maibaum. What we teach software engineers in the university: do we take engineering seriously? ACM SIGSOFT Software Engineering Notes, 22(6):40–50, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mili:1997:ANC

[MAKGM97] Hafedh Mili, Estelle Ah-Ki, Robert Godin, and Hamid Mcheick. Another nail to the coffin of faceted controlledvocabulary component classification and retrieval. ACM SIGSOFT Software Engineering Notes, 22(3):89–98, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Magnusson:1993:FGR

[MAM93] Boris Magnusson, Ulf Asklund, and Sten Minör. Fine-grained revision control for collaborative software development. ACM SIGSOFT Software Engineering Notes, 18(5):33–41, December 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mamone:1994:ISS

[Mam94] Salvatore Mamone. The IEEE standard for software maintenance. ACM SIGSOFT Software Engineering Notes, 19(1): 75–76, January 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Martinis:1990:SCC

[Mar90] J. Martinis. Softool change/ configuration management. ACM SIGSOFT Software Engineering Notes, 15(3):51, July 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marsh:1991:VPM

[Mar91] Shaun Marsh. The v project manager tools. ACM SIGSOFT Software Engineering Notes, 16 (2):58–61, April 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marovac:1994:GES

[Mar94] Nenad Marovac. Guidelines for embedded software documentation. ACM SIGSOFT Software Engineering Notes, 19 (2):22–28, April 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marovac:1997:SDE

[Mar97] Nenad Marovac. Software development environment based on HyperNet. ACM SIGSOFT Software Engineering Notes, 22 (2):68–71, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marovac:1998:EDS

[Mar98a]  Nenad Marovac.  Embed-

ded documentation for semiautomatic program construction and software reuse. ACM SIGSOFT Software Engineering Notes, 23(3):70–74, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Martin:1998:PDU

[Mar98b] Declan Martin. Preliminary design use cases: combining use cases and event response lists for reuse and legacy system enhancement. ACM SIGSOFT Software Engineering Notes, 23 (1):68–71, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Matthews:1992:WWB

[Mat92] Peter J. D. Matthews. When to white box test. ACM SIGSOFT Software Engineering Notes, 17 (1):43, January 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mathalone:1997:BBM

[Mat97] Sigal Mathalone. A behaviorallybased methodology for modeling system specifications. ACM SIGSOFT Software Engineering Notes, 22(2):39–42, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marre:1996:UDT

[MB96] Martina Marré and Antonia Bertolino. Unconstrained duals and their use in achieving alluses coverage. ACM SIGSOFT Software Engineering Notes, 21 (3):147–157, May 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McLaughlin:1991:SNP

[McL91] Robert McLaughlin. Some notes on program design. ACM SIGSOFT Software Engineering Notes, 16(4):53–54, October 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McLaughlin:1992:SNS

[McL92] Robert McLaughlin. Some notes on software design: reply to a reaction. ACM SIGSOFT Software Engineering Notes, 17 (2):70, April 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [vR92].

McLaughlin:1993:DCM

[McL93]  Robert McLaughlin.  Does

CASE make the customer happier. ACM SIGSOFT Software Engineering Notes, 18(2):18, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McGregor:1993:NIS

[MD93] John D. McGregor and Douglas M. Dyer. A note on inheritance and state machines. ACM SIGSOFT Software Engineering Notes, 18(4):61–69, October 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meijler:1997:MDP

[MDE97] Theo Dirk Meijler, Serge Demeyer, and Robert Engel. Making design patterns explicit in FACE: a frame work adaptive composition environment. ACM SIGSOFT Software Engineering Notes, 22(6):94–110, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mester:1998:AFP

[Mes98] Arnulf Mester. ACM Fellow profile: James Jay (Jim) Horning. ACM SIGSOFT Software Engineering Notes, 23(4):7–8, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marx:1996:PWA

[MF96] Delia I. S. Marx and Phyllis G. Frankl. The path-wise approach to data flow testing with pointer variables. ACM SIGSOFT Software Engineering Notes, 21(3):135–146, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mambella:1995:IAS

[MFCS95] E. Mambella, R. Ferrari, F. D.

Carli, and A. L. Surdo. An integrated approach to software reuse practice. ACM SIGSOFT Software Engineering Notes, 20(SI):63–80, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mili:1997:CPM

[MFDM97] R. Mili, M. Frappier, J. Desharnais, and A. Mili. A calculus of program modifications. ACM SIGSOFT Software Engineering Notes, 22(3):157–168, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mashayekhi:1994:CCA

[MFR94] Vahid Mashayekhi, Chris Feulner, and John Riedl. CAIS: collaborative asynchronous inspection of software. ACM SIGSOFT Software Engineering Notes, 19(5):21–34, December 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meekel:1997:DMA

[MHF^{+}97] Jacques Meekel, Thomas B. Horton, Robert B. France, Charlie Mellone, and Sajid Dalvi. From domain models to architecture frameworks. ACM SIGSOFT Software Engineering Notes, 22(3):75–80, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Musa:1991:ETN

[MI91] John D. Musa and A. Iannino. Estimating the total number of software failures using an exponential model. ACM SIGSOFT Software Engineering Notes, 16 (3):80–84, July 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miller:1993:EST

[Mil93] Edward F. Miller. Exploitation of software test technology. ACM SIGSOFT Software Engineering Notes, 18(3):159, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maruyama:1997:MAD

[MiS97] Katsuhisa Maruyama and Ken ichi Shima. A mechanism for automatically and dynamically changing software components. ACM SIGSOFT Software Engineering Notes, 22(3):169–180, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mannion:1995:SR

[MK95] Mike Mannion and Barry Keepence. SMART requirements. ACM SIGSOFT Software Engineering Notes, 20 (2):42–47, April 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Magee:1996:DSS

[MK96] Jeff Magee and Jeff Kramer.

Dynamic structure in software architectures. ACM SIGSOFT Software Engineering Notes, 21 (6):3–14, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mannion:1998:PR

[MK98] Mike Mannion and Barry Keepence. Partnering for reuse. ACM SIGSOFT Software Engineering Notes, 23(1): 77–80, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mahler:1990:ICM

[ML90] Axel Mahler and Andreas Lampen. Integrating configuration management into a generic environment. ACM SIGSOFT Software Engineering Notes, 15 (6):229–237, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maeda:1997:OIA

[MLMK97] Chris Maeda, Arthur Lee, Gail Murphy, and Gregor Kiczales. Open implementation analysis and design. ACM SIGSOFT Software Engineering Notes, 22 (3):44–52, May 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morell:1996:UPA

[MM96] Larry Morell and Branson Murrill. Using perturbation analysis to measure variation in the information content of test sets. ACM SIGSOFT Software Engineering Notes, 21(3):92–97, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morasca:1996:GFT

[MMS96] Sandro Morasca, Angelo Morzenti, and Pieluigi SanPietro. Generating functional test cases inthe-large for time-critical systems from logic-based specifications. ACM SIGSOFT Software Engineering Notes, 21(3): 39–52, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Murphy:1995:LSM

[MN95] Gail C. Murphy and David Notkin. Lightweight source model extraction. ACM SIG-SOFT Software Engineering Notes, 20(4):116–127, October 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Murphy:1995:SRM

[MNS95] Gail C. Murphy, David Notkin, and Kevin Sullivan. Software reflexion models: bridging the gap between source and highlevel models. ACM SIGSOFT Software Engineering Notes, 20 (4):18–28, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mody:1992:PA

[Mod92] R. P. Mody. Is programming an art? ACM SIGSOFT Software Engineering Notes, 17(4): 19–21, October 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Modes:1993:SIV

[Mod93] Ronald W. Modes. Structured IV & V for the space shuttle flight software. ACM SIGSOFT Software Engineering Notes, 18(3):81, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Montangero:1997:RFE

[Mon97] Carlo Montangero. Report on the Fifth European Workshop on Software Process Technology. ACM SIGSOFT Software Engineering Notes, 22(1): 24–27, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Medvidovic:1996:UOO

[MORT96] Nenad Medvidovic, Peyman Oreizy, Jason E. Robbins, and Richard N. Taylor. Using object-oriented typing to support architectural design in the C2 style. ACM SIGSOFT Software Engineering Notes, 21(6): 24–32, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Medvidovic:1997:RSC

[MOT97] Nenad Medvidovic, Peyman Oreizy, and Richard N. Taylor. Reuse of off-the-shelf components in c2-style architectures. ACM SIGSOFT Software Engineering Notes, 22(3):190–198, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mascolo:1999:FGM

[MPR99] Cecilia Mascolo, Gian Pietro Picco, and Gruia-Catalin Roman. A fine-grained model for code mobility. ACM SIGSOFT Software Engineering Notes, 24 (6):39–56, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Moriconi:1994:CCS

[MQ94] Mark Moriconi and Xiaolei Qian. Correctness and composition of software architectures. ACM SIGSOFT Software Engineering Notes, 19(5): 164–174, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Minsky:1990:CMC

[MR90] Naftaly H. Minsky and David Rozenshtein. Configuration management by consensus: an application of law-governed systems. ACM SIGSOFT Software Engineering Notes, 15(6): 44–55, December 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meyers:1992:ESM

[MR92] Scott Meyers and Steven P.

Reiss. An empirical study of multiple-view software development. ACM SIGSOFT Software Engineering Notes, 17(5): 47–57, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mrdalj:1990:BOO

[Mrd90] S. Mrdalj. Biblio of objectoriented system development. ACM SIGSOFT Software Engineering Notes, 15(5):60– 63, October 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Muller:1998:PCS

[MRS98] Hausi Müller, Thomas Reps, and Gregor Snelting. Program comprehension and software reengineering. ACM SIGSOFT Software Engineering Notes, 23 (5):36–44, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mazhindu-Shumba:1996:CNW

[MS96] Rose Mazhindu-Shumba. A comment on the notation of the Wirfs–Brock et al. objectoriented design method. ACM SIGSOFT Software Engineering Notes, 21(5):99, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McCoog:1997:MRA

[MS97a] Phillip McCoog and Rick Smith. Mostly reuse: another code sharing option. ACM SIGSOFT Software Engineering Notes, 22(3):39–43, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morzenti:1997:ATA

[MS97b] Angelo Morzenti and Fabio A. Schreiber. Analysis of techniques in the assessment of ultra-reliable computerized systems. ACM SIGSOFT Software Engineering Notes, 22(2): 87–91, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mili:1997:RQR

[MSB97] Hafedh Mili, Houari Sahraoui, and Ilham Benyahia. Representing and querying reusable object frameworks. ACM SIGSOFT Software Engineering Notes, 22(3):110–120, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Masuda:1998:ADP

[MSU98] Gou Masuda, Norihiro Sakamoto, and Kazuo Ushijima. Applying design patterns to decision tree learning system. ACM SIG-SOFT Software Engineering Notes, 23(6):111–120, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Medvidovic:1997:FCC

[MT97] Nenad Medvidovic and Richard N.

Taylor. A framework for classifying and comparing architecture description languages. ACM SIGSOFT Software Engineering Notes, 22(6):60–76, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Muller:1992:REE

[MTO^{+}92] H. A. Müller, S. R. Tilley, M. A. Orgun, B. D. Corrie, and N. H. Madhavji. A reverse engineering environment based on spatial and visual software interconnection models. ACM SIGSOFT Software Engineering Notes, 17(5):88–98, December 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mugisa:1997:RTS

[Mug97] Ezra K. Mugisa. A reuse triplet for systematic software reuse. ACM SIGSOFT Software Engineering Notes, 22(4):65–69, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Muhlhauser:1996:MWM

[Müh96] Max Mühlhäuser. MMSD’96: Workshop on Multimedia Software Development. ACM SIGSOFT Software Engineering Notes, 21(4):58–63, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Musa:1991:PAR

[Mus91a] John D. Musa. Performance analysis rat holes to avoid or how to stall a performance presentation. ACM SIGSOFT Software Engineering Notes, 16 (3):78, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Musa:1991:RFE

[Mus91b] John D. Musa. Rationale for fault exposure ratio K. ACM SIGSOFT Software Engineering Notes, 16(3):79, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morell:1991:IDS

[MV91] Larry J. Morell and Jeffrey Voas. On the inadequacies of data state space sampling as a measure of the trustworthiness of programs. ACM SIGSOFT Software Engineering Notes, 16 (2):73–74, April 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

McManus:1996:PMK

[MV96] Joe McManus and Renaat Vergruggen. A proposed methodology for knowledge based systems development. ACM SIGSOFT Software Engineering Notes, 21(2):22–31, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wei:1991:MCU

[mWE91] Yin min Wei and Klaus Eldridge. Module completeness as a useful guideline for programming. ACM SIGSOFT Software Engineering Notes, 16 (1):35, January 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mitchell:1998:EER

[MZ98] Brian Mitchell and Steven J.

Zeil. An experiment in estimating reliability growth under both representative and directed testing. ACM SIGSOFT Software Engineering Notes, 23 (2):32–41, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:1998:CDF

[NA98] Gleb Naumovich and George S. Avrunin. A conservative data flow algorithm for detecting all pairs of statements that may happen in parallel. ACM SIGSOFT Software Engineering Notes, 23(6):24–34, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:1999:EAC

[NAC99] Gleb Naumovich, George S. Avrunin, and Lori A. Clarke. An efficient algorithm for computing MHP information for concurrent Java programs. ACM SIGSOFT Software Engineering Notes, 24(6):338– 354, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:1997:ASA

[NACO97] Gleb Naumovich, George S. Avrunin, Lori A. Clarke, and Leon J. Osterweil. Applying static analysis to software architectures. ACM SIGSOFT Software Engineering Notes, 22 (6):77–93, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:1999:UPO

[NCC99] Gleb Naumovich, Lori A. Clarke, and Jamieson M. Cobleigh. Using partial order techniques to improve performance of data flow analysis based verification. ACM SIG-SOFT Software Engineering Notes, 24(5):57–65, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naumovich:1996:VCP

[NCO96] Gleb N. Naumovich, Lori A.

Clarke, and Leon J. Osterweil. Verification of communication protocols using data flow analysis. ACM SIGSOFT Software Engineering Notes, 21(6): 93–105, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Narayana:1990:IPD

[ND90] K. T. Narayana and Sanjeev Dharap. Invariant properties in a dialog system. ACM SIG-SOFT Software Engineering Notes, 15(4):67–79, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ness:1990:PET

[Nes90] Linda Ness. L.0: a parallel executable temporal logic language. ACM SIGSOFT Software Engineering Notes, 15(4): 80–89, September 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1990:RP

[Neu90a] P. G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 15 (5):3–17, October 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1990:RPCa

[Neu90b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 15 (1):4–23, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1990:RPCb

[Neu90c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 15 (2):3–22, April 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1991:RPCa

[Neu91a] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 16 (2):4–18, April 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1991:RPCb

[Neu91b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 16 (3):3–28, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1992:IRP

[Neu92] Peter G. Neumann. Illustrative risks to the public in the use of computer systems and related technology. ACM SIGSOFT Software Engineering Notes, 17 (1):23–32, January 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1993:RPC

[Neu93] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 18 (2):4–17, April 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1994:IRP

[Neu94] Peter G. Neumann. Illustrative risks to the public in the use of computer systems and related technology. ACM SIGSOFT Software Engineering Notes, 19 (1):16–29, January 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1995:FDC

[Neu95a] Peter Neumann. Fatal defect: Chasing killer computer bugs. ACM SIGSOFT Software Engineering Notes, 20(5): 91, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1995:SSS

[Neu95b] Peter Neumann. Safeware: System safety and computers. ACM SIGSOFT Software Engineering Notes, 20(5): 90–91, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1995:BRS

[Neu95c] Peter G. Neumann. Book review: Software Creativity by Robert L. Glass. ACM SIGSOFT Software Engineering Notes, 20(1):93, January 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1995:RPCa

[Neu95d] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 20 (1):16–21, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1995:RPCb

[Neu95e] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 20 (2):7–13, April 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1995:RPCc

[Neu95f] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 20 (3):7–12, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1995:RPCd

[Neu95g] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 20 (5):8–14, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1996:IRP

[Neu96a] Peter G. Neumann. Illustrative risks to the public in the use of computer systems and related technology. ACM SIGSOFT Software Engineering Notes, 21 (1):16–30, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1996:RPCa

[Neu96b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 21 (2):16–22, March 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1996:RPCb

[Neu96c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 21 (4):12–18, July 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1996:RPCc

[Neu96d] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 21(5):13–20, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1997:RPCa

[Neu97a] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 22 (1):16–22, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1997:RPCb

[Neu97b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 22 (2):19–24, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1997:RPCc

[Neu97c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 22 (4):25–34, July 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1997:RPCd

[Neu97d] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 22(5):13–15, September 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1998:RPCa

[Neu98a] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 23 (1):9–15, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1998:RPCb

[Neu98b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 23 (3):23–29, May 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1998:RPCc

[Neu98c] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 23 (4):21–25, July 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1998:RPCd

[Neu98d] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 23 (5):25–29, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1999:RPCa

[Neu99a] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 24 (1):31–35, January 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:1999:RPCb

[Neu99b] Peter G. Neumann. Risks to the public in computers and related systems. ACM SIGSOFT Software Engineering Notes, 24 (4):26–29, July 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nord:1990:FMM

[NLS90] Robert L. Nord, Peter Lee, and William L. Scherlis. Formal manipulation of modular software systems. ACM SIGSOFT Software Engineering Notes, 15(4):90–99, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nord:1997:SSA

[Nor97]  Robert L. Nord.  System

and software architecture track third IEEE international conference on engineering of complex computer systems (ICECCS’97). ACM SIGSOFT Software Engineering Notes, 22 (2):5, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:1993:NHR

[Not93] David Notkin. Nico Habermann remembered. ACM SIGSOFT Software Engineering Notes, 18(4):1, October 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:1996:LEC

[Not96] David Notkin. Letter from the executive committee. ACM SIGSOFT Software Engineering Notes, 21(1):1, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:1997:FSS

[Not97] David Notkin. FSE 96 special session: novel notions, wild ideas, and fun flames. ACM SIGSOFT Software Engineering Notes, 22(1):23–24, January 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nourani:1996:MAO

[Nou96] Cyrus F. Nourani. Multi-agent object level AI validation and verification. ACM SIGSOFT Software Engineering Notes, 21 (1):70–72, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nourani:1998:SAI

[Nou98] Cyrus F. Nourani. Software agents and intelligent object fusion. ACM SIGSOFT Software Engineering Notes, 23(1): 98, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ntafos:1998:RPT

[Nta98] Simeon Ntafos. On random and partition testing. ACM SIGSOFT Software Engineering Notes, 23(2):42–48, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ostrand:1998:VTD

[OAFG98] Thomas Ostrand, Aaron Anodide, Herbert Foster, and Tarak Goradia. A visual test development environment for GUI systems. ACM SIGSOFT Software Engineering Notes, 23 (2):82–92, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Offutt:1990:CT

[Off90] A. Jefferson Offutt. Comments on tai. ACM SIGSOFT Software Engineering Notes, 15 (1):67, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ossher:1990:SCR

[OH90] Harold Ossher and William Harrison. Support for change in RPDE3. ACM SIGSOFT Software Engineering Notes, 15(6): 218–228, December 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Offutt:1996:SMP

[OH96] A. Jefferson Offutt and J. Huffman Hayes. A semantic model of program faults. ACM SIGSOFT Software Engineering Notes, 21(3):195–200, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

OLaughlin:1996:BRB

[O’L96a] Brian O’Laughlin. Book review: Beyond Technology’s Promise. ACM SIGSOFT Software Engineering Notes, 21(4): 100–101, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

OLaughlin:1996:BRC

[O’L96b] Brian O’Laughlin. Book review: C/C++ Software Quality Tools. ACM SIGSOFT Software Engineering Notes, 21 (4):101–103, July 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

OLaughlin:1996:BRF

[O’L96c] Brian O’Laughlin. Book review: Foundations of Software Measurement. ACM SIGSOFT Software Engineering Notes, 21 (4):103, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

OLaughlin:1997:BRD

[O’L97a] Brian O’Laughlin. Book reviews: Dictionary of Object Technology, Donald G. Firesmith and Edward M. Eykholt. ACM SIGSOFT Software Engineering Notes, 22(1):107– 108, January 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

OLaughlin:1997:BRO

[O’L97b]  Brian O’Laughlin.  Book

reviews: Object Oriented Compiler Construction, Jim Holmes. ACM SIGSOFT Software Engineering Notes, 22 (1):108, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Olaughlin:1998:BRY

[Ola98] Brian Olaughlin. Book reviews: The Year 2000 Software Crisis: Challenge of the Century, William M. Ulrich and Ian S. Hayes. ACM SIGSOFT Software Engineering Notes, 23 (1):109, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Olenfeldt:1993:WSC

[Ole93] Lars Olenfeldt. Wish-4: a specification for a CASE-tool to wish for. ACM SIGSOFT Software Engineering Notes, 18 (2):20–22, April 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Omar:1991:SSF

[OM91] A. A. Omar and F. A. Mohammed. A survey of software functional testing methods. ACM SIGSOFT Software Engineering Notes, 16(2): 75–82, April 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ozcan:1995:VRV

[OM95] M. B. Ozcan and I. Morrey. A visual requirements validation environment for the reverse engineering of formal specifications from rapid prototypes. ACM SIGSOFT Software Engineering Notes, 20(5): 83–87, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

ONeill:1997:SVA

[O’N97] Don O’Neill. Software value added study. ACM SIGSOFT Software Engineering Notes, 22 (4):11–12, July 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Opdahl:1998:WSR

[OP98] Andreas L. Opdahl and Klaus Pohl. Workshop summary: REFSQ’98. ACM SIGSOFT Software Engineering Notes, 23 (5):44–50, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ogden:1994:PRF

[OSWZ94] William F. Ogden, Murali Sitaraman, Bruce W. Weide, and Stuart H. Zweben. Part i: the RESOLVE framework and discipline: a research synopsis. ACM SIGSOFT Software Engineering Notes, 19(4): 23–28, October 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Overmyer:1990:DSM

[Ove90]  S. P. Overmyer.  DoD-

Std-2167A and methodologies. ACM SIGSOFT Software Engineering Notes, 15(5): 50–59, October 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pomakis:1996:RAF

[PA96] Keith P. Pomakis and Joanne M. Atlee. Reachability analysis of feature interactions: a progress report. ACM SIGSOFT Software Engineering Notes, 21 (3):216–223, May 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Padberg:1999:PMS

[Pad99] Frank Padberg. A probabilistic model for software projects. ACM SIGSOFT Software Engineering Notes, 24(6):109–126, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1995:IMI

[Par95]  David Lorge Parnas.  On

ICSE’s “most influential” papers. ACM SIGSOFT Software Engineering Notes, 20(3): 29–32, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1997:SEE

[Par97] David Lorge Parnas. Software engineering (extended abstract): an unconsummated marriage. ACM SIGSOFT Software Engineering Notes, 22 (6):1–3, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1998:SSE

[Par98a] David Lorge Parnas. Successful software engineering research. ACM SIGSOFT Software Engineering Notes, 23(3):64–68, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1998:WTM

[Par98b] David Lorge Parnas. Who taught me about software engineering research? ACM SIGSOFT Software Engineering Notes, 23(4):26–28, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1999:CR

[Par99a] David Lorge Parnas. On code reuse. ACM SIGSOFT Software Engineering Notes, 24 (3):4, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See response [GMR99].

Parnas:1999:PPL

[Par99b] David Lorge Parnas. Parnas on parnas: a life of indecision. ACM SIGSOFT Software Engineering Notes, 24(4):47–49, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ponder:1994:PCH

[PB94] Carl Ponder and Bill Bush. Polymorphism considered harmful. ACM SIGSOFT Software Engineering Notes, 19 (2):35–37, April 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poulin:1999:WIS

[PBL99] Jeffrey S. Poulin, Don Batory, and Larry Latour. 9th Workshop on Institutionalizing Software Reuse (WISR ’9) workshop summary. ACM SIGSOFT Software Engineering Notes, 24(3):29–31, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prieto-Diaz:1990:DAI

[PD90] Rubén Prieto-Díaz. Domain analysis: an introduction. ACM SIGSOFT Software Engineering Notes, 15(2):47–54, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prieto-Diaz:1991:MSR

[PD91] Rubén Prieto-Díaz. Making software reuse work: an implementation model. ACM SIGSOFT Software Engineering Notes, 16(3):61–68, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prieto-Diaz:1995:SRS

[PD95] Rubén Prieto-Díaz. Systematic reuse: a scientific or an engineering method? ACM SIGSOFT Software Engineering Notes, 20(SI):9–10, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poulin:1998:ICS

[PD98] Jeffrey S. Poulin and Prem Devanbu. 5th international conference on software reuse (ICSR’5) conference summary. ACM SIGSOFT Software Engineering Notes, 23(5):30–35, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Person:1996:BRO

[Per96] Suzette Person. Book review: The Object Primer. ACM SIGSOFT Software Engineering Notes, 21(4):104, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Person:1997:BRW

[Per97] Suzette Person. Book reviews: What Every Software Manager MUST KNOW TO SUCCEED With Object Technology, John Williams. ACM SIGSOFT Software Engineering Notes, 22 (1):109, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Peterson:1991:CTS

[Pet91] A. Spencer Peterson. Coming to terms with software reuse terminology: a model-based approach. ACM SIGSOFT Software Engineering Notes, 16 (2):45–51, April 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Petroski:1996:EBC

[Pet96] Henry Petroski. Engineering bridges: from concept to reality (abstract). ACM SIGSOFT Software Engineering Notes, 21 (6):1, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Paredes:1995:RRS

[PF95] Carlos Paredes and José Luiz Fiadeiro. Reuse of requirements and specifications: a formal framework. ACM SIG-SOFT Software Engineering Notes, 20(SI):263–266, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfleeger:1994:DAS

[Pfl94] Shari Lawrence Pfleeger. Design and analysis in software engineering: the language of case studies and formal experiments. ACM SIGSOFT Software Engineering Notes, 19(4): 16–20, October 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfleeger:1995:EDAc

[Pfl95a] Shari Lawrence Pfleeger. Experimental design and analysis in software engineering, part 4: choosing an experimental design. ACM SIGSOFT Software Engineering Notes, 20(3): 13–15, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfleeger:1995:EDAd

[Pfl95b] Shari Lawrence Pfleeger. Experimental design and analysis in software engineering, part 5: analyzing the data. ACM SIGSOFT Software Engineering Notes, 20(5):14–17, December 1995. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfleeger:1995:EDAb

[Pfl95c] Shari Lawrence Pfleeger. Experimental design and analysis in software engineering: Types of experimental design. ACM SIGSOFT Software Engineering Notes, 20(2):14–16, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfleeger:1995:EDAa

[Pfl95d] Shari Lawrence Pfleeger. Experimental design and analysis in software engineering:Part 2: how to set up and experiment. ACM SIGSOFT Software Engineering Notes, 20(1): 22–26, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pillai:1996:FMI

[Pil96] Krish Pillai. The fountain model and its impact on project schedule. ACM SIGSOFT Software Engineering Notes, 21(2): 32–38, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Putkonen:1997:CTS

[PK97] A. Putkonen and M. Kiekara. A case-tool for supporting navigation in the class hierarchy. ACM SIGSOFT Software Engineering Notes, 22(1): 77–84, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Platek:1990:FMM

[Pla90] Richard Platek. Formal methods in mathematics. ACM SIG-SOFT Software Engineering Notes, 15(4):100–103, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pitts:1996:WWR

[PM96] David R. Pitts and Barbara H.

Miller. The wild-west revisited. ACM SIGSOFT Software Engineering Notes, 21(2): 84–86, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parissis:1996:SBT

[PO96] Ioannis Parissis and Farid Ouabdesselam. Specificationbased testing of synchronous software. ACM SIGSOFT Software Engineering Notes, 21(6): 127–134, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Polack:1990:PAC

[Pol90] Alexander J. Polack. Practical applications of CASE tools on DoD projects. ACM SIGSOFT Software Engineering Notes, 15 (1):73–78, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poore:1996:HMP

[Poo96] Jesse H. Poore. Harlan Mills’ passing. ACM SIGSOFT Software Engineering Notes, 21 (2):6, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poulin:1993:ISS

[Pou93] Jeffrey S. Poulin. Integrated support for software reuse in computer-aided software engineering (CASE). ACM SIG-SOFT Software Engineering Notes, 18(4):75–82, October 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poulin:1995:BRO

[Pou95] Jeffrey S. Poulin. Book reviews: Object Databases — The Essentials by Mary E. S. Loomis. ACM SIGSOFT Software Engineering Notes, 20(1): 90–91, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poulin:1997:SSR

[Pou97] Jeff Poulin. Symposium on software reuse information. ACM SIGSOFT Software Engineering Notes, 22(2):5, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poulin:1998:BRH

[Pou98] Jeffrey S. Poulin. Book reviews: High Performance Oracle8 Object-Oriented Design, David A. Anstey. ACM SIGSOFT Software Engineering Notes, 23(3):127–128, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pohl:1996:WSS

[PP96] Klaus Pohl and Peter Peters.

Workshop summary Second International Workshop on Requirements Engineering: Foundation of Software Quality. ACM SIGSOFT Software Engineering Notes, 21(1):31–34, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pastor:1997:UMM

[PP97] Esteban A. Pastor and R. T.

Price. Using metamodels of methodologies to determine the needs for reusability support. ACM SIGSOFT Software Engineering Notes, 22(3):121–129, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

pratimPal:1998:FAL

[pP98] Partha pratim Pal. A flexible, applet-like software distribution mechanism for Java applications. ACM SIGSOFT Software Engineering Notes, 23 (4):56–60, July 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Polo:1999:RMP

[PPRC99] Macario Polo, Mario Piattini, Francisco Ruiz, and Coral Calero. Roles in the maintenance process. ACM SIGSOFT Software Engineering Notes, 24 (4):84–86, July 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Philip:1995:RFS

[PR95] Thomas Philip and Ramani Ramsundar. A reengineering framework for small scale software. ACM SIGSOFT Software Engineering Notes, 20(5): 51–55, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

P:1999:MDT

[PR99] Puppydog P. O. P. and L. B. S. Raccoon. Multiple downcasting techniques. ACM SIGSOFT Software Engineering Notes, 24 (3):69–75, May 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prechelt:1999:SDS

[Pre99a] Lutz Prechelt. The surprising dynamics of a simple year 2000 bug. ACM SIGSOFT Software Engineering Notes, 24 (3):56–57, May 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Preston:1999:NTS

[Pre99b] Nicky Williams Preston. New type signatures for legacy Fortran subroutines. ACM SIG-SOFT Software Engineering Notes, 24(5):76–85, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Picco:1997:ECM

[PRM97] Gian Pietro Picco, Gruia-Catalin Roman, and Peter J. McCann. Expressing code mobility in mobile UNITY. ACM SIGSOFT Software Engineering Notes, 22(6):500– 518, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pohl:1995:WSF

[PSP95] Klaus Pohl, Gernot Starke, and Peter Peters. Workshop summary: First International Workshop on Requirements Engineering: Foundation of Software Quality (REFSQ;94). ACM SIGSOFT Software Engineering Notes, 20 (1):39–45, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Porter:1995:EAC

[PSTV95] A. Porter, H. Siy, C. A. Toman, and L. G. Votta. An experiment to assess the costbenefits of code inspections in large scale software development. ACM SIGSOFT Software Engineering Notes, 20(4): 92–103, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prince:1999:DIO

[PSTZ99] Robert Prince, Jianwen Su, Hong Tang, and Yonggang Zhao. The design of an interactive online help desk in the Alexandria digital library. ACM SIGSOFT Software Engineering Notes, 24(2): 217–226, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poulin:1994:WAW

[PT94] Jeff Poulin and Will Tracz. WISR’93: 6th Annual Workshop on Software Reuse: summary and working group reports. ACM SIGSOFT Software Engineering Notes, 19(1): 55–71, January 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Piattini:1997:ASM

[PT97] Mario G. Piattini and Hervé Tigréat. Applying the “STU-DIO” method to the interface design of an environmental software system. ACM SIGSOFT Software Engineering Notes, 22(4):81–83, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perry:1997:TTP

[PV97] Dewayne E. Perry and Lawrence G.

Votta. The tale of two projects (abstract). ACM SIGSOFT Software Engineering Notes, 22 (6):521–522, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Perry:1992:FSS

[PW92] Dewayne E. Perry and Alexander L. Wolf. Foundations for the study of software architecture. ACM SIGSOFT Software Engineering Notes, 17(4): 40–52, October 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Poulin:1995:MSA

[PW95] Jeffrey S. Poulin and Keith J.

Werkman. Melding structured abstracts and world wide Web for retrieval of reusable components. ACM SIGSOFT Software Engineering Notes, 20 (SI):160–168, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pohl:1997:CAP

[PW97] Klaus Pohl and Klaus Weidenhaupt. A contextual approach for process-integrated tools. ACM SIGSOFT Software Engineering Notes, 22(6): 176–192, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Podgurski:1993:PTS

[PY93] Andy Podgurski and Charles Yang. Partition testing, stratified sampling, and cluster analysis. ACM SIGSOFT Software Engineering Notes, 18(5): 169–181, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pezze:1996:GMF

[PY96] Mauro Pezzè and Michal Young. Generation of multiformalism state-space analysis tools. ACM SIGSOFT Software Engineering Notes, 21 (3):172–179, May 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Qian:1990:CMP

[QJD90] Xiaolei Qian, Richard Jullig, and Marilyn Daum. Consistency management in a project management assistant. ACM SIGSOFT Software Engineering Notes, 15(6):34–43, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Quester:1992:OCF

[Que92]  Remco Quester.  obTIOS:

a CAx-framework service for building concurrent engineering environments. ACM SIG-SOFT Software Engineering Notes, 17(5):32–40, December 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:1995:CMC

[Rac95a]  L. B. S. Raccoon.  The

chaos model and the chaos cycle. ACM SIGSOFT Software Engineering Notes, 20(1): 55–66, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:1995:CS

[Rac95b] L. B. S. Raccoon. The chaos strategy. ACM SIGSOFT Software Engineering Notes, 20(5): 40–47, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:1995:CG

[Rac95c] L. B. S. Raccoon. The complexity gap. ACM SIGSOFT Software Engineering Notes, 20 (3):37–44, July 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:1996:LCP

[Rac96] L. B. S. Raccoon. A learning curve primer for software engineers. ACM SIGSOFT Software Engineering Notes, 21(1): 77–86, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:1997:FYP

[Rac97] L. B. S. Raccoon. Fifty years of progress in software engineering. ACM SIGSOFT Software Engineering Notes, 22(1):

88–104, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:1998:TTS

[Rac98] L. B. S. Raccoon. Toward a tradition of software engineering. ACM SIGSOFT Software Engineering Notes, 23(3):105–110, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rajanna:1998:TDG

[Raj98] V. Rajanna. Test data generation. ACM SIGSOFT Software Engineering Notes, 23(5): 65–68, September 1, 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramachandran:1996:RDS

[Ram96] Muthu Ramachandran. Requirementsdriven software test: a processoriented approach. ACM SIG-SOFT Software Engineering Notes, 21(4):66–70, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raynham:1996:BRR

[Ray96] Peter Raynham. Book review: Rapid Software Development with Smalltalk. ACM SIGSOFT Software Engineering Notes, 21(4):104, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raynham:1997:BRB

[Ray97a] Peter Raynham. Book review:

Book two of Object Oriented Knowledge: The Working Object. ACM SIGSOFT Software Engineering Notes, 22(2): 101–102, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raynham:1997:BRP

[Ray97b] Peter Raynham. Book reviews:

Programming from Specifications, Second Edition, Carroll Morgan. ACM SIGSOFT Software Engineering Notes, 22(1): 108–109, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raynham:1998:BRO

[Ray98a] Peter Raynham. Book reviews:

Object Oriented COBOL, Edmund C. Arranga and Frank P. Coyle. ACM SIGSOFT Software Engineering Notes, 23 (3):128–129, May 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raynham:1998:BRS

[Ray98b] Peter Raynham. Book reviews: Smalltalk, Objects, and Design, by Chamond Liu. ACM SIGSOFT Software Engineering Notes, 23(1):110, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reps:1997:UPP

[RBDL97] Thomas Reps, Thomas Ball, Manuvir Das, and James Larus. The use of program profiling for software maintenance with applications to the year 2000 problem. ACM SIGSOFT Software Engineering Notes, 22 (6):432–449, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ram:1998:IMD

[RDO98] D. Janaki Ram, R. A. Dwivedi, and Ramakrishna Ongole. An implementation mechanism for design patterns. ACM SIG-SOFT Software Engineering Notes, 23(5):52–56, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rechtin:1996:SSA

[Rec96]  Eberhardt Rechtin.  Soft-

ware systems architecting (abstract). ACM SIGSOFT Software Engineering Notes, 21 (6):1, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reddy:1990:FMT

[Red90] Uday S. Reddy. Formal methods in transformational derivation of programs. ACM SIG-SOFT Software Engineering Notes, 15(4):104–114, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reifer:1997:RAC

[Rei97] Donald J. Reifer. Report: 4th ACM conference on computer and communications security. ACM SIGSOFT Software Engineering Notes, 22(4):32–33, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rementeria:1997:SPO

[Rem97] Santiago Rementeria. Software process in the organizational context: more data on European practices. ACM SIGSOFT Software Engineering Notes, 22(2):6–7, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Runciman:1990:FDS

[RF90] Colin Runciman and Mike Firth. Formalised development of software by machine assisted transformation. ACM SIG-SOFT Software Engineering Notes, 15(4):115–117, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rozman:1992:USE

[RGR92] I. Rozman, J. Györkös, and K. Rizmañ. Understandability of the software engineering method as an important factor for selecting a case tool. ACM SIGSOFT Software Engineering Notes, 17(3):43–46, July 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ratel:1991:PVC

[RHR91] Christophe Ratel, Nicolas Halbwachs, and Pascal Raymond. Programming and verifying critical systems by means of the synchronous data-flow language LUSTRE. ACM SIG-SOFT Software Engineering Notes, 16(5):112–119, December 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reps:1994:SS

[RHSR94] Thomas Reps, Susan Horwitz, Mooly Sagiv, and Genevieve Rosay. Speeding up slicing. ACM SIGSOFT Software Engineering Notes, 19(5): 11–20, December 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richardson:1999:RIW

[RI99] Debra Richardson and Paola Inverardi. ROSATEA: International Workshop on the Role of Software Architecture in Analysis (and) Testing. ACM SIGSOFT Software Engineering Notes, 24(4):33–42, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rine:1991:PSS

[Rin91a]  David C. Rine.  A pro-

posed standard set of principles for object-oriented development. ACM SIGSOFT Software Engineering Notes, 16(1):

43–49, January 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rine:1991:SOH

[Rin91b]  David C. Rine.  A short

overview of a history of software maintenance: as it pertains to reuse. ACM SIGSOFT Software Engineering Notes, 16 (4):60–63, October 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rine:1992:ECT

[Rin92] David Rine. EIFFEL: a case tool supporting object-oriented software construction. ACM SIGSOFT Software Engineering Notes, 17(1):57, January 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rine:1993:NSE

[Rin93] David Rine. A note on software engineering, software systems engineering and software design. ACM SIGSOFT Software Engineering Notes, 18(4): 16–18, October 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rine:1996:SDO

[Rin96] David Rine. Structural defects in object-oriented programming. ACM SIGSOFT Software Engineering Notes, 21 (2):86–88, March 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ransom:1995:SSR

[RM95] Keith J. Ransom and Chris D. Marlin. Supporting software reuse within an integrated software development environment (position paper). ACM SIG-SOFT Software Engineering Notes, 20(SI):233–237, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramloll:1999:MEU

[RM99a] Rameshsharma Ramloll and John A. Mariani. Moksha: exploring ubiquity in event filtration-control at the multiuser desktop. ACM SIGSOFT Software Engineering Notes, 24 (2):207–216, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robillard:1999:AEF

[RM99b] Martin P. Robillard and Gail C. Murphy. Analyzing exception flow in Java programs. ACM SIGSOFT Software Engineering Notes, 24(6):322–337, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Roche:1994:SMM

[Roc94] John M. Roche. Software metrics and measurement principles. ACM SIGSOFT Software Engineering Notes, 19(1): 77–85, January 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rolling:1994:PAB

[Rol94] Walter A. Rolling. A preliminary annotated bibliography on domain engineering. ACM SIGSOFT Software Engineering Notes, 19(3):82–84, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Richardson:1992:DIP

[ROMA92] Debra J. Richardson, T. Owen O’Malley, Cynthia Tittle Moore, and Stephanie Leif Aha. Developing and integrating ProDAG in the Arcadia environment. ACM SIGSOFT Software Engineering Notes, 17(5):109– 119, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosenblum:1996:FMT

[Ros96] David S. Rosenblum. Formal methods and testing: why the state-of-the art is not the state-of-the practice. ACM SIGSOFT Software Engineering Notes, 21(4):64–66, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosen:1997:PCS

[Ros97] Clive Rosen. PLUNGE DA:

a case study. ACM SIGSOFT Software Engineering Notes, 22 (2):82–83, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:1998:MCH

[RP98] L. S. B. Raccoon and Puppydog P. O. P. A middle-out concept of hierarchy (or the problem of feeding the animals). ACM SIGSOFT Software Engineering Notes, 23(3):111–119, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reps:1995:PIC

[RR95] Thomas Reps and Genevieve Rosay. Precise interprocedural chopping. ACM SIGSOFT Software Engineering Notes, 20 (4):41–52, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ram:1997:POT

[RRG97] D. Janaki Ram, K. N. Anantha Raman, and K. N. Guruprasad. A pattern oriented technique for software design. ACM SIGSOFT Software Engineering Notes, 22(4):70–73, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rountev:1999:DFA

[RRL99] Atanas Rountev, Barbara G. Ryder, and William Landi. Data-flow analysis of program fragments. ACM SIGSOFT Software Engineering Notes, 24 (6):235–252, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rada:1997:NSE

[RS97] Roy Rada and James Schoening. New standards for educational technology. ACM SIGSOFT Software Engineering Notes, 22(4):9–10, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rushby:1997:SS

[Rus97] John Rushby. Subtypes for specifications. ACM SIGSOFT Software Engineering Notes, 22 (6):4–19, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rushby:1991:FVA

[RvH91] John Rushby and Friedrich von Henke. Formal verification of algorithms for critical systems. ACM SIGSOFT Software Engineering Notes, 16(5): 1–15, December 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosenblum:1996:PCE

[RW96] David S. Rosenblum and Elaine J. Weyuker. Predicting the cost-effectiveness of regression testing strategies. ACM SIGSOFT Software Engineering Notes, 21(6):118– 126, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosenblum:1997:DFI

[RW97] David S. Rosenblum and Alexander L. Wolf. A design framework for Internet-scale event observation and notification. ACM SIGSOFT Software Engineering Notes, 22(6): 344–360, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ryant:1995:CAS

[Rya95] Ivan Ryant. The correctly analyzed system which behaves incorrectly. ACM SIGSOFT Software Engineering Notes, 20 (2):58–61, April 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ryant:1996:LSM

[Rya96] Ivan Ryant. 7-layered (small mental exercise). ACM SIG-SOFT Software Engineering Notes, 21(4):93, July 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sockut:1997:IDI

[SAMS97] Gary H. Sockut, Helen P. Arzu, Robert W. Matthews, and David E. Shough. Issues in designing an information model for application development. ACM SIGSOFT Software Engineering Notes, 22 (2):25–34, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saradhi:1991:SDP

[Sar91]  Motamarri Saradhi.  Sys-

tems deployment planning and scheduling. ACM SIGSOFT Software Engineering Notes, 16 (3):85–87, July 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saradhi:1992:DVI

[Sar92a] Motamarri Saradhi. A diagnostic view on information technology. ACM SIGSOFT Software Engineering Notes, 17(4): 68–70, October 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saradhi:1992:SEP

[Sar92b] Motamarri Saradhi. Software engineering from a practical perspective. ACM SIGSOFT Software Engineering Notes, 17 (3):21–26, July 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saradhi:1992:SMD

[Sar92c]  Motamarri Saradhi.  Sys-

tems modelling and description. ACM SIGSOFT Software Engineering Notes, 17 (2):57–63, April 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saradhi:1993:DCP

[Sar93] Motamarri Saradhi. Database conversion planning. ACM SIGSOFT Software Engineering Notes, 18(1):59–64, January 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saradhi:1994:REI

[Sar94]  Motamarri Saradhi.  Re-

engineer: If it wasn’t engineered! ACM SIGSOFT Software Engineering Notes, 19 (2):17, April 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sogaard-Anderson:1991:SKD

[SARL91] Jørgen F. Søgaard-Anderson, Camilla Østerberg Rump, and Hans Henrik Løvengreen. A systematic kernel development. ACM SIGSOFT Software Engineering Notes, 16(5):55–65, December 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:1998:MWNa

[SB98a] Yogesh Singh and Pradeep Bhatia. Module weakness: a new measure. ACM SIGSOFT Software Engineering Notes, 23 (4):82, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:1998:MWNb

[SB98b] Yogesh Singh and Pradeep Bhatia. Module weakness: a new measure. ACM SIGSOFT Software Engineering Notes, 23 (5):81, September 1, 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharble:1993:OOB

[SC93a] Robert C. Sharble and Samuel S.

Cohen. The object-oriented brewery: a comparison of two object-oriented development methods. ACM SIGSOFT Software Engineering Notes, 18(2): 60–73, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stocks:1993:TTF

[SC93b] P. Stocks and D. Carrington.

Test template framework: a specification-based testing case study. ACM SIGSOFT Software Engineering Notes, 18(3): 11–18, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Strens:1998:URM

[SC98] M. R. Strens and J. S. Chudge.

Using responsibility modeling to match organizational change to changes in the systems development process. ACM SIGSOFT Software Engineering Notes, 23(3):102–104, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scarlato:1991:DAS

[Sca91] Philip P. Scarlato. Developing Ada software without target hardware. ACM SIGSOFT Software Engineering Notes, 16 (1):36–40, January 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Siff:1999:CTC

[SCB^{+}99] Michael Siff, Satish Chandra, Thomas Ball, Krishna Kunchithapadam, and Thomas Reps. Coping with type casts in C. ACM SIGSOFT Software Engineering Notes, 24(6): 180–198, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:1992:DMS

[Sch92a]  Wilhelm Schaefer.  Design

methods and software processes. ACM SIGSOFT Software Engineering Notes, 17(1): 50–51, January 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaschinger:1992:EES

[Sch92b] Harald Schaschinger. ESA — an expert supported OOA method and tool. ACM SIG-SOFT Software Engineering Notes, 17(2):50–56, April 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schauer:1997:EFI

[Sch97] Helmut Schauer. ESEC97/ FSE5 invitation. ACM SIG-SOFT Software Engineering Notes, 22(1):6, January 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scott:1992:CSE

[Sco92] Gregory J. Scott. Can software engineering afford to improve the process? ACM SIGSOFT Software Engineering Notes, 17 (2):39–42, April 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Slovin:1994:EEE

[SD94] Malcolm Slovin and Donn Di Nunno. Engineering the enterprise: evolving quality systems. ACM SIGSOFT Software Engineering Notes, 19(3):45–47, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sears:1997:WGH

[Sea97] Andrew Sears. Working group on HCI education (identifying & disseminating resources). ACM SIGSOFT Software Engineering Notes, 22(1):6, January 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sefcik:1994:CSF

[Sef94] John G. Sefcik. Critical success factors for implementing software quality plans. ACM SIGSOFT Software Engineering Notes, 19(1):72–74, January 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sejnowski:1995:RDS

[Sej95] Matt Sejnowski. Revolutionary DCF system to replace CMM. ACM SIGSOFT Software Engineering Notes, 20(2):16–17, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Spanoudakis:1997:VIW

[SFE97] George Spanoudakis, Anthony Finkelsteiin, and Wolfgang Emmerich. Viewpoints 96: International Workshop on Multiple Perspectives in Software Development (SIGSOFT 96) workshop report. ACM SIGSOFT Software Engineering Notes, 22 (1):39–41, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sheth:1997:RNW

[SGJ^{+}97] Amit Sheth, Dimitrios Georgakopoulos, Stef M. M. Joosten, Marek Rusinkiewicz, Walt Scacchi, Jack Wileden, and Alexander L. Wolf. Report from the NSF Workshop on Workflow and Process Automation in Information Systems. ACM SIGSOFT Software Engineering Notes, 22(1): 28–38, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sloane:1996:BTP

[SH96] Anthony M. Sloane and Jason Holdsworth. Beyond traditional program slicing. ACM SIGSOFT Software Engineering Notes, 21(3):180–186, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Samaraweera:1998:EFO

[SH98] L. G. Samaraweera and R. Harrison. Evaluation of the functional and object-oriented programming paradigms: a replicated experiment. ACM SIGSOFT Software Engineering Notes, 23(4):38–43, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:1994:FDC

[Sha94]  Anil Sharma.  Framework

to define CASE tool requirements for distributed environment. ACM SIGSOFT Software Engineering Notes, 19(1): 86–89, January 1994. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaw:1995:AIS

[Sha95a] Mary Shaw. Architectural issues in software reuse: it’s not just the functionality, it’s the packaging. ACM SIG-SOFT Software Engineering Notes, 20(SI):3–6, 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaw:1995:BOS

[Sha95b] Mary Shaw. Beyond objects: a software design paradigm based on process control. ACM SIGSOFT Software Engineering Notes, 20(1):27–38, January 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See response [AA95].

Shaw:1998:PSE

[Sha98] Mary Shaw. A profession of software engineering: is there a need? YES: are we ready? NO. ACM SIGSOFT Software Engineering Notes, 23(6): 207–208, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shimomura:1991:ARM

[Shi91]  Takao Shimomura.  Auto-

matic resumption mechanism for program debugging. ACM SIGSOFT Software Engineering Notes, 16(4):37–41, October 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sutton:1990:LCM

[SHO90] Stanley M. Sutton, Jr., Dennis Heimbigner, and Leon J. Osterweil. Language constructs for managing change in process-centered environments.

ACM SIGSOFT Software Engineering Notes, 15(6):206–217, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shu:1993:IRC

[Shu93] William S. Shu. Inertia — the reluctance of code motion? ACM SIGSOFT Software Engineering Notes, 18(2):106–108, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Simos:1995:ODM

[Sim95] Mark A. Simos. Organization domain modeling (ODM): formalizing the core domain modeling life cycle. ACM SIGSOFT Software Engineering Notes, 20 (SI):196–205, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Simos:1996:DMR

[Sim96] Mark Simos. Domain modeling representation strategies: Towards a comparative framework. ACM SIGSOFT Software Engineering Notes, 21(1): 42–46, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sitaraman:1996:ISR

[Sit96] Murali Sitaraman. 4th international software reuse conference overview. ACM SIGSOFT Software Engineering Notes, 21 (1):15, January 1, 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sitaraman:1997:FIC

[Sit97] Murali Sitaraman. Fourth international conference on software reuse (ICSR4). ACM SIGSOFT Software Engineering Notes, 22(1):27, January 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shoemaker:1996:ESE

[SJ96] Dan Shoemaker and Vladan Jovanovic. Educating software executives: a doctorate in software management. ACM SIGSOFT Software Engineering Notes, 21(2):45–46, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sadler:1996:ESE

[SK96] Chris Sadler and Barbara Ann Kitchenham. Evaluating software engineering methods and tool — part 4: the influence of human factors. ACM SIG-SOFT Software Engineering Notes, 21(5):11–13, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Svoboda:1996:SWS

[SMDP96] Frank Svoboda, Fred Maymir-Ducharme, and Jeff Poulin. SRI workshop summary: “domain analysis in the DoD”. ACM SIGSOFT Software Engineering Notes, 21(1):55–67, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Simone:1999:IMA

[SMG99] Carla Simone, Gloria Mark, and Dario Giubbilei. Interoperability as a means of articulation work. ACM SIGSOFT Software Engineering Notes, 24 (2):39–48, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smith:1991:QAT

[Smi91] Richard E. Smith. Quality assurance through direct implementation of analysis and design constructs. ACM SIG-SOFT Software Engineering Notes, 16(3):93–94, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Smith:1996:FIC

[Smi96] Graeme Smith. First ISEW Cleanroom workshop summary. ACM SIGSOFT Software Engineering Notes, 21(4):41–44, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sullivan:1990:REI

[SN90] Kevin Sullivan and David Notkin. Reconciling environment integration and component independence. ACM SIG-SOFT Software Engineering Notes, 15(6):22–33, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schuster:1999:CMA

[SNS99] Hans Schuster, Jens Neeb, and Ralf Schamburger. A configuration management approach for large workflow management systems. ACM SIGSOFT Software Engineering Notes, 24(2): 177–186, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sutton:1997:DNG

[SO97] Stanley M. Sutton, Jr. and Leon J. Osterweil. The design of a next-generation process language. ACM SIGSOFT Software Engineering Notes, 22 (6):142–158, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Soloman:1998:NBD

[Sol98] Ad Soloman. A new bachelor’s degree program in software engineering. ACM SIGSOFT Software Engineering Notes, 23 (3):98–101, May 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sommerville:1996:SIW

[Som96] Ian Sommerville. Sixth International Workshop on Software Configuration Management. ACM SIGSOFT Software Engineering Notes, 21(4): 54–57, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Song:1995:FUI

[Son95] Xiping Song. A framework for understanding the integration of design methodologies. ACM SIGSOFT Software Engineering Notes, 20(1):46–54, January 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stevens:1998:SRP

[SP98] Perdita Stevens and Rob Pooley. Systems reengineering patterns. ACM SIGSOFT Software Engineering Notes, 23(6): 17–23, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Souter:1999:ICD

[SPH99] Amie L. Souter, Lori L. Pollock, and Dixie Hisley. Interclass def-use analysis with partial class representations. ACM SIGSOFT Software Engineering Notes, 24(5):47–56, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Seiter:1996:EOB

[SPL96] Linda M. Seiter, Jens Palsberg, and Karl J. Lieberherr. Evolution of object behavior using context relations. ACM SIGSOFT Software Engineering Notes, 21(6):46–57, November 1996. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Sprague:1991:RSC

[Spr91] Kristopher G. Sprague. The role of software configuration management in a measurement-based software engineering program. ACM SIGSOFT Software Engineering Notes, 16(2):62–66, April 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Siff:1996:PGS

[SR96] Michael Siff and Thomas Reps. Program generalization for software reuse: from C to C++. ACM SIGSOFT Software Engineering Notes, 21(6): 135–146, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stocks:1998:CFC

[SRLZ98] Philip A. Stocks, Barbara G. Ryder, William A. Landi, and Sean Zhang. Comparing flow and context sensitivity on the modification-side-effects problem. ACM SIGSOFT Software Engineering Notes, 23(2): 21–31, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schuetze:1997:PBA

[SRZ97] Martin Schuetze, Jan Peter Riegel, and Gerhard Zimmermann. A pattern-based application generator for building simulation. ACM SIGSOFT Software Engineering Notes, 22 (6):468–482, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Snodgrass:1990:FGD

[SS90] Richard Snodgrass and Karen Shannon. Fine grained data management to achieve evolution resilience in a software development environment. ACM SIGSOFT Software Engineering Notes, 15(6): 144–156, December 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Salasin:1995:EDC

[SS95] John Salasin and Howie Shrobe. Evolutionary design of complex software (EDCS). ACM SIGSOFT Software Engineering Notes, 20(5):18–22, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schappert:1995:ASS

[SSP95] Albert Schappert, Peter Sommerlad, and Wolfgang Pree. Automated support for software development with frameworks. ACM SIGSOFT Software Engineering Notes, 20 (SI):123–127, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Snelting:1998:RCH

[ST98] Gregor Snelting and Frank Tip. Reengineering class hierarchies using concept analysis. ACM SIGSOFT Software Engineering Notes, 23(6): 99–110, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Starke:1993:URI

[Sta93] Gernot Starke. Urgent research issues in software process engineering. ACM SIGSOFT Software Engineering Notes, 18(4): 13–15, October 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Steward:1993:SSM

[Ste93] Donald V. Steward. A simple straightforward method for software development. ACM SIGSOFT Software Engineering Notes, 18(4):36–43, October 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Steindl:1999:BDF

[Ste99] Christoph Steindl. Benefits of a data flow-aware programming environment. ACM SIGSOFT Software Engineering Notes, 24 (5):105–109, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sugiyama:1995:OMT

[Sug95] Yasuhiro Sugiyama. Object make: a tool for constructing software systems from existing software components. ACM SIGSOFT Software Engineering Notes, 20(SI):128– 136, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Suryanarayana:1997:BRO

[Sur97] Manjunath Suryanarayana.

Book review: Object-Orientation and Prototyping in Software Engineering. ACM SIGSOFT Software Engineering Notes, 22 (2):102, March 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Seemann:1998:PBD

[SvG98] Jochen Seemann and Jürgen Wolff von Gudenberg. Pattern-based design recovery of Java software. ACM SIGSOFT Software Engineering Notes, 23(6): 10–16, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stuurman:1998:LCM

[SvK98] Sylvia Stuurman and Jan van Katwijk. On-line change mechanisms: the software architectural level. ACM SIGSOFT Software Engineering Notes, 23 (6):80–86, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Srinath:1997:ISB

[SVR97] S. Srinath, K. Venkatesh, and D. Janaki Ram. An integrated solution based approach to software development using unified reuse artifacts. ACM SIGSOFT Software Engineering Notes, 22(4):56–60, July 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sitariman:1994:CBS

[SW94] Marulli Sitariman and Bruce Weide. Component-based software using RESOLVE. ACM SIGSOFT Software Engineering Notes, 19(4):21–22, October 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Swartz:1996:AAB

[Swa96]  A. John Swartz.  Airport

95: automated baggage system? ACM SIGSOFT Software Engineering Notes, 21(2): 79–83, March 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Samadzadeh:1996:SSR

[SZ96] M. Samadzadeh and M. K. Zand. SSR’95 summary report. ACM SIGSOFT Software Engineering Notes, 21(4):39–40, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Takao:1999:ENB

[Tak99] Shinji Takao. The effects of narrow-band width multipoint videoconferencing on group decision making and turn distribution. ACM SIGSOFT Software Engineering Notes, 24(2): 109–116, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tatsuta:1995:PVR

[Tat95]  Tanehiro Tatsuta.  Practi-

cal verification for requirements model of Hatley/Pirbhai. ACM SIGSOFT Software Engineering Notes, 20(3):94–97, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tatsuta:1996:CCA

[Tat96] Tanehiro Tatsuta. Conference on computer-aided software engineering summary report. ACM SIGSOFT Software Engineering Notes, 21(5): 36–39, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:1997:IPU

[Tay97] Richard Taylor. ICSE 97 program update. ACM SIGSOFT Software Engineering Notes, 22 (1):7, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thery:1992:RTP

[TBK92] Laurent Théry, Yves Bertot, and Gilles Kahn. Real theorem provers deserve real userinterfaces. ACM SIGSOFT Software Engineering Notes, 17 (5):120–129, December 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tarr:1993:POM

[TC93] Peri Tarr and Lori A. Clarke. PLEIADES: an object management system for software engineering environments. ACM SIGSOFT Software Engineering Notes, 18(5):56–70, December 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thibault:1997:FAG

[TC97] Scott Thibault and Charles Consel. A framework for application generator design. ACM SIGSOFT Software Engineering Notes, 22(3):131–135, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracey:1998:APF

[TCM98] Nigel Tracey, John Clark, and Keith Mander. Automated program flaw finding using simulated annealing. ACM SIGSOFT Software Engineering Notes, 23(2):73–81, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1993:DSS

[TCY93] Will Tracz, Lou Coglianese, and Patrick Young. A domainspecific software architecture engineering process outline. ACM SIGSOFT Software Engineering Notes, 18(2):40–49, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terwilliger:1990:OBE

[Ter90] Robert B. Terwilliger. An overview and bibliography of ENCOMPASS: an environment for incremental software development using executable, logicbased specification. ACM SIGSOFT Software Engineering Notes, 15(1):93–94, January 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terashima:1993:TAC

[Ter93a] Nobuyoshi Terashima. Toward automated communication software development. ACM SIGSOFT Software Engineering Notes, 18(1):49– 51, January 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terwilliger:1993:TTS

[Ter93b] Robert B. Terwilliger. Toward tools to support the Gries/ Dijkstra design process. ACM SIGSOFT Software Engineering Notes, 18(2):50–59, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thevenod-Fosse:1993:SAS

[TFW93] P. Thévenod-Fosse and H. Waeselynck. STATEMATE applied to statistical software testing. ACM SIGSOFT Software Engineering Notes, 18(3):99–109, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thompson:1999:SBP

[THM99] Jeffrey M. Thompson, Mats P. E. Heimdahl, and Steven P. Miller. Specification-based prototyping for embedded systems. ACM SIGSOFT Software Engineering Notes, 24(6): 163–179, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thompson:1998:WCS

[Tho98] Craig Thompson. Workshop on Compositional Software Architectures: workshop report. ACM SIGSOFT Software Engineering Notes, 23(3):44–63, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tichy:1993:SDW

[THP93] Walter F. Tichy, Nico Habermann, and Lutz Prechelt. Summary of the Dagstuhl Workshop on Future Directions in Software Engineering: February 17–21, 1992, Schloß Dagstuhl. ACM SIGSOFT Software Engineering Notes, 18 (1):35–48, January 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terry:1994:OTD

[THRE^{+}94] Allan Terry, Frederick Hayes-Roth, Lee Erman, Norman Coleman, Mary Devito, George Papanagopoulos, and Barbara Hayes-Roth. Overview of Teknowledge’s domain-specific software architecture program. ACM SIGSOFT Software Engineering Notes, 19(4):68– 76, October 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terry:1990:TSE

[TL90] B. Terry and D. Logee. Terminology for software engineering environment (SEE) and computer-aided software engineering (CASE). ACM SIG-SOFT Software Engineering Notes, 15(2):83–94, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tomic:1994:PAO

[Tom94] Marijana Tomic. A possible approach to object-oriented reengineering of Cobol programs. ACM SIGSOFT Software Engineering Notes, 19(2): 29–34, April 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1990:WDR

[Tra90] Will Tracz. Where does reuse start? ACM SIGSOFT Software Engineering Notes, 15(2): 42–46, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1991:CMM

[Tra91] Will Tracz. A conceptual model for megaprogramming. ACM SIGSOFT Software Engineering Notes, 16(3):36–45, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1992:DAW

[Tra92] Will Tracz. Domain analysis working group report: First International Workshop on Software Reusability. ACM SIG-SOFT Software Engineering Notes, 17(3):27–34, July 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1994:DSS

[Tra94] Will Tracz. Domain-specific software architecture (DSSA) frequently asked questions (FAQ). ACM SIGSOFT Software Engineering Notes, 19(2): 52–56, April 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1995:ICS

[Tra95a] Will Tracz. 17th international conference on software engineering: Window on the world. ACM SIGSOFT Software Engineering Notes, 20(3):16–28, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1995:BRC

[Tra95b] Will Tracz. Book review: Computer Related Risks by Peter G.

Neumann (ACM Press Books 1995). ACM SIGSOFT Software Engineering Notes, 20(1): 93, January 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1995:BRS

[Tra95c] Will Tracz. Book review: Software Engineering Standards and Specifications: an Annotated Index and Directory by Stan Magee and Leanard L. Trip. ACM SIGSOFT Software Engineering Notes, 20(1): 93, January 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1995:CUP

[Tra95d]  Will Tracz.  Confessions

of a used-program salesman: lessons learned. ACM SIG-SOFT Software Engineering Notes, 20(SI):11–13, 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1995:DDS

[Tra95e] Will Tracz. DSSA (domainspecific software architecture): pedagogical example. ACM SIGSOFT Software Engineering Notes, 20(3):49–62, July 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1995:TIC

[Tra95f] Will Tracz. Third international conference on software reuse summary. ACM SIGSOFT Software Engineering Notes, 20 (2):21–22, April 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1996:EDC

[Tra96a]  Will Tracz.  Evolutionary

Design of Complex Software (EDCS) kick off workshop summary. ACM SIGSOFT Software Engineering Notes, 21(5): 40–42, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1996:TAS

[Tra96b] Will Tracz. Test and analysis of software architectures. ACM SIGSOFT Software Engineering Notes, 21(3):1–3, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1997:BRA

[Tra97a] Will Tracz. Book review: The Art of Systems Architecting. ACM SIGSOFT Software Engineering Notes, 22(2): 102, March 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1997:DRJ

[Tra97b]  Will Tracz.  Developing

reusable Java components. ACM SIGSOFT Software Engineering Notes, 22(3):100–103, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1997:SPS

[Tra97c] Will Tracz. SIGSOFT policy statement for conferences and workshops sponsored by SIGSOFT. ACM SIGSOFT Software Engineering Notes, 22 (1):3–4, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1998:BRD

[Tra98a] Will Tracz. Book reviews:

The Deadline, A Novel about Project Management, Tom De-Marco. ACM SIGSOFT Software Engineering Notes, 23(1): 109, January 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:1998:HDP

[Tra98b] Will Tracz. Highlights from the DoD product line practice workshop product lines: bridging the gap — commercial success to DoD practice. ACM SIGSOFT Software Engineering Notes, 23(3):29–31, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thompson:1993:IFM

[TRC93] Margaret C. Thompson, Debra J. Richardson, and Lori A. Clarke. An information flow model of fault detection. ACM SIGSOFT Software Engineering Notes, 18(3):182–192, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tripp:1990:CES

[Tri90] Leonard L. Tripp. CASE evaluation and selection bibliography. ACM SIGSOFT Software Engineering Notes, 15(2):95, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tripp:1991:RFS

[Tri91] Leonard L. Tripp. Report on future software engineering standards direction. ACM SIGSOFT Software Engineering Notes, 16(3):77, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tripp:1992:WFS

[Tri92] Leonard L. Tripp. What is the future of software engineering standards?: discussion paper. ACM SIGSOFT Software Engineering Notes, 17(1): 58–61, January 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tripp:1993:SDP

[Tri93]  Leonard L. Tripp.  Stan-

dards development preference survey. ACM SIGSOFT Software Engineering Notes, 18(2): 109, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tahvanainen:1990:ACB

[TS90] Veli-Pekka Tahvanainen and Kari Smolander. An annotated CASE bibliography. ACM SIGSOFT Software Engineering Notes, 15(1):79–92, January 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tan:1990:CCO

[TSK90] Lichao Tan, Yoichi Shinoda, and Takuya Katayama. Coping with changes in an object management system based on attribute grammars. ACM SIGSOFT Software Engineering Notes, 15(6):56–65, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tedjini:1990:QSS

[TTB^{+}90] Mohamed Tedjini, Ian Thomas, Guy Benoliel, Fernando Gallo, and Régis Minot. A query service for a software engineering database system. ACM SIG-SOFT Software Engineering Notes, 15(6):238–248, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taylor:1995:SDU

[TTC95] Richard N. Taylor, Will Tracz, and Lou Coglianese. Software development using domainspecific software architectures: CDRl a011 — a curriculum module in the SEI style. ACM SIGSOFT Software Engineering Notes, 20(5):27–38, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Turpin:1993:LAD

[Tur93]  Russell Turpin.  A logical

approach to data structures. ACM SIGSOFT Software Engineering Notes, 18(5):138–148, December 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Toetenel:1990:SAF

[TvKP90] Hans Toetenel, Jan van Katwijk, and Nico Plat. Structured analysis — formal design, using stream and object oriented formal specifications. ACM SIGSOFT Software Engineering Notes, 15(4): 118–127, September 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Toleman:1992:EIM

[TWC92] Mark A. Toleman, Jim Welsh, and Alan J. Chapman. An empirical investigation of menu design in language-based editors. ACM SIGSOFT Software Engineering Notes, 17(5): 41–46, December 1992. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ulrich:1999:IWT

[UCD99] Andreas Ulrich and Gunther Chrobok-Diening. International Workshop on Testing Distributed Component-Based Systems. ACM SIGSOFT Software Engineering Notes, 24(4): 43–46, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Uhl:1997:WWE

[Uhl97] Jürgen Uhl. What we expect from software engineers in the industry. ACM SIGSOFT Software Engineering Notes, 22 (6):51, November 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Untch:1993:MAU

[UOH93] Roland H. Untch, A. Jefferson Offutt, and Mary Jean Harrold. Mutation analysis using mutant schemata. ACM SIGSOFT Software Engineering Notes, 18(3):139–148, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vanooteghem:1995:MMA

[Van95] Hubert Vanooteghem. The maturity movement and acedia. ACM SIGSOFT Software Engineering Notes, 20(2):18–19, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vazquez:1993:UOO

[Vaz93] Federico Vazquez. Using object oriented structured development to implement a hybrid system. ACM SIGSOFT Software Engineering Notes, 18(4): 44–53, October 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vazquez:1994:ICD

[Vaz94] Federico Vazquez. Identification of complete data flow diagrams. ACM SIGSOFT Software Engineering Notes, 19(3): 36–40, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vazquez:1996:AAD

[Vaz96] Federico Vazquez. An algebra approach to the deduction of data flow diagrams and object oriented diagrams from algebraic specifications. ACM SIGSOFT Software Engineering Notes, 21(4):71–80, July 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Villarreal:1997:RGD

[VB97] E. E. Villarreal and Don Batory. Rosetta: a generator of data language compilers. ACM SIGSOFT Software Engineering Notes, 22(3):146–156, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vandenBrand:1997:RES

[vdBKV97] M. G. J. van den Brand, P. Klint, and C. Verhoef. Reverse engineering and system renovation — an annotated bibliography. ACM SIGSOFT Software Engineering Notes, 22 (1):57–68, January 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vanderHoek:1997:SRM

[vdHHHW97] André van der Hoek, Richard S.

Hall, Dennis Heimbigner, and Alexander L. Wolf. Software release management. ACM SIG-SOFT Software Engineering Notes, 22(6):159–175, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Veryard:1996:ICM

[Ver96] Richard Veryard. Information coordination — the management of information model, systems, and organizations. ACM SIGSOFT Software Engineering Notes, 21(1):87–89, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vesterinen:1998:UID

[Ves98] Petri Vesterinen. Using inconsistent data for software process improvement. ACM SIGSOFT Software Engineering Notes, 23(4):53–55, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vesterinen:1999:ICE

[Ves99] Petri Vesterinen. Issues in calibrating effort estimation models. ACM SIGSOFT Software Engineering Notes, 24(3): 63–65, May 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vandevoorde:1994:USP

[VG94] Mark T. Vandevoorde and John V. Guttag. Using specialized procedures and specification-based analysis to reduce the runtime costs of modularity. ACM SIGSOFT Software Engineering Notes, 19 (5):121–127, December 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vazquez:1993:ASF

[VH93] Federico Vazquez and Guy Hanacek. Automatic system file generator. ACM SIGSOFT Software Engineering Notes, 18 (1):65–73, January 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Viljamaa:1995:PBI

[Vil95] Panu Viljamaa. The patterns business: impressions from PLoP-94. ACM SIGSOFT Software Engineering Notes, 20 (1):74–78, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vonMayrhauser:1991:TED

[vM91] Anneliese von Mayrhauser. Testing and evolutionary development. ACM SIGSOFT Software Engineering Notes, 16 (4):31–36, October 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

VanHilst:1996:DCD

[VN96] Michael VanHilst and David Notkin. Decoupling change from design. ACM SIGSOFT Software Engineering Notes, 21 (6):58–69, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vogel:1993:IGP

[Vog93] Peter A. Vogel. An integrated general purpose automated test environment. ACM SIGSOFT Software Engineering Notes, 18 (3):61–69, July 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Votta:1993:DEI

[Vot93] Lawrence G. Votta, Jr. Does every inspection need a meeting? ACM SIGSOFT Software Engineering Notes, 18(5): 107–114, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Valenti:1998:OCO

[VPC98] S. Valenti, M. Panti, and A. Cucchiarelli. Overcoming communication obstacles in user-analyst interaction for functional requirements elicitation. ACM SIGSOFT Software Engineering Notes, 23(1): 50–55, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Voas:1995:STE

[VPMM95] J. Voas, J. Payne, R. Mills, and J. McManus. Software testability: an experiment in measuring simulation reusability. ACM SIGSOFT Software Engineering Notes, 20(SI):247– 255, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vanReeken:1992:SNP

[vR92] A. J. van Reeken. Some notes on program design. ACM SIGSOFT Software Engineering Notes, 17(1):42, January 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See reply [McL92].

Vishik:1999:KSQ

[VW99] Claire Vishik and Andrew B. Whinston. Knowledge sharing, quality, and intermediation. ACM SIGSOFT Software Engineering Notes, 24(2): 157–166, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wahl:1999:ORT

[Wah99] Nancy J. Wahl. An overview of regression testing. ACM SIGSOFT Software Engineering Notes, 24(1):69–73, January 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Walters:1992:UHS

[Wal92] Neal Walters. Using harel statecharts to model objectoriented behavior. ACM SIG-SOFT Software Engineering Notes, 17(4):28–31, October 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ward:1992:NGC

[War92] Aidan Ward. The next generation of computer assistance for software engineering. ACM SIGSOFT Software Engineering Notes, 17(3):39–42, July 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wartik:1996:SYR

[War96] Steve Wartik. Slow down, you read too fast. ACM SIGSOFT Software Engineering Notes, 21 (1):73–74, January 1, 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wasmund:1995:SIR

[Was95] Michael Wasmund. The spinoff illusion: reuse is not a by-product. ACM SIGSOFT Software Engineering Notes, 20 (SI):219–221, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Waugh:1997:DET

[Wau97]  Doug Waugh.  Description

of EDCS technology clusters. ACM SIGSOFT Software Engineering Notes, 22(5):33–42, September 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Warboys:1999:CCI

[WBG^{+}99] B. C. Warboys, D. Balasubramaniam, R. M. Greenwood, G. N. C. Kirby, K. Mayes, R. Morrison, and D. S. Munro. Collaboration and composition: issues for a second generation process language. ACM SIGSOFT Software Engineering Notes, 24(6):75–90, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weyuker:1995:UCF

[Wey95] Elaine J. Weyuker. Using the consequence of failures for testing and reliability assessment. ACM SIGSOFT Software Engineering Notes, 20(4): 81–91, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weiss:1993:ISA

[WF93] Stewart N. Weiss and Vladimir N. Fleyshgakker. Improved serial algorithms for mutation analysis. ACM SIGSOFT Software Engineering Notes, 18 (3):149–158, July 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wermelinger:1999:ASA

[WF99] Michel Wermelinger and José Luiz Fiadeiro. Algebraic software architecture reconfiguration. ACM SIGSOFT Software Engineering Notes, 24(6):393– 409, November 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:1990:ELP

[WG90] Jeannette M. Wing and Chun Gong. Experience with the larch prover. ACM SIGSOFT Software Engineering Notes, 15 (4):140–143, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weske:1999:RMW

[WGHS99] Mathias Weske, Thomas Goesmann, Roland Holten, and Rüdiger Striemer. A reference model for workflow application development processes. ACM SIGSOFT Software Engineering Notes, 24(2):1–10, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Whiting:1999:SAH

[WH99] Liz Whiting and Mike Hill.

Safety analysis of Hawk In Flight monitor. ACM SIG-SOFT Software Engineering Notes, 24(5):32–38, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Whittle:1995:MLC

[Whi95] Ben Whittle. Models and languages for component description and reuse. ACM SIGSOFT Software Engineering Notes, 20 (2):76–89, April 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wiener:1992:TRS

[Wie92] Lauren Wiener. A trip report on SIGSOFT ’91. ACM SIGSOFT Software Engineering Notes, 17(2):23–38, April 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wile:1990:ARA

[Wil90] David S. Wile. Adding relational abstraction to programming languages. ACM SIG-SOFT Software Engineering Notes, 15(4):128–139, September 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Whittaker:1999:WSF

[WJ99] James A. Whittaker and Alan Jorgensen. Why software fails. ACM SIGSOFT Software Engineering Notes, 24(4):81–83, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:1997:TOO

[WKC^{+}97] Y. Wang, G. King, I. Court, M. Ross, and G. Staples. On testable object-oriented programming. ACM SIGSOFT Software Engineering Notes, 22 (4):84–90, July 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:1998:BTR

[WKP^{+}98] Yingxu Wang, Graham King, Dilip Patel, Ian Court, Geoff Staples, Margaret Ross, and Shushma Patel. On builtin tests and reuse in objectoriented programming. ACM SIGSOFT Software Engineering Notes, 23(4):60–64, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Woit:1993:SOP

[Woi93] Denise M. Woit. Specifying operational profiles for modules. ACM SIGSOFT Software Engineering Notes, 18(3):2–10, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wolf:1997:SSI

[Wol97]  Alexander L. Wolf.  Suc-

ceedings of the Second International Software Architecture Workshop (ISAW-2). ACM SIGSOFT Software Engineering Notes, 22(1):42–56, January 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wood:1990:AFM

[Woo90] William G. Wood. Application of formal methods to system and software specification. ACM SIGSOFT Software Engineering Notes, 15(4): 144–146, September 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wood:1993:PAS

[Woo93] Kenneth R. Wood. A practical approach to software engineering using z and the refinement calculus. ACM SIGSOFT Software Engineering Notes, 18 (5):79–88, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

White:1992:IHC

[WP92] Elizabeth L. White and James M. Purtilo. Integrating the heterogeneous control properties of software modules. ACM SIG-SOFT Software Engineering Notes, 17(5):99–108, December 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wood:1997:CCS

[WRBM97] Murray Wood, Marc Roper, Andrew Brooks, and James Miller. Comparing and combining software defect detection techniques: a replicated empirical study. ACM SIGSOFT Software Engineering Notes, 22 (6):262–277, November 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Waters:1991:TDA

[WT91] Richard C. Waters and Yang Meng Tan. Toward a design apprentice: supporting reuse and evolution in software design. ACM SIGSOFT Software Engineering Notes, 16(2):33–34, April 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:1991:AAB

[WTW^{+}91] Chih-Kan Wang, Duu-Chung Tsou, Rwo-Hsi Wang, James C. Browne, and Aloysius K. Mok. Automated analysis of bounded response time for two NASA expert systems. ACM SIG-SOFT Software Engineering Notes, 16(5):147–161, December 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:1995:MCS

[WVF95] Jeannette M. Wing and Mandana Vaziri-Farahani. Model checking software systems: a case study. ACM SIGSOFT Software Engineering Notes, 20(4):128–139, October 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:1998:LBE

[WW98] Chung-Hua Wang and Feng-Jian Wang. A languagebased editing process for visual object-oriented programming. ACM SIGSOFT Software Engineering Notes, 23(4): 70–75, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wong:1998:UDS

[WWBJ98] W. Eric Wong, Norman Wilde, Kristin Blackwell, and Rendy Justice. Understanding datasensitive code: one piece of the year 2000 puzzle. ACM SIG-SOFT Software Engineering Notes, 23(5):75–80, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wybolt:1990:ECO

[Wyb90] Nicholas Wybolt. Experiences with C++ and object-oriented software development. ACM SIGSOFT Software Engineering Notes, 15(2):31–39, April 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wybolt:1991:PCT

[Wyb91] Nicholas Wybolt. Perspectives on CASE tool integration. ACM SIGSOFT Software Engineering Notes, 16(3):56–60, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xiaoguang:1999:OOD

[XDB99] Zhu Xiaoguang, Wang Dongmu, and Hong Bingrong. An objectoriented data framework for virtual environments with hierarchical modeling. ACM SIGSOFT Software Engineering Notes, 24(1):65–68, January 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xiaodong:1998:CCO

[XDH^{+}98] Yuan Xiaodong, Hu Deqiang, Xu Hao, Li Yong, and Zheng Guoliang. COOZ: a complete object-oriented extension to z. ACM SIGSOFT Software Engineering Notes, 23(4):76–81, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xiaocong:1998:RCA

[XGG98] Fan Xiaocong, Chen Guanling, and Zheng Guoliang. Research on concurrent actions in multiagent systems. ACM SIGSOFT Software Engineering Notes, 23 (4):83, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xia:1998:WWS

[Xia98] Franck Xia. What’s wrong with software engineering research methodology. ACM SIGSOFT Software Engineering Notes, 23 (1):62–65, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xiaodong:1998:TDS

[XJG98] Yuan Xiaodong, Chen Jiajun, and Zheng Guoliang. Twodimensional software development model combining objectoriented method with formal method. ACM SIGSOFT Software Engineering Notes, 23(1): 81–85, January 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:1991:STC

[XP91] Jia Xu and David Lorge Parnas. On satisfying timing constraints in hard-real-time systems. ACM SIGSOFT Software Engineering Notes, 16(5): 132–146, December 1991. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:1995:ISS

[XZ95] Dianxiang Xu and Guoliang Zheng. Inheritance as sublation and its semantics. ACM SIGSOFT Software Engineering Notes, 20(5):62–66, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xu:1995:DJD

[XZLY95] Manwu Xu, Fancong Zeng, Jianfeng Lu, and Dong Yang. The development of JR-DSSG: a case study of software reuse. ACM SIGSOFT Software Engineering Notes, 20(5):67–72, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yuan:1998:DCC

[YCZ98] Xiaodong Yuan, Jiajun Chen, and Guoliang Zheng. Duration calculus in COOZ. ACM SIGSOFT Software Engineering Notes, 23(3):78, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ying:1996:SCS

[YH96] Jing Ying and Zhijun He. Some comments on software methodology versus CASE. ACM SIG-SOFT Software Engineering Notes, 21(5):58–59, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yang:1990:PIA

[YHR90] Wuu Yang, Susan Horwitz, and Thomas Reps. A program integration algorithm that accommodates semantics-preserving transformations. ACM SIG-SOFT Software Engineering Notes, 15(6):133–143, December 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Young:1990:VPS

[You90]  William D. Young.  Veri-

fied program support environments. ACM SIGSOFT Software Engineering Notes, 15(4): 147–149, September 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yang:1996:TSL

[YP96] Cheer-Sun D. Yang and Lori L.

Pollock. Towards a structural load testing tool. ACM SIGSOFT Software Engineering Notes, 21(3):201–208, May 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yang:1998:ADP

[YSP98] Cheer-Sun D. Yang, Amie L. Souter, and Lori L. Pollock. All-du-path coverage for parallel programs. ACM SIGSOFT Software Engineering Notes, 23 (2):153–162, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yeung:1990:IJD

[YT90] W. L. Yeung and G. Topping. Implementing JSD designs in Ada: a tutorial. ACM SIGSOFT Software Engineering Notes, 15(3):25–32, July 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yu:1991:VTR

[Yu91] Don Yu. A view on three R’s (3Rs): reuse, re-engineering, and reverse-engineering. ACM SIGSOFT Software Engineering Notes, 16(3):69, July 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zand:1997:RRD

[ZAD^{+}97] Mansour Zand, Gillermo Arango, Maggie Davis, Ralph Johnson, Jeffrey S. Poulin, and Andrew Watson. Reuse research and development: is it on the right track? ACM SIGSOFT Software Engineering Notes, 22 (3):212–216, May 1997. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zelkowitz:1994:AMR

[Zel94] Marvin V. Zelkowitz. Algebra and models (and reality). ACM SIGSOFT Software Engineering Notes, 19(4):79–81, October 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zelkowitz:1995:AMR

[Zel95a] Marvin V. Zelkowitz. Algebra and models (and reality). ACM SIGSOFT Software Engineering Notes, 20(2):55–57, April 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zeller:1995:UVM

[Zel95b] Andreas Zeller. A unified version model for configuration management. ACM SIGSOFT Software Engineering Notes, 20(4):151–160, October 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zeller:1999:YMP

[Zel99] Andreas Zeller. Yesterday, my program worked. today, it does not. why? ACM SIGSOFT Software Engineering Notes, 24 (6):253–267, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhou:1990:PSP

[Zho90] Wanlei Zhou. PM: a system for prototyping and monitoring remote procedure call programs. ACM SIGSOFT Software Engineering Notes, 15(1): 59–63, January 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zucconi:1996:BTS

[ZR96] Lin Zucconi and Karl Reed. Building testable software. ACM SIGSOFT Software Engineering Notes, 21(5):51–55, September 1996. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:1996:PDP

[ZRL96] Sean Zhang, Barbara G. Ryder, and William Landi. Program decomposition for pointer aliasing: a step toward practical analyses. ACM SIGSOFT Software Engineering Notes, 21 (6):81–92, November 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhao:1999:TWM

[ZS99] J. Leon Zhao and Edward A. Stohr. Temporal workflow management in a claim handling system. ACM SIGSOFT Software Engineering Notes, 24 (2):187–195, March 1999. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zucconi:1994:ICR

[Zuc94] Lin Zucconi. Issues concerning re-engineering of legacy software in the federal R&D environment. ACM SIGSOFT Software Engineering Notes, 19 (3):35, July 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zaremski:1993:SMK

[ZW93] Amy Moormann Zaremski and Jeannette M. Wing. Signature matching: a key to reuse. ACM SIGSOFT Software Engineering Notes, 18(5): 182–190, December 1993. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zaremski:1995:SMS

[ZW95] Amy Moormann Zaremski and Jeannette M. Wing. Specification matching of software components. ACM SIGSOFT Software Engineering Notes, 20 (4):6–17, October 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:1996:NOA

[ZW96] Jiazhong Zhang and Zhijian Wang. NDHORM: an OO approach to requirements modeling. ACM SIGSOFT Software Engineering Notes, 21(5): 65–69, September 1996. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zheng:1998:IFS

[ZZW98] Mingchun Zheng, Jiazhong Zhang, and Yanbing Wang.

Integrating a formal specification notation with HOOD. ACM SIGSOFT Software Engineering Notes, 23(5):57–61, September 1, 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).